### v1

# Large Language Models as Span Annotators

Abstract
For high-quality texts, single-score metrics seldom provide actionable feedback. In contrast, span annotation—pointing out issues in the text by annotating their spans—can guide improvements and provide insights. Until recently, span annotation was limited to human annotators or fine-tuned encoder models. In this study, we automate span annotation with large language models (LLMs). We compare expert or skilled crowdworker annotators with open and proprietary LLMs on three tasks: data-to-text generation evaluation, machine translation evaluation, and propaganda detection in human-written texts. In our experiments, we show that LLMs as span annotators are straightforward to implement and notably more cost-efficient than human annotators. The LLMs achieve moderate agreement with skilled human annotators, in some scenarios comparable to the average agreement among the annotators themselves. Qualitative analysis shows that reasoning models outperform their instruction-tuned counterparts and provide more valid explanations for annotations. We release the dataset of more than 40k model and human annotations for further research. [^1] Project website: https://llm-span-annotators.github.io
1 Introduction
Fine-grained aspects of texts, such as faithfulness or coherence, depend on local lexical choices. We need a detailed analysis in order to reflect them in the evaluation of natural language generation (NLG) systems or quality judgments on human-written texts. However, most automatic evaluation metrics are based on direct assessment, producing holistic scores for each evaluated aspect Gkatzia & Mahamood (2015); Sai et al. (2023); Schmidtová et al. (2024). Although numerical values make it easy to rank systems, these metrics are too simplistic and susceptible to biases or miscalibration issues of the underlying models Gehrmann et al. (2023); Gao et al. (2024); Liu et al. (2024); Wang et al. (2024).
The subject of our study, span annotation, offers an alternative approach. Figure 1 shows the difference between direct assessment and span annotation. Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans and classify them into predefined categories. In contrast to numerical rating, annotations are aligned to the evaluated text, which makes them more explainable and actionable. The annotations can also be examined post hoc, enabling a more modular evaluation process.
Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation. The method has traditionally relied on human annotators, making it costly and difficult to scale Da San Martino et al. (2019); Thomson & Reiter (2020); Popovic (2020); Kocmi et al. (2024b).
The situation has changed with LLM-as-a-judge paradigm, in which LLMs follow instructions for the evaluation process similarly to human annotators Zheng et al. (2023); Gu et al. (2024). However, the current LLM-as-a-judge applications still mostly focus on document-level classification or numerical ratings on the Likert scale, which are too coarse and unreliable for evaluating fine-grained aspects of texts.
Here, we investigate a specific LLM-as-a-judge application: using LLMs as span annotators. Instructing LLMs to annotate text spans with predefined categories is a way to produce span annotations automatically and robustly, lowering the barrier to using the method for NLG evaluation.
We focus on the following research questions:
  • RQ1: How to instantiate an LLM-based span annotator? We do a small-scale study of prompting techniques to arrive at a setup that is robust across diverse tasks and scenarios.
  • RQ2: How do LLMs compare to skilled human annotators? We compute inter-annotator agreement (IAA) between annotations produced by LLMs and skilled human annotators on both novel data and existing datasets.
  • RQ3: How valid are LLM annotations and their explanations? We manually analyze a subset of LLM and human annotations to assess the validity of annotations and their respective explanations.
For our experiments, we select three diverse tasks (cf. Section 3.1): evaluation of data-to-text generation Thomson & Reiter (2020), evaluation of machine translation Kocmi et al. (2024b), and propaganda detection Da San Martino et al. (2019). We base our span annotators on both open and proprietary state-of-the-art LLMs, including recent reasoning models. We collect more than 40k model and human annotations that we release for further investigation.
In the experimental part, we first analyze various prompting styles, demonstrating that zero-shot prompting with detailed guidelines is a robust way to set up LLMs as span annotators. Using automatic metrics, we show that LLMs are close to IAA of skilled human annotators, leveling or surpassing the performance of qualified crowdworkers in some scenarios. Qualitatively, we find that LLM annotations are completely correct in 49.5% of cases (56.5% in the case of reasoning models), which is similar to qualified crowdworkers.
Figure 1: Comparison between direct assesment of text (=conventional evaluation methods) and span annotation with pre-defined categories (=this paper). Span annotation provides more signal for diagnosing model outputs.
2 Related Work
LLMs for NLG Evaluation.
Automatic NLG metrics tradionally assess text quality by measuring similarity to human-written reference texts Sai et al. (2023); Schmidtová et al. (2024). As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgements Novikova et al. (2017); Reiter (2018). With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators across diverse tasks, either with simple numeric scoring (Bavaresco et al., 2024; Liu et al., 2023; Sottana et al., 2023; Leiter et al., 2023; Chiang & Lee, 2023) or with detailed feedback (Li et al., 2024; Kim et al., 2024a; b; Xu et al., 2023). However, the scores produced by LLMs still miss fine-grained text aspects and are influenced by LLM biases Stureborg et al. (2024); Koo et al. (2024); Wang et al. (2024).
Span Annotation Protocol.
In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024b), where annotators mark erroneous spans in translations. In data-to-text (D2T) generation, span annotation was applied by Thomson & Reiter (2020), who introduce a span-annotation-based evaluation protocol for annotation of generated basketball reports. Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection Da San Martino et al. (2019), and text summarization Subbiah et al. (2024). Unlike our work, these works conduct only human annotation.
Automating Span Annotation.
For all the aforementioned tasks, first attempts at automating span annotation were based on fine-tuned pre-trained encoder models. That includes MT (Guerreiro et al., 2024), D2T generation Kasner et al. (2021), text summarization Goyal et al. (2022), and propaganda detection Martino et al. (2020); Goffredo et al. (2023); Piskorski et al. (2023). Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities. While existing works (Kocmi & Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner & Dušek, 2024; Chang et al., 2024; Kartáč et al., 2025; Ramponi et al., 2025; Zouhar et al., 2024) applied LLMs as a tool for evaluating their systems, our study systematically compares LLMs to human annotators across tasks and domains.
See more discussion on related work in Appendix H.
3 Automating Span Annotation with LLMs
In this section, we first formally introduce the span annotation process in Section 3.1. Next, we discuss how to automate the process with LLMs in Section 3.2 and how to evaluate the quality of span annotations in Section 3.3.
3.1 Span Annotation: Task Definition
The goal of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$, i.e., to produce a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i}\rangle$:
  • $s_{i},e_{i}\in\{1,\ldots,n\},s_{i}<e_{i}$ are the start and end indices of the annotated span, [^2] Our definition allows overlapping annotations: $s_{i}\leq s_{j}\leq e_{i};i\neq j$.
  • $c_{i}\in C$ is the assigned annotation category from the pre-defined set of task-specific categories $C=\{c_{1},\ldots,c_{k}\}$.
The process is further specified in the guidelines $\mathcal{G}$: a set of fine-grained instructions for the annotation process regarding how to handle ambiguities or how to decide which parts of the text to include in the span. The annotator also considers the source $X$ (e.g., the translation source or the input data). Note that $X$ can be empty if we are annotating only intrinsic text aspects such as coherence or style.
3.2 Automating Span Annotations
Following our definition from Section 3.1, our goal is to collect the set of annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$. In our setup, we rely on the LLM-as-judge paradigm: prompting an LLM to follow annotation guidelines:
$A=\textsc{LLM}(\text{prompt}(Y,C,\mathcal{G},X)).$ (1)
To obtain the annotations, we closely follow the setup of Kasner & Dušek (2024). Specifically, we use structured decoding, asking for the list of annotations in JSON format. The output contains the fields reason (a short sentence justifying the annotation), text (the literal content of the matched span), and type (the integer index of the error category as given in the prompt). This approach enables us to robustly parse the model output and collect auxiliary explanations for annotations. [^3] See Appendix B for a detailed discussion of the selected approach and alternatives.
3.3 Evaluating Span Annotations
To determine and compare the quality of the human and the automatic annotation process, we need a notion of similarity between two sets of annotations $\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $Y=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$. Based on these similarities, we can then compute reliability metrics, such as IAA or comparison to gold annotations. We consider three similarity metrics.
Pearson correlation $\rho$ over counts.
The simplest correlation computation simply compares how many spans were annotated for each example.
$\mathrm{CountCorrelation}(A_{*},\hat{A}_{*})=\mathrm{Pearson}(\langle|A_{y}|,|\hat{A}_{y}|\rangle_{y\in Y})$ (2)
Precision, Recall, and $\mathbf{F_{1}}$.
To also quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined in Da San Martino et al. (2019). These measures are on matching annotations, adjusted to give partial credit to imperfect matches (which is necessary given the varying lengths of annotations): $\text{Precision}(A_{*},\hat{A}_{*})=\frac{1}{|A|}\sum_{a\in A}\max_{\hat{a}\in\hat{A}}\frac{|a\cap\hat{a}|}{|a|},$ (3) $\text{Recall}(A_{*},\hat{A}_{*})=\frac{1}{|\hat{A}|}\sum_{\hat{a}\in\hat{A}}\max_{a\in A}\frac{|a\cap\hat{a}|}{|\hat{a}|},$ (4) where $a\cap\hat{a}$ is the character overlap between two annotation spans and $|a|=e-s+1$ is the length of the annotation span in characters (see Section 3.1). Subsequently, we compute $F_{1}$-score as the harmonic mean of precision and recall.
For each of the metrics, we consider soft and hard variants. While the soft variant disregards span categories, the hard one only considers overlaps where the span category is matching. We consider the hard variant to be the default. Additionally, we report the difference $F_{1}\Delta=F_{1}(\text{soft})-F_{1}(\text{hard})$.
Gamma $\bm{\gamma}$.
The $F_{1}$ score is sensitive to varying span granularities and does not consider near matches with no overlap. It also does not account for agreement by chance. To this end, we follow Da San Martino et al. (2019) and Hasanain et al. (2024) in using the $\gamma$ score (Mathet et al., 2015) as another annotation similarity metric. This metric operates on whole sets of annotations for a single input $y$: $A_{*,y}=\{A,\hat{A}_{y},\ldots,A_{k,y}\}$. The metric builds the best possible alignment between the annotations $A_{i,y}$ and $A_{j,y}$ and computes their dissimilarity $\mathrm{Dissimilarity}(A_{i,y},A_{j,y})$. This is then compared to the expected dissimilarity by chance, which is created by sampling across all annotations: $\mathrm{DissimilarityChance}(A_{*,y})$. The final score is based on the average dissimilarity between annotations and the dissimilarity by chance:
$\gamma(A_{*,y})=1-\frac{\frac{1}{k^{2}}\sum_{i,j=1}^{k}\mathrm{Dissimilarity}(A_{i,y},A_{j,y})}{\mathrm{DissimilarityChance}(A_{*,y})}$ (5)
The score ranges from negative infinity to 1, where 1 is achieved when the annotations have perfect alignment. The $\gamma$ score extends Krippendorff’s $\alpha$ (Krippendorff, 1980), another popular metric, by computing the category-aware span alignments. We use the implementation of Titeux & Riad (2021).
Examples with empty annotation sets.
For an output $Y$, either one or both of the annotation sets $A_{y},\hat{A}_{y}$ can be empty. Note that this case is natural and may even be desirable: e.g., if the goal is to annotate errors in an output that is entirely correct. However, neither of the introduced scores accounts for these cases accordingly. [^4] The $\gamma$ score is undefined for less than two non-empty annotation sets: these examples thus need to be skipped during the $\gamma$ computation. The F1 score does not account for examples with no annotations. To compensate for that, we introduce a score $S_{\emptyset}$ that is computed for examples with less than two non-empty sets of annotations as follows:
$S_{\emptyset}=1/(1+|A|)$ (6)
where $|A|$ is the number of annotations from the remaining annotator. The score is equal to 1 for the cases where no annotator produced any annotations, i.e., a perfect match, and decreases proportionally to the number of annotations from the remaining annotator.
4 Experiments
Task # Cat. # Texts Avg. Len Task Setup Novel Data
D2T-Eval 6 1296 118/715 Kasner & Dušek (2024)
MT-Eval 2 2,854 26/185 Kocmi et al. (2024a) ×
Propaganda 18 100 914/4,659 Da San Martino et al. (2019) ×
Table 1: Overview of span annotation tasks used in our experiments. # Cat. denotes the number of categories used in the task (see Appendix E for their listings), # Texts the number of texts annotated, and Avg. Len the average number of words/characters in the output.
We cover three span annotation tasks of different quantitative and qualitative aspects: evaluation of data-to-text generation (D2T-Eval; Section 4.1), evaluation of machine translation (MT-Eval; Section 4.2), and propaganda detection (Propaganda; Section 4.3). In Table 1, we provide a quantitative overview of our tasks. [^5] See also Figure 1 and Appendix I for the visualization of examples of our tasks. Additionally, we discuss how we prevent the issue of data contamination for our tasks in Appendix C. We further describe our process of collecting annotations in Section 4.4.
4.1 Task: Evaluation of Data-to-text Generation
In D2T-Eval, we evaluate semantic accuracy and stylistic aspects of texts generated from structured data Celikyilmaz et al. (2020); Sharma et al. (2022). We apply span annotation to the text $Y$ generated from the input structured data $X$.
Subtasks and input data.
We select three subtasks of D2T generation: generating five-day weather forecasts (openweather), generating product descriptions (gsmarena), and generating reports of soccer games (football). We use the approach of Kasner & Dušek (2024) to download novel structured data from publicly available APIs, to make sure they are unseen for all LLMs (cf. Appendix C on data contamination). We provide more details on the data collection process in Section D.1.
Collecting model outputs.
We generate outputs for the structured inputs using LLMs. Specifically, we select two state-of-the-art models, Llama 3.3 70B Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 2B Team et al. (2024) and Phi 3.5 3.8B Abdin et al. (2024). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. [^6] See Sections A.2 and A.3 for details on the models and Appendix F for the prompts.
Data splits.
For further experiments, we divide the data into multiple subsets:
  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – the subset for LLM evaluation, annotated with crowdworkers.
  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – the subset for the study of prompt variants, annotated internally.
  • $\mathcal{D}_{\texttt{iaa}}$ (12 outputs) – the control subset for human crowdworkers, annotated internally.
Crowdsourcing annotations
We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com. We apply best practices for gathering human annotations, including an iterative process of refining annotation guidelines and pre-selecting best-performing annotators using a qualification task Tseng et al. (2020); Iskender et al. (2020); Huang et al. (2023); Zhang et al. (2023). Of the 230 annotators that participated in the qualification task, we selected 50 best annotators to annotate our data. We provide more details on the process of collecting human annotations in Section D.2.
4.2 Task: Evaluation of Machine Translation
For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024a). The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024b) using professional translators. We focus on character-level span annotations of major and minor translation errors provided in the dataset. The inputs $X$ are the source texts used to generate the translation $Y$ [^7] Note the annotations in MT-Eval cannot overlap and need not to be aligned with word boundaries.
We use data for the three textual domains present in the WMT 2024 shared task: news, literary, and social. We focus on data translated from English into other languages, including Chinese, Czech, German, Hindi, Icelandic, Japanese, Russian, Spanish, Ukrainian.
The original dataset has nearly 50k model outputs, making it too costly to run the full evaluation campaign. Therefore, we use all available system outputs for 10 sampled inputs for each language, giving us 2,854 outputs which we use for the experiments.
4.3 Task: Propaganda Technique Detection
For the Propaganda task, we use the dataset of Da San Martino et al. (2019). The dataset contains news collected from mostly propagandistic on-line sources. The token-level annotations in the dataset were created by expert annotators and cover 18 categories of logical fallacies and persuasion techniques. We use the test split for our experiments. Note that there are no inputs $X$ for the task, as all the annotated categories are intrinsic to the evaluated text $Y$.
4.4 Collecting LLM annotations
Models
For our span annotator LLMs, we use a mix of open and proprietary state-of-the-art models: (1) instruction-tuned models Llama 3.3 70B Grattafiori et al. (2024), GPT-4o Hurst et al. (2024), Claude 3.7 Sonnet Anthropic (2025), and (2) reasoning models: DeepSeek-R1 70B DeepSeek-AI et al. (2025), o3-mini OpenAI (2025) and Gemini 2.0 Flash Thinking Deepmind (2025).
We run the open models (Llama 3.3, DeepSeek-R1) locally in 4-bit quantization using ollama.
Prompts
We define several prompt variants for our experiments. $\mathcal{P}_{\texttt{base}}$ is the base prompt that includes the guidelines $\mathcal{G}$ as given to human annotators. By extending $\mathcal{P}_{\texttt{base}}$, we implement few-shot ($\mathcal{P}_{\texttt{5shot}}$) and chain-of-thought ($\mathcal{P}_{\texttt{cot}}$) prompts. We also ablate $\mathcal{P}_{\texttt{base}}$ by removing extended guidelines ($\mathcal{P}_{\texttt{noguide}}$) and not asking for explanations ($\mathcal{P}_{\texttt{noreason}}$). The full prompts can be found in Appendix F.
Annotation Tool
We use factgenie Kasner et al. (2024), a tool which supports both collecting span annotations from humans via a web interface and from LLMs via API calls.
5 Results
In this section, we describe the results of our experiments on span annotation with LLMs. We evaluated the LLM annotations both automatically using the metrics from Section 3.3 (Section 5.2) and manually (Section 5.3).
5.1 Prompting Techniques
To answer RQ1, we study the differences between prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ split using the open models (Llama 3.3 and DeepSeek-R1). The results are shown in Table 2.
Llama 3.3 DeepSeek-R1
Prompt $F_{1}$ $\gamma$a/o $F_{1}$ $\gamma$a/o
$\mathcal{P}_{\texttt{base}}$ 0.16 0.13 2.4 0.23 0.20 1.0
$\mathcal{P}_{\texttt{noguide}}$ 0.09 0.08 3.4 0.19 0.16 1.6
$\mathcal{P}_{\texttt{noreason}}$ 0.19 0.13 2.2 0.22 0.18 1.1
$\mathcal{P}_{\texttt{cot}}$ 0.08 0.10 0.8 0.21 0.19 1.1
$\mathcal{P}_{\texttt{5shot}}$ 0.23 0.18 2.5 0.20 0.16 1.4
Table 2: Comparison of different prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ split with Llama 3.3 and DeepSeek-R1: $F_{1}$ score, $\gamma$ score, and the average number of annotations per output (#a/o).
Including detailed guidelines seems beneficial: omitting the guidelines ($\mathcal{P}_{\texttt{noguide}}$) lowers the performance of both models. In contrast, not letting the model explain the annotation ($\mathcal{P}_{\texttt{noreason}}$) does not have a substantial effect on the model performance. For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce less annotations per example than the base variant (0.76 vs 2.42), leading to lower F1 and $\gamma$ scores. [^8] The model with $\mathcal{P}_{\texttt{cot}}$ often “overthinks” the annotations, deciding not to annotate cases of errors against which it can find some arguments. We provide an example of this behavior in Table 27. Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.999This observation is aligned with DeepSeek-AI et al. (2025), who note that few-shot prompting consistently degrades the performance of DeepSeek-R1. Following this preliminary study, we decided to focus on prompting with $\mathcal{P}_{\texttt{base}}$ for our main experiments.
5.2 LLM vs. Human Annotations
Figure 2: Comparison of LLMs and human annotators using $\mathcal{P}_{\texttt{base}}$. The stripe pattern denotes proprietary models, darker shades denote reasoning models. For $F_{1}$ score, the grey bar denotes its soft variant. The dashed horizontal lines are the average value of agreement between human annotators. Detailed results are included in Tables 8, 16 and 20.
To answer RQ2, we compare LLM annotations to human annotations using metrics described in Section 3.3. Wherever possible, we compare model results with an average IAA of human annotators. [^10] For D2T-Eval and MT-Eval, we estimate the bounds from the examples annotated by pairs of annotators. For Propaganda, we use the $\gamma$ score reported by Da San Martino et al. (2019) for agreement of the annotators before consolidation ($\gamma=0.31$). The overall results for all the tasks are outlined in Figure 2. We provide detailed results for individual tasks in Appendix G.
Figure 3: Confusion matrix of category counts averaged across models for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other, cf. Table 5).
Model football gsmarena openweath.
$F_{1}$ $\gamma$ $F_{1}$ $\gamma$ $F_{1}$ $\gamma$
Llama 3.3 0.27 0.20 0.08 0.08 0.10 0.05
GPT-4o 0.27 0.19 0.07 0.06 0.16 0.12
Claude 3.7 0.36 0.25 0.13 0.14 0.27 0.19
DeepS. R1 0.30 0.28 0.09 0.16 0.16 0.12
o3-mini 0.38 0.33 0.11 0.19 0.29 0.26
Gem. 2-FT 0.35 0.31 0.13 0.16 0.22 0.16
Table 3: Performance of models by domain (football, gsmarena, openweather) in D2T-Eval: $F_{1}$ score (hard), $\gamma$ score. The best results are bolded.
Performance is task-dependent.
LLMs provide annotations that are aligned with human annotations to a degree, with exact agreement depending on the task. For D2T-Eval, all models surpass the average $F_{1}$-scores (both soft and hard) between human annotators (0.13 and 0.21, respectively), although only o3-mini surpasses the $\gamma$ score (0.25). Performance is notably lower for Propaganda, where all models remain below the $\gamma$ agreement of human annotators. It should be noted that expert annotators were used in this task, which can make it more difficult for the models to reach their agreement level. This task also has the largest number of categories, manifested by a large increase in the soft $F_{1}$-score as opposed to its hard variant. On average, the models are also below human performance for MT-Eval, with Claude 3.7 showing the best results (0.22 $F_{1}$-score vs. 0.25 of human annotators).
Reasoning models outperform instruction-tuned models.
DeepSeek-R1, which is a reasoning model, outperforms the same-sized instruction-tuned Llama 3.3. The superiority of DeepSeek-R1 is most pronounced on D2T-Eval ($F_{1}$-score of 0.19 vs. 0.15, $\gamma$ score of 0.19 vs. 0.11). We hypothesize that this is because statements in texts generated from structured data often involve numerical reasoning. The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o despite o3-mini’s price per token being approximately 2x lower. Both findings demonstrate the potential of test-time scaling Snell et al. (2024); Welleck et al. (2024). A notable exception from this trend is Claude 3.7 Sonnet, which is mostly on par with o3-mini (although its price per token is approximately 3x higher).
Models mostly confuse related categories.
The results of category classification on D2T-Eval suggest that the models mostly tend to confuse annotation categories such as Contradictory, Not checkable and Misleading that are all related to semantic accuracy (cf. Figure 3 and Section G.8). This suggests that categorization errors may be less serious and may be related to category ambiguity or subjective understanding of category definitions. The results also show that the models use a less diverse distribution of categories, resorting to the Contradictory category more often than human annotators.
LLMs are more cost-efficient than human annotators.
Using LLMs, as opposed to human annotators, is notably different once the financial aspect is taken into account: For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance model o3-mini LLM costs \$3.60. We provide a more detailed cost- and time-wise analysis in Section A.1.
5.3 Manual Analysis of LLM Annotations
To answer RQ3, we manually analyzed the quality of LLM annotations on 216 samples. For each model, we sampled 3 annotations per error category in D2T-Eval and 1 annotation per error category in Propaganda. We show the results in Figure 4.
Overall, 49.5% of LLM-generated annotations and 50.5% of reasons were marked as correct for both tasks (with 9.2% of annotations and 12.5% of reasons additionally marked as partially correct). Reasoning models perform better, with 56.4% of their annotations and 58.3% marked as completely correct. The annotations made by Gemini 2.0 Flash Thinking and DeepSeek-R1 were the most accurate in D2T-Eval. o3-mini performed well across both tasks, even though Propaganda proved to be challenging for all models.
What are the sources of model errors?
Qualitatively, we find that the models often select wrong error categories despite identifying real issues (e.g. labeling contradictory statements as "incoherent"). Models also tend to be overly attentive, flagging noise in the data (quotes or non-article content) as errors, or marking slight numerical variations (such as rounded values) as misleading. All of these cases could be tackled by more descriptive guidelines or additional examples. However, in some cases, the models also misread or misinterpret data (e.g., claiming wind speed measurements do not exist when they do), which hints at deeper issues with understanding the data. Sources of incorrect reasons vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., counting omissions as sources of errors), to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors. Occasionally, the model admits that it marked a correct span as an error, such as in “The description of the game’s duration aligns with the data, providing coherent information”. [^11] Note that this typically happens to GPT-4o. Since OpenAI API ensures JSON key ordering, the reason had to be generated before the annotation (same as for the other models).
How good are human annotations?
Interestingly, the LLM annotations that were marked as correct have only 24% hard character-level overlap (51% soft) with human annotations. This fact led us to analyze the quality of human annotations in D2T-Eval. On a sample of 108 human annotations, we found that only 45.3% of annotations and 54.6% of reasons were marked as correct, which is comparable to LLM annotations. This suggests that the task is hard even for human annotators, and the quality of annotations from crowdworkers varies, even if they are preselected using a qualification task.
Figure 4: Results of our manual analysis. We analyzed 18 annotations and their explanations for each model and task (216 annotations in total). The color bars show annotations that we classified as correct, partially correct, assigned wrong category, incorrect, and undecidable cases. Detailed results are provided in Tables 22 and 23 in Appendix G.
6 Discussion
Can LLMs substitute human annotators?
Our results paint a complex picture in this regard. The IAA between LLMs and human annotators is generally moderate. However, we also need to take into account a realistic upper bound: the average IAA between human annotators themselves. In that case, the strongest LLMs have already reached the level of human annotators in D2T-Eval and are not far behind in MT-Eval and Propaganda. As we also showed, not all human annotations are “gold”, as even qualified crowdworkers make similar amounts of errors as LLMs. If we also factor in financial aspects, LLMs provide better flexibility, shorter response times, and much lower costs. In light of this, we argue that the benefits of automating span annotation with LLMs already outweigh its issues. Moreover, LLM-based span annotation will benefit from future increases in LLM capabilities.
Recommendations
To instantiate an LLM as a span annotator, it is beneficial to provide detailed guidelines describing how to handle all edge cases. In contrast, we note that providing specific examples (i.e., a few-shot setup) does not bring consistent improvements due to the length and complexity of typical examples for this task. When using LLMs in a new setup, we recommend validating model’s annotations against examples hand-annotated by experts on a sample of the data. As for choosing the model, we recommend using a reasoning model (as opposed to an equivalent model without reasoning), since the reasoning models provide more reliable annotations, although at the cost of slightly higher response times and token count.
Limitations
Span annotation as an evaluation method is not suitable for certain NLG evaluation tasks, such as annotating omissions or rating the overall text style. Although we tried to experiment with a representative sample of models, prompts, and tasks, our choice is necessarily constrained by our limited time frame and budget. Our estimates of the upper-bound IAA for each task are difficult to establish and depend on many factors, such as the chosen annotation categories, their inherent ambiguity, the annotation guidelines, or the qualification level of human annotators.
7 Conclusion
We demonstrated that LLMs can serve as span annotators for three span annotation tasks: evaluation of data-to-text generation, evaluation of machine translation, and propaganda detection in human-written texts. Our experiments show that LLMs achieve moderate agreement with skilled human annotators. The models perform best in D2T-Eval, where they are comparable to qualified crowdworkers. Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions. Overall, automating span annotation with LLMs offers a promising alternative to human annotation, opening the way for more fine-grained NLG evaluation methods.
Author Contributions
DG and SB first came up with the idea for the project, with SB further coordinating and overseeing the research process. ZK led the experimental design and execution part, including conducting both preliminary and main experiments, organizing the crowdsourcing campaigns, and processing the collected data. Multiple authors (DG, IK, KO, SB, SM, VZ, ZK) participated in the collection of gold data for D2T-Eval. Similarly, multiple authors (IK, KO, OD, OP, PS, SB, ZK) were involved in manual evaluation of the model outputs. DG provided financial resources for the Prolific campaigns. SM and SB provided expertise in preparing annotation guidelines and structuring the Prolific campaigns. Data processing and analysis were handled mainly by ZK, VZ, and PS, with VZ providing extra support with the WMT data. The paper was written by ZK, VZ, IK, PS, OD, and SB.
Acknowledgments
This work was funded by the European Union (ERC, NG-NLG, 101039303) and Charles University project SVV 260 698. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
We thank David M. Howcroft for his early input and contributions to the research methodologies adopted in this study.
Ethics Statement
Our human annotators were hired over Prolific and paid the platform-recommended wage of 9 GBP/hour (adjusted to slightly higher rates to account for real annotation times). The annotators were preselected based on their primary language (English). All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research (see Appendix D.2 for details). The domains were selected so that they do not contain any sensitive or potentially offensive content. We do not collect any demographic data about the participants.
Appendix A Implementation Details
A.1 Costs and Time Requirements
We estimate the cost and time requirements for a representative subset of annotators on D2T-Eval in Table 4.
Metric crowdw. Llama 3.3 DeepSeek-R1 Claude 3.7 Sonnet o3-mini
Cost per 1k outputs (\$) 500 - - 10.5 3.6
Time per output (s) 129.1 21.6 227.5 9.0 21.8
Table 4: Estimate of costs and time requirements for different kinds of annotators on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and two out of the proprietary models (Claude 3.7 Sonnet, and o3-mini).
Note that we do not estimate the costs of the models we run locally. In practice, the costs of infrastructure acquisition and inference (GPUs, electricity) needs to be taken into account. The time estimates for the open models were computed on the infrastructure we describe in Section A.2.
A.2 Open Models
We run the local models using the ollama framework in 4-bit quantization. Specifically, we use the following local models for span annotations:
We also use the following models for generating texts in D2T-Eval:
For better reproducibility, we set the seed to 42 and temperature to 0 for the local models. We do not use these parameters for proprietary models, as these parameters are generally not supported.
We run the models on a set of GPUs, including NVIDIA H100 NVL (95G), AMD MI210 (64G), and NVIDIA RTX 3090 (24G).
A.3 Proprietary Models
We use the following proprietary model versions:
  • GPT-4o: gpt-4o-2024-11-20
  • Claude 3.7 Sonnet: claude-3-7-sonnet-20250219
  • o3-mini: o3-mini-2025-01-31
  • Gemini 2.0 Flash Thinking: gemini-2.0-flash-thinking-exp-01-21
Appendix B Extracting Annotations from LLM Outputs
B.1 Tagging the Input Text
To the best of our knowledge, there is no established way of tagging the input text with decoder-only models. In the following, we discuss some of the existing approaches along with their pros and cons.
  1. 1.
    Labeling the text with special tokens Wang et al. (2023); Treviso et al. (2024): This approach instructs the LLM to generate the full copy of the input text. Spans are denoted using special XML-like tokens added to this copy. The approach is unambiguous in case the model copies the text correctly. However, it relies on the ability of the LLM to produce an exact copy of the (potentially very long) input text, and offers no fallback if the model fails to do that.
  2. 2.
    Listing start and end indices of spans Hasanain et al. (2024); Ramponi et al. (2025): This approach asks the LLM to generate numerical indices of the characters or tokens in the span. The approach would also make the output unambiguous, but it was discouraged by Hasanain et al., who notes that “[GPT-4] frequently generated indices not matching the corresponding spans in a paragraph”. Similarly, when models were asked to produce the output in CoNLL format Sang & Meulder (2003), “the models struggled to provide consistent outputs” Ramponi et al. (2025). A potential solution, providing the identifier for each token in the input text Ramponi et al. (2025), could be a confounding factor as it modifies the input text.
  3. 3.
    Listing textual content of spans Kasner & Dušek (2024); Kocmi & Federmann (2023); Zouhar et al. (2024): This approach asks the LLM to produce the textual content of the labeled spans along with their labels. The start and end indices are then found heuristically by case-insensitive string matching on the input text. This approach is more flexible, as the model does not need to copy the entire text, although it still relies on the ability of the models to literally copy text snippets. It can also misalign the model output by assigning the annotation to the first match in case there are multiple matches.
Of the available options, we select listing textual content of spans as our method of choice. Although we are aware of its disadvantages, our preliminary experiments showed that the failures are infrequent: the models are generally able to copy text snippets literally and string matching collisions are rare. This approach is also the most robust, as failing to match a single annotation does not influence other matches.
    "annotations": [{
      "reason": [REASON],
      "text": [SPAN_TEXT],
      "type": [CATEGORY_IDX]
Figure 5: Annotation output format for LLMs.
B.2 Structured Output
Having a high-level idea of the output format, we need to determine the exact format required from the model. Existing works collect model annotations using unstructured lists Kocmi & Federmann (2023); Zouhar et al. (2024) and structured output in the JSON format Kasner & Dušek (2024).
Following our preliminary experiments, we decide to closely follow the setup of Kasner & Dušek (2024), using structured decoding with a fixed JSON scheme as we describe in Section 3.2 (see Figure 5). Our rationale is that JSON outputs are easier and more robust to parse, while being a common format that is natural for the LLMs. Structured output is now also getting available in LLM frameworks. [^12] See, e.g., https://ollama.com/blog/structured-outputs.
Can structured output hurt model performance?
Some authors voice concerns about the diminished performance of LLMs when using JSON output Tam et al. (2024); Beurer-Kellner et al. (2024). However, reactions to Tam et al. Kurt (2024); Castillo (2024) suggest that these issues do not manifest if several factors—such as describing the JSON format in the prompt—are controlled for. We leave experiments on this issue as a separate topic for future work.
Parsing the output with reasoning chains
At the time we performed our experiments, some of the reasoning models either did not support structured output (Gemini 2.0 Flash Thinking) or it would cancel out their ability to perform reasoning chains (ollama models [^13] https://github.com/ollama/ollama/issues/8529). Therefore, we use the following heuristic: we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response. This approach turns out to be able to robustly parse model responses containing reasoning chains with minimum failures.
Appendix C Discussion on Data Contamination
It is common practice not to disclose the data used for LLM training. That applies not only to the prorietary models, but also to the majority of the open models. As discussed in several recent works Balloccu et al. (2024); Dong et al. (2024); Jiang et al. (2024b), it is mostly certain that the LLMs were trained on some test sets of benchmarks that were available before their knowledge cut-off date. Memorization of these test sets may then artificially inflate the model performance on these benchmarks.
In the context of D2T generation, Kasner & Dušek (2024) suggests to download a new set of structured data from publicly available sources – e.g., downloading up-to-date weather data for weather forecast generation. We use the method for D2T-Eval, where it creates only little overhead.
For the other tasks, we are not able to follow the same process due to the costs of data collection. Nevertheless, we believe that it does not influence the validity of our results:
  1. 1.
    For MT-Eval, we use the WMT 2024 dataset that was released at the end of 2024, which is after the knowledge cut-off date of the LLMs we are using.
  2. 2.
    For Propaganda, the dataset is in a format that cannot be easily memorized. The model would need to match the word-level annotations for token ids with the text, which (as we also discuss in Section B.2) is not typically within model capabilities.
  3. 3.
    Lastly, D2T-Eval serves as a control task. If the model would underperform on D2T-Eval, but perform well on other tasks where the datasets were known, then we could suspect that the models memorized the data. In practice, we see the opposite: the models have the lowest performance on Propaganda which is the only task with a publicly available dataset.
Appendix D D2T-Eval Details
In this section, we provide more details on the process of collecting data (Section D.1) and annotations (Section D.2) for D2T-Eval.
D.1 Collecting Data
Our goal was to collect a novel dataset for D2T-Eval that is not subject to data contamination (cf. Appendix C). To this end, we follow the approach of Kasner & Dušek (2024) to collect structured data from publicly available APIs.
Domains
The Quintd tool Kasner & Dušek (2024) enables downloading data for five domains: openweather, gsmarena, ice_hockey, our_world_in_data, and wikidata. We selected openweather and gsmarena as the most suitable for the data collection process. Additionally, we collect the new football domain from RapidAPI - API-Football. [^14] Our investigation showed that its structured data contain more detailed information about the game than ice_hockey and the sport itself is also more generally well-known.
Generating model outputs.
We leverage LLMs to get output texts for the dataset. Specifically, we prompt two state-of-the-art models, Llama 3.3 Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 Team et al. (2024) and Phi-3.5 Abdin et al. (2024) (cf. Section A.2). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. We provide the prompts in Appendix F.
Note on annotating models’ own outputs.
We point out that two of the four models that we use to collect annotations (Llama 3.3 and GPT-4o) are also later used to collect annotations. The fact that the models annotate their own outputs can reveal the self-preference bias Zheng et al. (2023); Wataoka et al. (2024). Since our main point is not to rank the quality of generated outputs, we do not consider it a flaw, but rather a suitable ground for future investigation. We hypothesize that since span annotations are more grounded than direct assessment (and not directly ranking models), the self-preference bias will not be as prominent here.
D.2 Collecting annotations
Here we describe our process of writing annotation guidelines and selecting crowdworkers in order to ensure high quality of the collected annotations for D2T-Eval.
Annotation guidelines
For the annotation guidelines, we (the authors of this paper) went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$. We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works Kasner & Dušek (2024); Thomson & Reiter (2020). Eventually, we settled on the list of annotation categories described in Appendix E (Table 5).
Our list is an extended and updated version of the error taxonomy of Kasner & Dušek (2024), covering also coherence categories (Incoherent, Repetitive) on top of semantic accuracy.
Additionally, we agreed upon a list of general recommendations for collecting the annotations:
  • The annotator should always annotate the longest continuous span (i.e., the whole fact instead of a single word).
  • The annotator should annotate according to their own knowledge and only the spans that they are sure about.
  • The annotator should ignore subjective statements such as "a lightweight smartphone".
  • The outputs in openweather can contain both precise numbers (e.g. 10.71°C) and the rounded ones (e.g. 11°C) as long as they agree with the data.
We provide these additional recommendations in $\mathcal{P}_{\texttt{base}}$, while we omit them in $\mathcal{P}_{\texttt{noguide}}$ (cf. Figures 7 and 8).
Gold annotations
Following the established annotation guidelines, we proceeded to collect our own internally-annotated gold data. The purpose of this data is two-fold: (1) to pre-select skilled crowdworkers, and (2) to create a smaller high-quality development set for the model prompting study. Since we also wanted to quantify our own internal IAA, we divided the data into two sets: $\mathcal{D}_{\texttt{dev}}$, containing 84 examples annotated individually by 7 annotators (12 examples per annotator), and $\mathcal{D}_{\texttt{dev}}$, containing 12 examples annotated commonly by each annotator. Our average IAA on $\mathcal{D}_{\texttt{dev}}$ was $F_{1}=0.433$ and $\gamma=0.399$ ([TODO: ] provides the average agreement of crowdworkers and models w.r.t. our annotations on this split).
Qualification task
To gather a pool of skilled annotators, we set up a qualification task on Prolific.com. For the study, we pre-selected workers whose first language is English, with >99% approval rate and more than 100 submissions. The workers were first presented with a detailed tutorial with annotation guidelines and examples of individual errors. After the tutorial, we tested the worker performance on five manually pre-selected examples from $\mathcal{D}_{\texttt{dev}}$. For annotating the data in $\mathcal{D}_{\texttt{test}}$, we selected workers that had $F_{1}$-score higher than 0.5 w.r.t. our ground truth solution to the qualification task. Of the 230 annotators who participated in the qualification task, 50 annotators (21.7%) passed this baseline. We pay all the annotators an average reward of 9.58 GBP / hour, regardless of the qualification outcome.
Annotating test data
We invited annotators who passed the qualification task to annotate $\mathcal{D}_{\texttt{test}}$. Out of 50 selected participants, 45 participated in the main task (=90% turnover rate). We presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{dev}}$ (i.e, the examples from $\mathcal{D}_{\texttt{dev}}$ that we did not use for the qualification task). All the 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator. Additionally, 475 outputs (39.6%) were annotated by an extra annotator. We payed all the annotators an average reward of 10.70 GBP / hour. We use examples with two annotators to compute the average IAA for D2T-Eval in Section 5.2.151515For other experiments, we use only the outputs from the first annotator as the reference data.
D.3 Annotation Interface
We implement our annotation interface using factgenie. Figure 6 shows the main annotation interface, including an example from the openweather domain.
Figure 6: Screenshot of the factgenie annotation interface used for collecting span annotations.
Appendix E Annotation Categories
Tables 5, 6 and 7 show an overview of annotation span categories that we used for our tasks along with their descriptions.
Category Name Description
Contradictory The fact contradicts the data.
Not checkable The fact cannot be verified from the data.
Misleading The fact is technically true, but leaves out important information or otherwise distorts the context.
Incoherent The text uses unnatural phrasing or does not fit the discourse.
Repetitive The fact has been already mentioned earlier in the text.
Other The text is problematic for another reason.
Table 5: Annotation categories for the D2T-Eval task.
Category Name Description
Major An error that disrupts the flow and makes the understandability of text difficult or impossible.
Minor An error that does not disrupt the flow significantly and what the text is trying to say is still understandable.
Table 6: Annotation categories for the MT-Eval task.
Category Name Description
Appeal to Authority Stating that a claim is true simply because a valid authority or expert on the issue said it was true, without any other supporting evidence offered. We consider the special case in which the reference is not an authority or an expert in this technique, altough it is referred to as Testimonial in literature
Appeal to fear-prejudice Seeking to build support for an idea by instilling anxiety and/or panic in the population towards an alternative. In some cases the support is built based on preconceived judgements
Bandwagon Attempting to persuade the target audience to join in and take the course of action because "everyone else is taking the same action"
Black-and-White Fallacy Presenting two alternative options as the only possibilities, when in fact more possibilities exist. As an the extreme case, tell the audience exactly what actions to take, eliminating any other possible choices (Dictatorship)
Causal Oversimplification Assuming a single cause or reason when there are actually multiple causes for an issue. It includes transferring blame to one person or group of people without investigating the complexities of the issue
Doubt Questioning the credibility of someone or something
Exaggeration, Minimisation Either representing something in an excessive manner: making things larger, better, worse (e.g., "the best of the best", "quality guaranteed") or making something seem less important or smaller than it really is (e.g., saying that an insult was just a joke)
Flag-Waving Playing on strong national feeling (or to any group; e.g., race, gender, political preference) to justify or promote an action or idea
Loaded Language Using specific words and phrases with strong emotional implications (either positive or negative) to influence an audience
Name Calling, Labeling Labeling the object of the propaganda campaign as either something the target audience fears, hates, finds undesirable or loves, praises
Obfuscation, Intentional Vagueness, Confusion Using words which are deliberately not clear so that the audience may have its own interpretations. For example when an unclear phrase with multiple definitions is used within the argument and, therefore, it does not support the conclusion
Red Herring Introducing irrelevant material to the issue being discussed, so that everyone“s attention is diverted away from the points made
Reductio ad hitlerum Persuading an audience to disapprove an action or idea by suggesting that the idea is popular with groups hated in contempt by the target audience. It can refer to any person or concept with a negative connotation
Repetition epeating the same message over and over again so that the audience will eventually accept it
Slogans A brief and striking phrase that may include labeling and stereotyping. Slogans tend to act as emotional appeals
Straw Men When an opponent“s proposition is substituted with a similar one which is then refuted in place of the original proposition
Thought-terminating Cliches Words or phrases that discourage critical thought and meaningful discussion about a given topic. They are typically short, generic sentences that offer seemingly simple answers to complex questions or that distract attention away from other lines of thought
Whataboutism A technique that attempts to discredit an opponent“s position by charging them with hypocrisy without directly disproving their argument
Table 7: Annotation categories for the MT-Eval task. The categories are adopted from Da San Martino et al. (2019).
Appendix F Prompts
Figures 7, 8, 9 and 10 show the prompts for the D2T-Eval that we use for the experiments in Section 5.1.
Figure 11 shows the base prompt we used for MT-Eval, and Figure 12 shows the base prompt we used for Propaganda.
Finally, Figure 13 shows the prompt we used for generating the outputs for D2T-Eval.
Your task is to identify errors in the text and classify them.
Output the errors as a JSON object with a single key "annotations". The value of "annotations" is a list in which each object contains fields "reason", "text", and "annotation_type". The value of "reason" is the short sentence justifying the annotation. The value of "text" is the literal value of the identified span (we will later identify the span using string matching). The value of "annotation_type" is an integer index of the error based on the following list:
{categories}
Examples:
- Contradictory: The lowest temperature does not drop below 4°C, but the text mentions 2°C. - Not checkable: The text mentions that "both teams display aggressive play", which cannot be checked from the data.
- Misleading: The tone of the text suggests there are many sensors out of which just a few are listed here. However, according to the data, the device has exactly these four sensors.
- Incoherent: The text states that both teams had "equal chances until the first half ended scoreless." While this is technically true, the expression sounds unnatural for a sport summary.
- Repetitive: The output text unnecessarily re-states information about a smartphone battery that was mentioned earlier.
- Other: Use this as a last resort when you notice something else not covered by the above categories.
Hints:
- Always try to annotate the longest continuous span (i.e., the whole fact instead of a single word).
- Annotate only the spans that you are sure about. If you are not sure about an annotation, skip it.
- Ignore subjective statements: for example "a lightweight smartphone" highly depends on the context: you should not annotate these statements.
- Numerical conventions: For weather forecasts, we accept both precise numbers (e.g. 10.71°C) and the rounded ones (e.g. 11°C) as long as they agree with the data.
- Annotate only according to your own knowledge. If you are considering using an external tool (Google, ChatGPT etc.), just skip that specific fact.
If there is nothing to annotate in the text, "annotations" will be an empty list.
Given the data:
{data} ```
annotate the errors in the corresponding text generated from the data:
{text} ```
Figure 7: The prompt $\mathcal{P}_{\texttt{base}}$ for D2T-Eval.
Your task is to identify errors in the text and classify them.
Output the errors as a JSON object with a single key "annotations". The value of "annotations" is a list in which each object contains fields "reason", "text", and "annotation_type". The value of "reason" is the short sentence justifying the annotation. The value of "text" is the literal value of the identified span (we will later identify the span using string matching). The value of "annotation_type" is an integer index of the error based on the following list:
{categories}
Given the data:
{data} ```
annotate the errors in the corresponding text generated from the data:
{text} ```
Figure 8: The prompt $\mathcal{P}_{\texttt{noguide}}$ for D2T-Eval.
Think about it step-by-step. You should enclose your chain of thoughts between the <think> and </think> tags. Once you are ready, output the JSON object in the required format.
Example:
<think> ... chain of thoughts ... </think> ... JSON object ... ```
Figure 9: The additional text added for $\mathcal{P}_{\texttt{cot}}$.
In order to help you with the task, we provide you with five examples of inputs, outputs and annotations:
Example #1: data:
{data} ```
text:
{text} ```
output:
{annotations} ``` (...)
Figure 10: The additional text added for $\mathcal{P}_{\texttt{5shot}}$.
Your task is to identify errors in the translation and classify them.
Output the errors as a JSON object with a single key "annotations". The value of "annotations" is a list in which each object contains fields "reason", "text", and "annotation_type". The value of "reason" is the short sentence justifying the annotation. The value of "text" is the literal value of the identified span (we will later identify the span using string matching). The value of "annotation_type" is an integer index of the error based on the following list:
{categories}
Error spans can include parts of the words, whole words, or multi-word phrases. Hint: errors are usually accuracy-related (addition, mistranslation, omission, untranslated text), fluency-related (character encoding, grammar, inconsistency, punctuation, register, spelling), style-related (awkward), terminology (inappropriate for context, inconsistent use).
Make sure that the annotations are not overlapping. If there is nothing to annotate in the text, "annotations" will be an empty list.
Given the source:
{source} ```
annotate its translation:
{text} ```
Figure 11: The prompt $\mathcal{P}_{\texttt{base}}$ for MT-Eval.
Your task is to identify spans of text that employ propaganda techniques.
Output the errors as a JSON object with a single key "annotations". The value of "annotations" is a list in which each object contains fields "reason", "text", and "annotation_type". The value of "reason" is the short sentence justifying the annotation. The value of "text" is the literal value of the identified span (we will later identify the span using string matching). The value of "annotation_type" is an integer index of the error based on the following list:
{categories}
Now annotate the following text:
{text} ```
Figure 12: The prompt $\mathcal{P}_{\texttt{base}}$ for Propaganda.
Given the structured summary of a football game:
{data} ``` Generate a match summary using approximately five sentences. The summary should sound natural, reporting on the important moments of the game. Avoid subjective statements, keep the tone of the summary neutral. Do not fabricate any facts that are not explicitly stated in the data.
Figure 13: The prompt used for generating outputs in the football domain for D2T-Eval. The prompts for the other domains are analogical. For more robust parsing, we initialize the model response with ’Sure, here is the summary: "’ .
Appendix G Results
In this section, we provide detailed results of our experiments.
The results for D2T-Eval are included in Sections G.1, G.2, G.3 and G.4:
The results for MT-Eval are included in Sections G.5 and G.6:
The results for Propaganda are included in Section G.7: Tables 20 and 21.
The confusion matrices are included in Section G.8: Figures 14, 15 and 16.
The manual evaluation results are included in Section G.9: Tables 22 and 23.
G.1 D2T-Eval – Main Results
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.307 0.132 0.276 0.185 0.388 0.154 0.323 0.169 0.109 0.418
GPT-4o 0.346 0.178 0.300 0.180 0.303 0.179 0.301 0.122 0.130 0.429
Claude 3.7 0.512 0.262 0.395 0.287 0.432 0.274 0.412 0.138 0.203 0.592
DeepS. R1 0.453 0.293 0.493 0.154 0.259 0.202 0.340 0.138 0.185 0.645
o3-mini 0.505 0.351 0.488 0.250 0.347 0.292 0.405 0.113 0.273 0.637
Gem. 2-FT 0.458 0.259 0.434 0.236 0.395 0.247 0.414 0.167 0.209 0.612
Table 8: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. See Figure 2 for visualizaton of this table.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 2981 2.5 28.8 50.3
Llama 3.3 3214 2.7 7.4 65.5
GPT-4o 2284 1.9 4.8 66.3
Claude 3.7 2865 2.4 22.5 57.2
DeepS. R1 1387 1.2 44.2 56.8
o3-mini 1836 1.5 35.6 58.0
Gem. 2-FT 2517 2.1 28.9 54.3
Table 9: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
G.2 D2T-Evalfootball domain
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.512 0.241 0.408 0.295 0.500 0.266 0.449 0.183 0.197 0.455
GPT-4o 0.422 0.271 0.394 0.263 0.382 0.267 0.388 0.121 0.190 0.427
Claude 3.7 0.612 0.330 0.466 0.384 0.542 0.355 0.501 0.146 0.254 0.587
DeepS. R1 0.561 0.414 0.643 0.237 0.368 0.301 0.468 0.167 0.276 0.659
o3-mini 0.610 0.431 0.558 0.344 0.445 0.382 0.495 0.113 0.328 0.645
Gem. 2-FT 0.588 0.387 0.571 0.314 0.464 0.347 0.512 0.165 0.307 0.647
Table 10: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the football domain of D2T-Eval.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 1053 2.6 26.8 54.1
Llama 3.3 977 2.4 6.2 71.5
GPT-4o 761 1.9 2.8 72.8
Claude 3.7 1051 2.6 13.8 63.1
DeepS. R1 486 1.2 41.2 67.1
o3-mini 642 1.6 30.5 70.8
Gem. 2-FT 755 1.9 29 61.3
Table 11: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the football domain of D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
G.3 D2T-Evalgsmarena domain
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.165 0.062 0.144 0.118 0.276 0.081 0.189 0.108 0.077 0.456
GPT-4o 0.127 0.068 0.169 0.078 0.195 0.072 0.181 0.109 0.062 0.453
Claude 3.7 0.214 0.146 0.290 0.121 0.240 0.133 0.263 0.130 0.140 0.656
DeepS. R1 0.223 0.183 0.373 0.059 0.121 0.090 0.183 0.093 0.160 0.686
o3-mini 0.172 0.201 0.429 0.078 0.167 0.113 0.240 0.127 0.192 0.689
Gem. 2-FT 0.234 0.179 0.352 0.098 0.193 0.127 0.249 0.122 0.161 0.675
Table 12: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the gsmarena domain of D2T-Eval.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 809 2.0 40.2 42.5
Llama 3.3 890 2.2 15 73.9
GPT-4o 670 1.7 10.5 59.2
Claude 3.7 584 1.5 47.2 48.9
DeepS. R1 208 0.5 69.2 53.8
o3-mini 272 0.7 66.2 49.1
Gem. 2-FT 419 1.1 54.2 45.1
Table 13: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the gsmarena domain of D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
G.4 D2T-Evalopenweather domain
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.209 0.092 0.270 0.118 0.346 0.103 0.303 0.200 0.053 0.272
GPT-4o 0.404 0.164 0.298 0.158 0.288 0.161 0.293 0.132 0.122 0.377
Claude 3.7 0.550 0.246 0.370 0.291 0.437 0.266 0.401 0.135 0.189 0.429
DeepS. R1 0.454 0.215 0.391 0.129 0.234 0.162 0.293 0.131 0.124 0.533
o3-mini 0.552 0.317 0.437 0.259 0.357 0.285 0.393 0.108 0.259 0.481
Gem. 2-FT 0.458 0.197 0.366 0.241 0.447 0.217 0.403 0.186 0.155 0.368
Table 14: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the openweather domain of D2T-Eval.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 1119 2.8 19.2 52.2
Llama 3.3 1347 3.4 1 55.7
GPT-4o 853 2.1 1.2 66.2
Claude 3.7 1230 3.1 6.5 56.1
DeepS. R1 693 1.7 22.2 50.5
o3-mini 922 2.3 10 51.8
Gem. 2-FT 1343 3.4 3.5 53.2
Table 15: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the openweather domain of D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
G.5 MT-Eval – Main Results
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.182 0.048 0.085 0.228 0.376 0.071 0.124 0.053 0.015 0.392
GPT-4o 0.158 0.063 0.107 0.169 0.291 0.081 0.137 0.057 0.076 0.428
Claude 3.7 0.301 0.117 0.174 0.304 0.435 0.151 0.223 0.071 0.131 0.628
DeepS. R1 0.177 0.077 0.121 0.119 0.187 0.084 0.132 0.047 0.058 0.631
o3-mini 0.197 0.085 0.143 0.137 0.234 0.096 0.162 0.066 0.100 0.646
Gem. 2-FT 0.257 0.104 0.172 0.117 0.211 0.102 0.174 0.071 0.066 0.710
Table 16: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 2405 0.62 70.04 13.97
Llama 3.3 6361 2.3 6.2 17.4
GPT-4o 4866 1.7 7.0 15.9
Claude 3.7 3782 1.4 30.6 15.9
DeepS. R1 2586 0.9 36.3 15.1
o3-mini 3039 1.1 35.8 13.8
Gem. 2-FT 2181 0.8 50.0 15.2
Table 17: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
G.6 MT-Eval – By Language
Language $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
en-cs 0.303 0.085 0.130 0.205 0.323 0.117 0.180 0.063 0.084 0.582
en-es 0.171 0.063 0.117 0.164 0.313 0.086 0.161 0.075 0.080 0.631
en-hi 0.170 0.042 0.059 0.160 0.237 0.064 0.091 0.027 -0.0 0.552
en-is 0.347 0.217 0.368 0.149 0.255 0.173 0.294 0.121 0.108 0.493
en-ja 0.127 0.033 0.046 0.256 0.320 0.058 0.078 0.020 0.063 0.569
en-ru 0.225 0.120 0.170 0.172 0.250 0.136 0.195 0.059 0.162 0.588
en-uk 0.192 0.063 0.097 0.172 0.276 0.089 0.139 0.050 0.031 0.542
en-zh 0.163 0.036 0.083 0.155 0.339 0.058 0.130 0.072 0.075 0.623
Table 18: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language (average across models).
Language Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
en-cs Model avg. 600 1.4 27.0 16.6
Human 399 0.7 66.1 13.0
en-es Model avg. 417 1.1 38.9 18.8
Human 248 0.6 70.3 10.3
en-hi Model avg. 396 1.3 26.2 19.0
Human 222 0.5 71.2 10.7
en-is Model avg. 563 1.9 14.3 15.7
Human 752 2.5 18.3 16.6
en-ja Model avg. 471 1.3 28.7 11.1
Human 118 0.2 87.5 14.8
en-ru Model avg. 500 1.3 25.9 18.2
Human 287 0.7 58.7 19.4
en-uk Model avg. 436 1.5 25.4 17.8
Human 208 0.7 64.3 12.3
en-zh Model avg. 420 1.2 34.6 7.2
Human 171 0.2 85.1 6.6
Table 19: Statistics of models (averaged) and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language. Ann=# of annotations, Ann/Ex=annotations per example. Ex. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
G.7 Propaganda – Main Results
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.336 0.069 0.253 0.063 0.231 0.066 0.242 0.176 0.092 0.343
GPT-4o 0.344 0.092 0.294 0.036 0.114 0.051 0.164 0.113 0.066 0.234
Claude 3.7 0.460 0.109 0.277 0.088 0.225 0.098 0.249 0.151 0.155 0.113
DeepS. R1 0.354 0.081 0.257 0.058 0.182 0.067 0.213 0.146 0.091 0.476
o3-mini 0.418 0.152 0.431 0.065 0.183 0.091 0.257 0.166 0.154 0.517
Gem. 2-FT 0.560 0.106 0.277 0.186 0.484 0.135 0.353 0.218 0.202 0.493
Table 20: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on Propaganda.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 1439 14.2 4.0 40.2
Llama 3.3 574 5.7 3.0 92.0
GPT-4o 246 2.4 8.9 91.1
Claude 3.7 803 8.0 7.9 58.5
DeepS. R1 459 4.5 9.9 89.3
o3-mini 376 3.7 10.9 65.3
Gem. 2-FT 1864 18.5 3.0 54.1
Table 21: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on Propaganda. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
G.8 Confusion Matrices
(a) Llama 3.3
(b) DeepSeek-R1
(c) GPT-4o
(d) o3-mini
(e) Gemini 2.0 Flash Thinking
(f) Claude 3.7 Sonnet
Figure 14: Confusion matrices comparing human annotations (rows) with model predictions (columns) for D2T-Eval. Categories: Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other (cf. Table 5). For the model average, see Figure 3.
(a) Llama 3.3
(b) DeepSeek-R1
(c) GPT-4o
(d) o3-mini
(e) Gemini 2.0 Flash Thinking
(f) Claude 3.7 Sonnet
Figure 15: Confusion matrices comparing human annotations (rows) with model predictions (columns) for MT-Eval. Categories: Major and Minor errors (cf. Table 6).
Figure 16: Confusion matrix of category counts averaged across models for Propaganda (see Table 7 for the description of categories).
G.9 Manual Analysis
Annotation Reason
Model C P W I U C P W I U
Llama 3.3 7 1 2 8 0 6 0 1 10 1
GPT-4o 5 2 1 10 0 6 1 0 11 0
Claude 3.7 7 2 3 6 0 9 2 0 7 0
DeepSeek 11 2 4 1 0 12 3 0 3 0
o3-mini 10 3 2 0 3 8 7 0 0 3
Gemini 2 F-T 12 4 1 1 0 12 5 1 0 0
Total 52 14 13 26 3 53 18 2 31 4
Table 22: Manual evaluation results for D2T-Eval domain. Categories for annotation and reason: C=Correct, P=Partially correct, W=Wrong category, I=Incorrect, U=Undecidable.
Annotation Reason
Model C P W I U C P W I U
Llama 3.3 9 1 2 5 1 8 2 1 6 1
GPT-4o 9 0 3 6 0 8 2 1 7 0
Claude 3.7 9 2 3 4 0 9 3 3 3 0
DeepSeek 6 0 4 8 0 7 0 3 8 0
o3-mini 15 0 0 2 1 15 0 0 2 1
Gemini 2 F-T 7 3 2 6 0 9 2 1 6 0
Total 55 6 14 31 2 56 9 9 32 2
Table 23: Manual evaluation results for Propaganda domain. Categories for annotation and reason: C=Correct, P=Partially correct, W=Wrong category, I=Incorrect, U=Undecidable.
Appendix H Related Work (extended)
This section provides more discussion on related work following Section 2.
Due to the complexity and often open-ended nature of generated text, NLG evaluation is inherently difficult (Celikyilmaz et al., 2020). Machine translation (MT) has historically played an important role in the development of both automated evaluation metrics and human annotation protocols. Although evaluation practices are less standardized in other tasks, human evaluation typically involves ranking outputs or assigning numeric scores for specific aspects such as fluency or faithfulness (Novikova et al., 2018; van der Lee et al., 2021; Kocmi et al., 2022).
Error Span Annotation.
As simple numeric scores can be subjective and unreliable, the field is shifting towards annotating erroneous spans. The MT community has adopted a detailed error taxonomy, known as MQM (Lommel et al., 2014; Mariana, 2014, Multidimensional Quality Metrics). In MQM, each erroneous span in the translation has an assigned category, such as accuracy/terminology and severity, such as minor/major/critical. The final translation score is defined as a weighted sum of the error spans based on the error categories and severities. However, this leads to many ties and is problematic when judging longer translations. In addition, this protocol requires a very expensive expert human labor. For these reasons, Kocmi et al. (2024b), based on prior work by Popovic (2020), introduce ESA (Error Span Annotation) for machine translation evaluation. This protocol drops the error categorization, which is primarily useful for diagnosis and not evaluation, simplifies severities, and has the annotators also assign the final score after they have done the span annotations. To make large-scale annotation more efficient, Zouhar et al. (2024) use a quality estimation system that pre-highlights potential error spans in the translation. The annotators then “post-edit” the error annotations and assign the final score.
In data-to-text, Thomson & Reiter (2020) introduce an evaluation protocol for high-quality annotation of semantic accuracy errors in generated basketball reports. Similarly to MQM, such an approach involves significant costs and time requirements, as the authors report. Beyond NLG evaluation, span annotation has also been applied in other areas, such as fine-grained propaganda detection. Da San Martino et al. (2019) build an annotated corpus with span annotations for 18 different propaganda techniques, followed by Hasanain et al. (2024) and Ramponi et al. (2025) who apply similar approach for propaganda detection in Arabic and Italian, respectively.
Automated Metrics.
Although human evaluation is considered the gold standard for assessing quality in NLG tasks, its substantial demands in terms of time and cost make it less scalable and practical. This has led to an overreliance on traditional automatic metrics (Schmidtová et al., 2024), which typically assess text quality by measuring lexical overlap with human-written reference texts (e.g., Papineni et al., 2002; Lin, 2004). Although efficient and scalable, these metrics focus on surface-level properties and are unable to measure fine-grained aspects of text quality (Gehrmann et al., 2023). In addition, traditional automatic metrics have been shown to generally correlate weakly with human judgements (Novikova et al., 2017; Reiter, 2018). Model-based metrics (Rei et al., 2020; Yuan et al., 2021; Zhong et al., 2022) have addressed these issues to some extent by fine-tuning pre-trained models to evaluate text beyond lexical overlap. While these metrics generally correlate better with human judgments compared to overlap-based metrics, they still suffer from several limitations. Specifically, they provide limited explainability by providing only numeric scores, require large amounts of labeled training data, and are often limited to evaluating specific tasks or aspects (e.g., Kryscinski et al., 2020; Mehri & Eskénazi, 2020).
To automate error span annotation in data-to-text, Kasner et al. (2021) introduce a combination of a rule-based NLG system with a neural retriever model and fine-tuned RoBERTa (Liu et al., 2019). Goyal et al. (2022) propose a protocol for fine-grained annotation of coherence errors in narrative summarization, applying it to both human and automatic span annotations. In MT evaluation, xCOMET (Guerreiro et al., 2024) is a recent model-based metric specialized in error span annotation.
LLM-as-a-judge.
With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators across diverse tasks. Although their generative capabilities initially have been applied mainly to simple numeric scoring Liu et al. (2023); Sottana et al. (2023), other approaches include generating textual feedback before their final decision (Li et al., 2024; Kim et al., 2024a; b).
Relatively few works have addressed evaluation as error span annotation with LLMs. To obtain error spans for translation evaluation, GEMBA-MQM (Kocmi & Federmann, 2023) promps an LLM with annotation guidelines based on MQM evaluation protocol. InstructScore (Xu et al., 2023) and TigerScore (Jiang et al., 2024a) fine-tune smaller LLMs on synthetic data generated by GPT-4 to generate error spans. Although computationally more efficient than prompting larger LLMs, these two metrics either require reference texts, or are limited to predefined evaluation aspects. Kartáč et al. (2025) apply an ensemble of open-weight LLMs as evaluators to a variety of tasks and aspects, generating both numeric scores and error spans with detailed explanations. To address potential data contamination, Kasner & Dušek (2024) collect novel structured data and model outputs in five domains, and collect span annotations for semantic accuracy errors with both crowdworkers and GPT-4.
Goel et al. (2023) apply LLMs for medical text annotation, identifying and extracting spans containing specific named entities.
Appendix I Examples
In Tables 26, 25 and 24, we show examples of the annotated outputs for our tasks.
In Table 27, we show an example of a reasoning trace of Llama 3.3 prompted with $\mathcal{P}_{\texttt{cot}}$.
In Tables 28 and 29, we show the examples of reasoning traces of DeepSeek-R1.
Annotated by Llama 3 When the left made Linda Sarsour into its role model, it climbed into bed withfear the worst of the worstlabelling. The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundloaded to commit school shootings, prosecutors alleged in court documents Wednesday.
Annotated by DeepSeek r1 When the left made Linda Sarsour into its role model, it climbed into bed with the worst of the worst.loaded The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compound to commit school shootings, prosecutors alleged in court documents Wednesday.exag.
Annotated by o3-mini When the left made Linda Sarsour into its role model, it climbed into bed with the worst of the worst.loaded The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundloaded to commit school shootings, prosecutors alleged in court documents Wednesday.
Annotated by human When the left made Linda Sarsour into its role modellabelling, it climbed into bedloaded with the worst of the worst.exag. The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundlabelling to commit school shootings, prosecutors alleged in court documents Wednesday.
Table 24: Two examples for span annotation of Propaganda outputs with appeal-to-fearfear, name-calling-labellinglabelling, loaded-languageloaded, and exaggerationexag. as span categories.
Source Stealthily I made my way towards the back of the store, murmurs and talking heard beyond a metal grate which took a few yanks to tear off. Climbing into the vent was the safest and quietest option for me right now; climbing my way through, I’d pick up on conversations and talks between strangely dressed men. I noticed Al inside, chained. Figures he’d get caught eventually.
Annotated output by Llama3.3 Plížil jsem semajor ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemiminor, které se podařilo po několika pokusech strhnout. Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorům mezi podivně oblečenými mužiminor. Všiml jsem si Ala uvnitř, v řetězech. Je jasné, že se nakonec chytil.major
Annotated output by DeepSeek r1 Plížil jsem se ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemi, které se podařilo po několika pokusech strhnout. Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorůmminor mezi podivně oblečenými muži. Všiml jsem si Ala uvnitř, v řetězech. Je jasné, že se nakonec chytil.minor
Annotated output by o3-mini Plížil jsem se ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemiminor, které se podařilo po několika pokusech strhnout. Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorůmminor mezi podivně oblečenými muži. Všiml jsem si Ala uvnitř, v řetězech. Je jasné, že se nakonec chytil.major
Annotated output by human Plížil jsem se ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemi, které se podařilo po několika pokusech strhnout. Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorům mezi podivně oblečenými muži. Všiml jsem si Ala uvnitř, v řetězech. Je jasné, že se nakonec chytil.minor
Source “It’s your birthday soon, isn’t it?” Ivory asked, remembering that the princess’ birthday was in a few days. Kari nodded, a sad glint in her light blue eyes.
Annotated output by Llama 3 “¿Es su cumpleaños pronto, no lo es?”major Ivory le preguntó,minor recordando que el cumpleaños de la princesa era en unos días. Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
Annotated output by DeepSeek r1 “¿Es su cumpleaños pronto, no lo es?”major Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días. Kari nodó, un deslumbramientominor triste en sus ojos azules claros.
Annotated output by o3-mini “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días. Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
Annotated output by human “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días. Kari nodómajor, un deslumbramiento triste en sus ojos azules claros.
[…]
Annotated by Llama 3 Borsec will experience a period of overcast skies with light snow expected for the next few days.not check. Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.not check. There’s a chance of scattered clouds and occasional clear periods throughout the week. Winds will be moderate, ranging from 1-3 m/s.mislead. Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
Annotated by DeepSeek r1 Borsec will experience a period of overcast skiescont. with light snow expected for the next few days. Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.mislead. There’s a chance of scattered clouds and occasional clear periods throughout the week. Winds will be moderate, ranging from 1-3 m/s. Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
Annotated by o3-mini Borsec will experience a period of overcast skies with light snow expected for the next few days. Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.mislead. There’s a chance of scattered clouds and occasional clear periods throughout the week. Winds will be moderate, ranging from 1-3 m/s. Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
Annotated by human Borsec will experience a period of overcast skies with light snow expected for the next few days. Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.cont. There’s a chance of scattered clouds and occasional clear periods throughout the week. Winds will be moderate,cont. ranging from 1-3 m/s. Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
Table 26: Example for span annotation of D2T-Eval in the weather domain with contradictorycont, misleadingmislead., and not-checkeablenot check. as error span categories. In the annotation interface, the weather panel is interactive and hovering over specific time point shows the temperature and wind conditions.
Annotated by human annotator The Ituano team won 4-3 against Vila Nova in a match played on November 17th, 2024. Ituano scored four goals throughout the game, with Leozinho and Salatiel contributing to their victory.mislead. Vila Nova fought back with three goals from Luciano Naninho, Alesson, and Jemmes, but ultimately fell short.
Annotated by Llama 3 with $\mathcal{P}_{\texttt{base}}$ The Ituano team won 4-3 against Vila Nova in a match played on November 17th, 2024. Ituano scored four goals throughout the game, with Leozinho and Salatiel contributing to their victory.mislead. Vila Nova fought back with three goals from Luciano Naninho, Alesson, and Jemmesmislead., but ultimately fell short.
Annotated by Llama 3 with $\mathcal{P}_{\texttt{cot}}$ The Ituano team won 4-3 against Vila Nova in a match played on November 17th, 2024. Ituano scored four goals throughout the game, with Leozinho and Salatiel contributing to their victory. Vila Nova fought back with three goals from Luciano Naninho, Alesson, and Jemmes, but ultimately fell short.
Reasoning trace of Llama 3 with $\mathcal{P}_{\texttt{cot}}$ (emphasis added) To identify errors in the text, we need to compare it with the provided data. The data contains information about a football match between Ituano and Vila Nova, including the score, goals, and events during the game. First, let’s check if the score mentioned in the text matches the data. According to the data, the score is indeed 4-3 in favor of Ituano, so this part of the text is correct. Next, we look at the goals scored by each team. The text mentions that Ituano scored four goals and attributes contributions to their victory to Leozinho and Salatiel. From the data, we see that Ituano’s goals were scored by José Aldo, Leozinho, Salatiel, and Vinícius Paiva. While Leozinho and Salatiel are correctly mentioned as contributors, the text does not mention José Aldo and Vinícius Paiva, who also scored for Ituano. However, this omission does not necessarily make the statement incorrect or misleading since it doesn’t claim to list all scorers. The text then mentions that Vila Nova scored three goals from Luciano Naninho, Alesson, and Jemmes. According to the data, these players indeed scored for Vila Nova, with their goals being part of the events listed. Therefore, this information is accurate based on the provided data. Given the information in the text and comparing it with the data, there doesn’t seem to be any direct contradictions (Contradictory), unverifiable claims (Not checkable), misleading statements (Misleading), incoherent phrasing (Incoherent), repetitive information (Repetitive), or other issues that would require annotation. Since the text accurately reflects the data provided without introducing any errors or problematic statements, there are no annotations to make.
Table 27: Example of a reasoning trace of Llama 3.3 prompted with $\mathcal{P}_{\texttt{cot}}$ in the D2T-Eval football domain. The model prompted with $\mathcal{P}_{\texttt{cot}}$ dismisses the annotation of the sentence “with Leozinho” that the model prompted with $\mathcal{P}_{\texttt{base}}$ (and the human annotator) annotate as misleadingmislead.. The screenshot of the data input is truncated.
Annotated by DeepSeek-R1 The Posh Equal Plus X700 is a 7-inch smartphone with an IPS LCD display and a resolution of 600 x 1024 pixels. It features a 5MP main camera with LED flash and supports video recording. The device runs on Android 6.0 Marshmallow and has a Qualcomm MT8321 chipsetcont., along with 1GB RAM and 16GB internal storage. The Posh Equal Plus X700 is powered by a Li-Po battery with a capacity of 2650 mAh.
Reasoning trace of DeepSeek-R1 (truncated) Alright, I need to identify errors in the given text based on the provided data. Let me go through each sentence one by one. First sentence: "The Posh Equal Plus X700 is a 7-inch smartphone with an IPS LCD display and a resolution of 600 x 1024 pixels." Checking the data, under ’Display’, it says size is 7.0 inches and resolution 600x1024. So that’s correct. Next: "It features a 5MP main camera with LED flash and supports video recording." Data shows Main Camera has 5 MP and LED flash, Video is ’Yes’. Correct again. Then: "The device runs on Android 6.0 Marshmallow…" Wait, in the data under Platform, OS is Android 6.0 (Marshmallow), correct. But chipset is Mediatek MT8321, not Qualcomm. The text says Qualcomm MT8321, which is wrong. So that’s an error. Next part: "…and has a Qualcomm MT8321 chipset, along with 1GB RAM and 16GB internal storage." As mentioned, chipset is incorrect. Also, Memory section shows Internal as 16GB 1GB RAM, so that part is correct. Last sentence: "The Posh Equal Plus X700 is powered by a Li-Po battery with a capacity of 2650 mAh." Battery details match exactly. So no issue here. So the only error is the chipset being incorrectly stated as Qualcomm instead of Mediatek. This contradicts the data, so it’s a Contradictory error.
Table 28: The reasoning trace of DeepSeek-R1 prompted with $\mathcal{P}_{\texttt{base}}$ in D2T-Eval gsmarena domain with a contradictorycont annotation. The visualization of the input data is truncated, the relevant field is highlighted.
Input text annotated by DeepSeek-R1 (truncated) Dallas Chief Says She Can’t Fire Amber Guyger for Killing Botham Jean - Dept Rules Imply Otherwiseb-w-fall. Dallas, TX — As citizens call for justice in the case of Dallas police officer Amber Guyger — the cop who killed the innocentBotham Jean in his own home — the controversy continues to grow. In a town hall meeting this week, citizens showed that they still have a lot of questions for the Dallas Police Department in regards to the cold-bloodedkilling of Botham Jean. On top of questions as to why it took several days to arrest the cop — despite the fact that Guyger admitted to it — were questions like why this person who has proven to be a threat to societyhas yet to be fired from the police force. During the meeting, as Chief Renee Hall dodged several questions over the details of the arrest, she did go out of her way to answer the question of why Guyger is still collecting a paycheck from their department. The short of it, according to Hall, is she can’t fire her.
[…2,849 characters skipped …]
Reasoning trace of DeepSeek-R1 (truncated) Okay, so I need to analyze the given text and identify spans that use propaganda techniques. Let me go through each paragraph step by step. Starting with the title: "Dallas Chief Says She Can’t Fire Amber Guyger for Killing Botham Jean - Dept Rules Imply Otherwise." Hmm, this seems to present two opposing viewpoints without a middle ground. It’s implying either she can’t fire her or the dept rules say otherwise, which might be a Black-and-White Fallacy. In the first paragraph, it mentions citizens calling for justice and the controversy growing. Not sure about propaganda here yet. The second paragraph talks about a town hall meeting with questions about the arrest delay and why Guyger hasn’t been fired. Still more informative than propagandistic. Moving on, Chief Hall dodges questions but explains she can’t fire Guyger due to laws. Then experts say this isn’t true. This sets up a conflict where the authority (Chief) is contradicted by other experts, which might relate to Appeal_to_Authority if the article is using the experts’ opinions to discredit the Chief without evidence.
[…1,385 characters skipped …]
So, I think several propaganda techniques are present: Appeal_to_Authority (using experts), Black-and-White Fallacy in the title, Exaggeration/Minimisation with other officers’ cases, Doubt when questioning Hall’s actions, and Loaded_Language throughout.
Table 29: Excerpt from the reasoning trace of DeepSeek-R1 prompted with $\mathcal{P}_{\texttt{base}}$ in Propaganda. Both the input text (containing the annotated Black-and-White Fallacyb-w-fall.) and the reasoning trace are truncated.

### v2

@@
 Abstract
-For high-quality texts, single-score metrics seldom provide actionable feedback.
-In contrast, span annotation—pointing out issues in the text by annotating their spans—can guide improvements and provide insights.
-Until recently, span annotation was limited to human annotators or fine-tuned encoder models.
-In this study, we automate span annotation with large language models (LLMs).
-We compare expert or skilled crowdworker annotators with open and proprietary LLMs on three tasks: data-to-text generation evaluation, machine translation evaluation, and propaganda detection in human-written texts.
-In our experiments, we show that LLMs as span annotators are straightforward to implement and notably more cost-efficient than human annotators.
-The LLMs achieve moderate agreement with skilled human annotators, in some scenarios comparable to the average agreement among the annotators themselves.
-Qualitative analysis shows that reasoning models outperform their instruction-tuned counterparts and provide more valid explanations for annotations.
+Span annotation is the task of localizing and classifying text spans according to custom guidelines.
+Annotated spans can be used to analyze and evaluate high-quality texts for which single-score metrics fail to provide actionable feedback.
+Until recently, span annotation was limited to human annotators or fine-tuned models.
+In this study, we show that large language models (LLMs) can serve as flexible and cost-effective span annotation backbones.
+To demonstrate their utility, we compare LLMs to skilled human annotators on three diverse span annotation tasks: evaluating data-to-text generation, identifying translation errors, and detecting propaganda techniques.
+We demonstrate that LLMs achieve inter-annotator agreement (IAA) comparable to human annotators at a fraction of a cost per output annotation.
+We also manually analyze model outputs, finding that LLMs make errors at a similar rate to human annotators.
 We release the dataset of more than 40k model and human annotations for further research.
@@
 Fine-grained aspects of texts, such as faithfulness or coherence, depend on local lexical choices.
-We need a detailed analysis in order to reflect them in the evaluation of natural language generation (NLG) systems or quality judgments on human-written texts.
-However, most automatic evaluation metrics are based on direct assessment, producing holistic scores for each evaluated aspect Gkatzia & Mahamood (2015); Sai et al.
+In order to reflect these aspects in quality judgments of texts, we need analysis techniques that provide the appropriate amount of detail.
+However, most automatic evaluation metrics for natural language generation (NLG) are based on direct assessment, producing holistic scores for each evaluated aspect Gkatzia & Mahamood (2015); Sai et al.
 (2023); Schmidtová et al.
 (2024).
-Although numerical values make it easy to rank systems, these metrics are too simplistic and susceptible to biases or miscalibration issues of the underlying models Gehrmann et al.
+Although numerical values make it easy to rank systems, these metrics are too simplistic and susceptible to biases or miscalibration of the underlying models Gehrmann et al.
 (2023); Gao et al.
@@
 (2024).
-The subject of our study, span annotation, offers an alternative approach.
-Figure 1 shows the difference between direct assessment and span annotation.
-Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans and classify them into predefined categories.
+The subject of our study, span annotation (Figure 1), offers an alternative approach.
+Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans and classify them according to task-specific guidelines.
 In contrast to numerical rating, annotations are aligned to the evaluated text, which makes them more explainable and actionable.
-The annotations can also be examined post hoc, enabling a more modular evaluation process.
+The annotations can also be examined post hoc, which brings more accountability into the evaluation process.
 Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation.
-The method has traditionally relied on human annotators, making it costly and difficult to scale Da San Martino et al.
+The method traditionally required human annotators, making it costly and difficult to scale Da San Martino et al.
 (2019); Thomson & Reiter (2020); Popovic (2020); Kocmi et al.
-(2024b).
-The situation has changed with LLM-as-a-judge paradigm, in which LLMs follow instructions for the evaluation process similarly to human annotators Zheng et al.
+(2024c).
+The situation may change with the upcoming LLM-as-a-judge paradigm, in which LLMs follow instructions for the evaluation process similarly to human annotators.
+However, current LLM-as-a-judge applications still focus mainly on document-level classification or numerical ratings on the Likert scale Zheng et al.
 (2023); Gu et al.
 (2024).
-However, the current LLM-as-a-judge applications still mostly focus on document-level classification or numerical ratings on the Likert scale, which are too coarse and unreliable for evaluating fine-grained aspects of texts.
-Here, we investigate a specific LLM-as-a-judge application: using LLMs as span annotators.
-Instructing LLMs to annotate text spans with predefined categories is a way to produce span annotations automatically and robustly, lowering the barrier to using the method for NLG evaluation.
-We focus on the following research questions:
-  • RQ1: How to instantiate an LLM-based span annotator? We do a small-scale study of prompting techniques to arrive at a setup that is robust across diverse tasks and scenarios.
-  • RQ2: How do LLMs compare to skilled human annotators?
-We compute inter-annotator agreement (IAA) between annotations produced by LLMs and skilled human annotators on both novel data and existing datasets.
-  • RQ3: How valid are LLM annotations and their explanations? We manually analyze a subset of LLM and human annotations to assess the validity of annotations and their respective explanations.
-For our experiments, we select three diverse tasks (cf.
-Section 3.1): evaluation of data-to-text generation Thomson & Reiter (2020), evaluation of machine translation Kocmi et al.
-(2024b), and propaganda detection Da San Martino et al.
+In this work, we investigate whether LLMs can be applied as a substitute for human annotators on span annotation tasks.
+First, we focus on how to produce span annotations with LLMs.
+We discuss input tagging methods and experiment with prompting techniques, looking for a setup that is robust and widely applicable.
+Next, we compare LLMs with skilled human annotators in terms of IAA.
+Where possible, we contrast the IAA between LLMs and human annotators with the upper bound of the IAA between the annotators themselves.
+Finally, we manually analyze a subset of LLM and human annotations to assess the validity of the annotations and their respective explanations.
+For our experiments, we select three diverse span annotation tasks (cf.
+Section 3.1): evaluating data-to-text generation Thomson & Reiter (2020), identifying errors in machine translation Kocmi et al.
+(2024a), and detecting propaganda techniques in human-written texts Da San Martino et al.
 (2019).
-We base our span annotators on both open and proprietary state-of-the-art LLMs, including recent reasoning models.
-We collect more than 40k model and human annotations that we release for further investigation.
-In the experimental part, we first analyze various prompting styles, demonstrating that zero-shot prompting with detailed guidelines is a robust way to set up LLMs as span annotators.
-Using automatic metrics, we show that LLMs are close to IAA of skilled human annotators, leveling or surpassing the performance of qualified crowdworkers in some scenarios.
-Qualitatively, we find that LLM annotations are completely correct in 49.5% of cases (56.5% in the case of reasoning models), which is similar to qualified crowdworkers.
-Figure 1: Comparison between direct assesment of text (=conventional evaluation methods) and span annotation with pre-defined categories (=this paper).
-Span annotation provides more signal for diagnosing model outputs.
+Note that unlike established span annotation tasks such as named entity recognition Tjong Kim Sang & De Meulder (2003); Wang et al.
+(2025a), there are no general purpose fine-tuned models for these tasks since annotation guidelines are specified ad hoc, making their automation difficult.
+We base our span annotators on open and proprietary state-of-the-art LLMs, including recent reasoning LLMs that leverage test-time compute for better performance on challenging tasks Muennighoff et al.
+(2025); Marjanović et al.
+(2025).
+Our contributions are as follows:
+  1. 1.
+    We demonstrate that zero-shot prompting in combination with structured outputs can be a robust way to set up LLMs as span annotators, given that the model is provided detailed annotation guidelines (Sections 3.2 and 5.1)
+  2. 2.
+    Using a combination of specialized IAA metrics and manual analysis, we show that performance of LLMs is close to skilled human annotators, in some scenarios leveling or surpassing the performance of qualified crowdworkers (Sections 5.2 and 5.3).
+  3. 3.
+    We release a dataset of more than 40k human and model annotations for all our tasks and setups, including reasoning traces from reasoning LLMs and annotations collected from skilled crowdworkers.
+Figure 1: Examples of span annotation tasks that we automate with LLMs. We unify the setup for evaluation tasks (D2T-Eval, MT-Eval) and text analysis tasks (Propaganda).
 2 Related Work
 LLMs for NLG Evaluation.
-Automatic NLG metrics tradionally assess text quality by measuring similarity to human-written reference texts Sai et al.
+Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts Sai et al.
 (2023); Schmidtová et al.
 (2024).
-As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgements Novikova et al.
+As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgments Novikova et al.
 (2017); Reiter (2018).
@@
 Span Annotation Protocol.
-In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024b), where annotators mark erroneous spans in translations.
-In data-to-text (D2T) generation, span annotation was applied by Thomson & Reiter (2020), who introduce a span-annotation-based evaluation protocol for annotation of generated basketball reports.
+In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024c), where annotators mark erroneous spans in translations.
+In data-to-text (D2T) generation, span annotation is applied by Thomson & Reiter (2020), who introduce a span-based evaluation protocol for annotation of generated basketball reports.
 Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection Da San Martino et al.
-(2019), and text summarization Subbiah et al.
+(2019) and text summarization Subbiah et al.
 (2024).
-Unlike our work, these works conduct only human annotation.
+Note that unlike our work, these works perform span annotation only using human annotators.
 Automating Span Annotation.
@@
 Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities.
-While existing works (Kocmi & Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner & Dušek, 2024; Chang et al., 2024; Kartáč et al., 2025; Ramponi et al., 2025; Zouhar et al., 2024) applied LLMs as a tool for evaluating their systems, our study systematically compares LLMs to human annotators across tasks and domains.
-See more discussion on related work in Appendix H.
+We build on recent work that applies LLMs as a task-specific evaluation tool (Kocmi & Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner & Dušek, 2024; Chang et al., 2024; Zouhar et al., 2025; Kartác et al., 2025; Ramponi et al., 2025).
+Our work is the first that unifies diverse span annotation tasks within a single framework and systematically compares the performance of LLMs to human annotators.
+See more discussion on related work in Appendix F.
 3 Automating Span Annotation with LLMs
@@
 3.1 Span Annotation: Task Definition
-The goal of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$, i.e., to produce a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i}\rangle$:
-  • $s_{i},e_{i}\in\{1,\ldots,n\},s_{i}<e_{i}$ are the start and end indices of the annotated span, [^2] Our definition allows overlapping annotations: $s_{i}\leq s_{j}\leq e_{i};i\neq j$.
-  • $c_{i}\in C$ is the assigned annotation category from the pre-defined set of task-specific categories $C=\{c_{1},\ldots,c_{k}\}$.
-The process is further specified in the guidelines $\mathcal{G}$: a set of fine-grained instructions for the annotation process regarding how to handle ambiguities or how to decide which parts of the text to include in the span.
-The annotator also considers the source $X$ (e.g., the translation source or the input data).
-Note that $X$ can be empty if we are annotating only intrinsic text aspects such as coherence or style.
-3.2 Automating Span Annotations
-Following our definition from Section 3.1, our goal is to collect the set of annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$.
-In our setup, we rely on the LLM-as-judge paradigm: prompting an LLM to follow annotation guidelines:
+The goal of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$ given:
+  • the set of categories $C=\{c_{1},\ldots,c_{k}\}$,
+  • the annotation guidelines $\mathcal{G}$ (including category definitions and other instructions),
+  • the source $X$ (such as the translation source; empty if we are annotating only intrinsic text aspects).
+The goal is to produce a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i},r_{i}\rangle$:
+  • $s_{i},e_{i}\in\{1,\ldots,n\},s_{i}<e_{i}$ are the start and end indices of the annotated span,
+  • $c_{i}\in C$ is the assigned annotation category,
+  • $r_{i}$ is an explanation for the annotation (optional). [^2] We collect the explanations from LLMs, but not from human annotators due to time considerations.
+3.2 Automating Span Annotation with LLMs
+To automate span annotation, we collect the annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$ from an LLM:
 $A=\textsc{LLM}(\text{prompt}(Y,C,\mathcal{G},X)).$ (1)
-To obtain the annotations, we closely follow the setup of Kasner & Dušek (2024).
-Specifically, we use structured decoding, asking for the list of annotations in JSON format.
-The output contains the fields reason (a short sentence justifying the annotation), text (the literal content of the matched span), and type (the integer index of the error category as given in the prompt).
-This approach enables us to robustly parse the model output and collect auxiliary explanations for annotations.
-[^3] See Appendix B for a detailed discussion of the selected approach and alternatives.
+How to Identify the Spans?
+An annotation needs to contain the start and end indices of the span.
+For encoder models, the task would be trivially solved by classifying tokens on top of the encoder’s hidden states.
+However, to the best of our knowledge, there is no established way of tagging the input text with decoder-only models.
+Of the methods from previous work, we discarded surrounding spans with special XML-like tokens Wang et al.
+(2025b); Treviso et al.
+(2024).
+This method requires the LLM to generate full copy of the input text and does not offer a fallback if the model fails to do so.
+We also rule out listing start and end indices of spans Hasanain et al.
+(2024); Ramponi et al.
+(2025), as Hasanain et al.
+note that the model frequently generates indices that do not match the corresponding spans in a paragraph.
+Instead, we adopted the method of listing textual content of the spans Kocmi & Federmann (2023); Zouhar et al.
+(2025); Kasner & Dušek (2024).
+Here, the start and end indices of the span are found heuristically by string matching of the span’s textual content.
+With this approach, we do not need to modify the input and failing to match a single annotation does not influence other matches.
+[^3] We do note that this approach can misalign the model output by assigning the annotation to the first match, but our preliminary experiments showed that the collisions are rare.
+Parsing Model Output
+To obtain the annotations, we request the list of annotations in JSON format, following the setup of Kasner & Dušek (2024).
+This setup allows us to use constrained decoding with a defined JSON scheme, ensuring that the output is syntactically valid.
+We require each annotation to contain the fields reason (the explanation $r_{i}$), text (the textual content of the span), and type (the integer index of the error category $c_{i}$).
+[^4] Given the considerations of Castillo (2024), we ensure that the reason field has precedence.
 3.3 Evaluating Span Annotations
-To determine and compare the quality of the human and the automatic annotation process, we need a notion of similarity between two sets of annotations $\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $Y=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$.
-Based on these similarities, we can then compute reliability metrics, such as IAA or comparison to gold annotations.
-We consider three similarity metrics.
+To compare annotations automatically, we need a notion of similarity between two sets of annotations $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\hat{\mathcal{A}}=\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $\mathcal{Y}=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$.
+We consider three similarity metrics:
 Pearson correlation $\rho$ over counts.
-The simplest correlation computation simply compares how many spans were annotated for each example.
-$\mathrm{CountCorrelation}(A_{*},\hat{A}_{*})=\mathrm{Pearson}(\langle|A_{y}|,|\hat{A}_{y}|\rangle_{y\in Y})$ (2)
+This metric simply compares how many spans were annotated for each example:
+$\mathrm{CountCorrelation}(\mathcal{A},\hat{\mathcal{A}})=\mathrm{Pearson}(\langle|A_{Y}|,|\hat{A}_{Y}|\rangle_{Y\in\mathcal{Y}})$ (2)
+This metric serves as a sanity check: a low value would suggest that an annotator either skips examples or over-annotates, indicating unclear annotation guidelines.
 Precision, Recall, and $\mathbf{F_{1}}$.
-To also quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined in Da San Martino et al.
+To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined in Da San Martino et al.
 (2019).
-These measures are on matching annotations, adjusted to give partial credit to imperfect matches (which is necessary given the varying lengths of annotations): $\text{Precision}(A_{*},\hat{A}_{*})=\frac{1}{|A|}\sum_{a\in A}\max_{\hat{a}\in\hat{A}}\frac{|a\cap\hat{a}|}{|a|},$ (3) $\text{Recall}(A_{*},\hat{A}_{*})=\frac{1}{|\hat{A}|}\sum_{\hat{a}\in\hat{A}}\max_{a\in A}\frac{|a\cap\hat{a}|}{|\hat{a}|},$ (4) where $a\cap\hat{a}$ is the character overlap between two annotation spans and $|a|=e-s+1$ is the length of the annotation span in characters (see Section 3.1).
-Subsequently, we compute $F_{1}$-score as the harmonic mean of precision and recall.
+These measures are on matching annotations, adjusted to give partial credit to imperfect matches: $\text{Precision}(A_{Y},\hat{A}_{Y})=\frac{1}{|A_{Y}|}\sum_{a\in A_{Y}}\frac{|a\cap\hat{a}|}{|a|},$ (3) $\text{Recall}(A_{Y},\hat{A}_{Y})=\frac{1}{|\hat{A}_{Y}|}\sum_{\hat{a}\in\hat{A}_{Y}}\frac{|a\cap\hat{a}|}{|\hat{a}|},$ (4) where $a\cap\hat{a}$ is the character overlap between two annotation spans and $|a|=e-s+1$ is the length of the annotation span in characters (see Section 3.1).
+Subsequently, we compute the $F_{1}$-score as the harmonic mean of precision and recall.
 For each of the metrics, we consider soft and hard variants.
-While the soft variant disregards span categories, the hard one only considers overlaps where the span category is matching.
+The hard variant only considers overlaps where the span category matches, while the soft variant disregards the categories.
 We consider the hard variant to be the default.
-Additionally, we report the difference $F_{1}\Delta=F_{1}(\text{soft})-F_{1}(\text{hard})$.
+In addition, we report the difference $F_{1}\Delta=F_{1}(\text{soft})-F_{1}(\text{hard})$.
 Gamma $\bm{\gamma}$.
-The $F_{1}$ score is sensitive to varying span granularities and does not consider near matches with no overlap.
-It also does not account for agreement by chance.
+The $F_{1}$ score is sensitive to varying span granularities and does not consider near-matches with no overlap or agreement by chance.
 To this end, we follow Da San Martino et al.
 (2019) and Hasanain et al.
-(2024) in using the $\gamma$ score (Mathet et al., 2015) as another annotation similarity metric.
-This metric operates on whole sets of annotations for a single input $y$: $A_{*,y}=\{A,\hat{A}_{y},\ldots,A_{k,y}\}$.
-The metric builds the best possible alignment between the annotations $A_{i,y}$ and $A_{j,y}$ and computes their dissimilarity $\mathrm{Dissimilarity}(A_{i,y},A_{j,y})$.
-This is then compared to the expected dissimilarity by chance, which is created by sampling across all annotations: $\mathrm{DissimilarityChance}(A_{*,y})$.
-The final score is based on the average dissimilarity between annotations and the dissimilarity by chance:
-$\gamma(A_{*,y})=1-\frac{\frac{1}{k^{2}}\sum_{i,j=1}^{k}\mathrm{Dissimilarity}(A_{i,y},A_{j,y})}{\mathrm{DissimilarityChance}(A_{*,y})}$ (5)
-The score ranges from negative infinity to 1, where 1 is achieved when the annotations have perfect alignment.
+(2024) in using the $\gamma$ score (Mathet et al., 2015) as a complementary metric.
+In short, the metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations.
+The score ranges from negative infinity to 1, where 1 is achieved when the annotations are perfectly aligned.
 The $\gamma$ score extends Krippendorff’s $\alpha$ (Krippendorff, 1980), another popular metric, by computing the category-aware span alignments.
 We use the implementation of Titeux & Riad (2021).
-Examples with empty annotation sets.
-For an output $Y$, either one or both of the annotation sets $A_{y},\hat{A}_{y}$ can be empty.
-Note that this case is natural and may even be desirable: e.g., if the goal is to annotate errors in an output that is entirely correct.
-However, neither of the introduced scores accounts for these cases accordingly.
+$S_{\emptyset}$ score.
+For an output $y$, one or both annotation sets $A_{Y},\hat{A}_{Y}$ can be empty.
+Having these cases is desirable: e.g.
+if the goal is to annotate errors in an output that is entirely correct.
+However, neither of the introduced scores accounts for these cases.
 [^5] The $\gamma$ score is undefined for less than two non-empty annotation sets: these examples thus need to be skipped during the $\gamma$ computation.
 The F1 score does not account for examples with no annotations.
-To compensate for that, we introduce a score $S_{\emptyset}$ that is computed for examples with less than two non-empty sets of annotations as follows:
-$S_{\emptyset}=1/(1+|A|)$ (6)
-where $|A|$ is the number of annotations from the remaining annotator.
-The score is equal to 1 for the cases where no annotator produced any annotations, i.e., a perfect match, and decreases proportionally to the number of annotations from the remaining annotator.
+To compensate for that, we introduce a score $S_{\emptyset}$ that is computed for examples with less than two nonempty sets of annotations as $S_{\emptyset}=1/(1+|A|)$, where:
+$|A|=\begin{cases}|A_{Y}|&\text{if }|\hat{A}_{Y}|=0,\\ |\hat{A}_{Y}|&\text{otherwise}.\end{cases}$ (5)
+The score is equal to 1 for the cases where no annotator produced any annotation (i.e., a perfect match) and decreases proportionally to the number of annotations from the annotator that provided a non-zero number of annotations.
 4 Experiments
-Task # Cat. # Texts Avg. Len Task Setup Novel Data
-D2T-Eval 6 1296 118/715 Kasner & Dušek (2024)
-MT-Eval 2 2,854 26/185 Kocmi et al. (2024a) ×
-Propaganda 18 100 914/4,659 Da San Martino et al. (2019) ×
+Task # Cat. # Texts Avg. Len Novel Data
+D2T-Eval 6 1,296 118/715
+MT-Eval 2 2,854 26/185
+Propaganda 18 100 914/4,659
 Table 1: Overview of span annotation tasks used in our experiments.
 # Cat.
-denotes the number of categories used in the task (see Appendix E for their listings), # Texts the number of texts annotated, and Avg.
+denotes the number of categories used in the task (see Appendix C for their listings), # Texts the number of texts annotated, and Avg.
 Len the average number of words/characters in the output.
-We cover three span annotation tasks of different quantitative and qualitative aspects: evaluation of data-to-text generation (D2T-Eval; Section 4.1), evaluation of machine translation (MT-Eval; Section 4.2), and propaganda detection (Propaganda; Section 4.3).
+We cover three span annotation tasks of different quantitative and qualitative aspects: evaluating data-to-text generation (D2T-Eval; Section 4.1), identifying errors in machine translation (MT-Eval; Section 4.2), and detecting propaganda techniques (Propaganda; Section 4.3).
 In Table 1, we provide a quantitative overview of our tasks.
-[^5] See also Figure 1 and Appendix I for the visualization of examples of our tasks.
-Additionally, we discuss how we prevent the issue of data contamination for our tasks in Appendix C.
-We further describe our process of collecting annotations in Section 4.4.
+[^6] See also Figure 1 and Appendix G for the visualization of examples of our tasks.
 4.1 Task: Evaluation of Data-to-text Generation
-In D2T-Eval, we evaluate semantic accuracy and stylistic aspects of texts generated from structured data Celikyilmaz et al.
+In D2T-Eval, we focus on evaluating the semantic accuracy and stylistic aspects of texts generated from structured data Celikyilmaz et al.
 (2020); Sharma et al.
 (2022).
-We apply span annotation to the text $Y$ generated from the input structured data $X$.
+The input is the text $Y$ generated from the structured data $X$.
 Subtasks and input data.
-We select three subtasks of D2T generation: generating five-day weather forecasts (openweather), generating product descriptions (gsmarena), and generating reports of soccer games (football).
-We use the approach of Kasner & Dušek (2024) to download novel structured data from publicly available APIs, to make sure they are unseen for all LLMs (cf.
-Appendix C on data contamination).
-We provide more details on the data collection process in Section D.1.
+We select three D2T generation subtasks: generating five-day weather forecasts (openweather), generating product descriptions (gsmarena), and generating reports of soccer games (football).
+We use the approach of Kasner & Dušek (2024) to download novel structured data from publicly available APIs to make sure that they are unseen for all LLMs (cf.
+Section 4.4 on data contamination).
+We provide more details on the data collection process in Section B.1.
 Collecting model outputs.
 We generate outputs for the structured inputs using LLMs.
-Specifically, we select two state-of-the-art models, Llama 3.3 70B Grattafiori et al.
+Specifically, we selected two state-of-the-art models, Llama 3.3 70B Grattafiori et al.
 (2024) and GPT-4o Hurst et al.
@@
 We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences.
-[^6] See Sections A.2 and A.3 for details on the models and Appendix F for the prompts.
+[^7] See Sections A.1 and A.2 for details on the models and Appendix D for the prompts.
 Data splits.
 For further experiments, we divide the data into multiple subsets:
-  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – the subset for LLM evaluation, annotated with crowdworkers.
-  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – the subset for the study of prompt variants, annotated internally.
+  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – the subset for LLM evaluation, annotated with crowdworkers,
+  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – the subset for the study of prompt variants, annotated internally,
   • $\mathcal{D}_{\texttt{iaa}}$ (12 outputs) – the control subset for human crowdworkers, annotated internally.
@@
 We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com.
-We apply best practices for gathering human annotations, including an iterative process of refining annotation guidelines and pre-selecting best-performing annotators using a qualification task Tseng et al.
+We apply best practices for gathering human annotations, including an iterative process of refining annotation guidelines and preselecting best-performing annotators using a qualification task Tseng et al.
 (2020); Iskender et al.
@@
 Of the 230 annotators that participated in the qualification task, we selected 50 best annotators to annotate our data.
-We provide more details on the process of collecting human annotations in Section D.2.
-4.2 Task: Evaluation of Machine Translation
-For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024a).
-The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024b) using professional translators.
+We provide more details on the process of collecting human annotations in Section B.2.
+4.2 Task: Identifying Errors in Machine Translation
+For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024b).
+The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024c) by professional translators.
 We focus on character-level span annotations of major and minor translation errors provided in the dataset.
-The inputs $X$ are the source texts used to generate the translation $Y$ [^7] Note the annotations in MT-Eval cannot overlap and need not to be aligned with word boundaries.
+The inputs $X$ in MT-Eval are the source texts used to generate the translation $Y$ [^8] Note unlike for the other tasks, the annotations in MT-Eval cannot overlap and need not to be aligned with word boundaries.
 We use data for the three textual domains present in the WMT 2024 shared task: news, literary, and social.
@@
 The original dataset has nearly 50k model outputs, making it too costly to run the full evaluation campaign.
-Therefore, we use all available system outputs for 10 sampled inputs for each language, giving us 2,854 outputs which we use for the experiments.
+Therefore, we used all available system outputs for 10 sampled inputs for each language, giving us 2,854 outputs that we use for the experiments.
 4.3 Task: Propaganda Technique Detection
@@
 The dataset contains news collected from mostly propagandistic on-line sources.
-The token-level annotations in the dataset were created by expert annotators and cover 18 categories of logical fallacies and persuasion techniques.
+The token-level annotations in the dataset created by expert annotators cover 18 categories of logical fallacies and persuasion techniques.
 We use the test split for our experiments.
 Note that there are no inputs $X$ for the task, as all the annotated categories are intrinsic to the evaluated text $Y$.
+4.4 Mitigating Data Contamination
+We took preventive measures to ensure that our results are not affected by data contamination Balloccu et al. (2024); Dong et al. (2024); Jiang et al. (2024b):
+  • For D2T-Eval, we follow the approach of Kasner & Dušek (2024) to download a new set of structured data from publicly available sources and generate new outputs.
+  • For MT-Eval, we use the WMT 2024 dataset that was released at the end of 2024, which is after the knowledge cut-off date of the LLMs we are using.
+  • For Propaganda, the dataset is publicly available since 2019, but due to its format cannot be easily memorized.
+[^9] The model would need to separately match the annotations for numerical token ids with the text.
 4.5 Collecting LLM annotations
@@
 We also ablate $\mathcal{P}_{\texttt{base}}$ by removing extended guidelines ($\mathcal{P}_{\texttt{noguide}}$) and not asking for explanations ($\mathcal{P}_{\texttt{noreason}}$).
-The full prompts can be found in Appendix F.
+The full prompts can be found in Appendix D.
 Annotation Tool
-We use factgenie Kasner et al. (2024), a tool which supports both collecting span annotations from humans via a web interface and from LLMs via API calls.
+We use factgenie Kasner et al. (2024), a tool that supports collecting span annotations from humans via a web interface and from LLMs via API calls.
 5 Results
 In this section, we describe the results of our experiments on span annotation with LLMs.
-We evaluated the LLM annotations both automatically using the metrics from Section 3.3 (Section 5.2) and manually (Section 5.3).
+We first investigate the effect of prompting techniques in Section 5.1.
+Next, we evaluate the LLM annotations using automatic metrics (Section 5.2) and manually analyze the errors in the model outputs (Section 5.3).
 5.1 Prompting Techniques
-To answer RQ1, we study the differences between prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ split using the open models (Llama 3.3 and DeepSeek-R1). The results are shown in Table 2.
+We perform a small-scale study on $\mathcal{D}_{\texttt{dev}}$ using open models (Llama 3.3 and DeepSeek-R1) to study the differences between prompting techniques. The results are shown in Table 2.
 Llama 3.3 DeepSeek-R1
 Prompt $F_{1}$ $\gamma$a/o $F_{1}$ $\gamma$a/o
-$\mathcal{P}_{\texttt{base}}$ 0.16 0.13 2.4 0.23 0.20 1.0
-$\mathcal{P}_{\texttt{noguide}}$ 0.09 0.08 3.4 0.19 0.16 1.6
-$\mathcal{P}_{\texttt{noreason}}$ 0.19 0.13 2.2 0.22 0.18 1.1
-$\mathcal{P}_{\texttt{cot}}$ 0.08 0.10 0.8 0.21 0.19 1.1
-$\mathcal{P}_{\texttt{5shot}}$ 0.23 0.18 2.5 0.20 0.16 1.4
-Table 2: Comparison of different prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ split with Llama 3.3 and DeepSeek-R1: $F_{1}$ score, $\gamma$ score, and the average number of annotations per output (#a/o).
+$\mathcal{P}_{\texttt{base}}$ 0.20 0.13 2.4 0.25 0.20 1.0
+$\mathcal{P}_{\texttt{noguide}}$ 0.11 0.08 3.4 0.20 0.16 1.6
+$\mathcal{P}_{\texttt{noreason}}$ 0.22 0.13 2.2 0.24 0.18 1.1
+$\mathcal{P}_{\texttt{cot}}$ 0.09 0.10 0.8 0.24 0.19 1.1
+$\mathcal{P}_{\texttt{5shot}}$ 0.25 0.18 2.5 0.21 0.16 1.4
+Table 2: Comparison of prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ (#a/o is the average number of annotations per output).
 Including detailed guidelines seems beneficial: omitting the guidelines ($\mathcal{P}_{\texttt{noguide}}$) lowers the performance of both models.
-In contrast, not letting the model explain the annotation ($\mathcal{P}_{\texttt{noreason}}$) does not have a substantial effect on the model performance.
-For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce less annotations per example than the base variant (0.76 vs 2.42), leading to lower F1 and $\gamma$ scores.
-[^8] The model with $\mathcal{P}_{\texttt{cot}}$ often “overthinks” the annotations, deciding not to annotate cases of errors against which it can find some arguments.
+In contrast, not letting the model explain the annotation ($\mathcal{P}_{\texttt{noreason}}$) does not have a substantial effect.
+For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce less annotations per example than the base variant (0.8 vs.
+2.4), leading to lower F1 and $\gamma$ scores.
+[^10] The model with $\mathcal{P}_{\texttt{cot}}$ tends to “overthink” the annotations, deciding not to annotate cases of errors against which it can find some arguments.
 We provide an example of this behavior in Table 27.
-Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.999This observation is aligned with DeepSeek-AI et al.
+Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.111111This observation is aligned with DeepSeek-AI et al.
 (2025), who note that few-shot prompting consistently degrades the performance of DeepSeek-R1.
-Following this preliminary study, we decided to focus on prompting with $\mathcal{P}_{\texttt{base}}$ for our main experiments.
+Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for our main experiments.
 5.2 LLM vs. Human Annotations
-Figure 2: Comparison of LLMs and human annotators using $\mathcal{P}_{\texttt{base}}$.
+Figure 2: Comparison of LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators.
 The stripe pattern denotes proprietary models, darker shades denote reasoning models.
 For $F_{1}$ score, the grey bar denotes its soft variant.
-The dashed horizontal lines are the average value of agreement between human annotators.
+The dashed horizontal lines for D2T-Eval denote agreement between our human annotators.
 Detailed results are included in Tables 8, 16 and 20.
-To answer RQ2, we compare LLM annotations to human annotations using metrics described in Section 3.3.
-Wherever possible, we compare model results with an average IAA of human annotators.
-[^10] For D2T-Eval and MT-Eval, we estimate the bounds from the examples annotated by pairs of annotators.
-For Propaganda, we use the $\gamma$ score reported by Da San Martino et al.
-(2019) for agreement of the annotators before consolidation ($\gamma=0.31$).
-The overall results for all the tasks are outlined in Figure 2.
-We provide detailed results for individual tasks in Appendix G.
-Figure 3: Confusion matrix of category counts averaged across models for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other, cf. Table 5).
+Next, we compute IAA between LLM and human annotations using the metrics described in Section 3.3.
+The overall results for all tasks are given in Figure 2.
+We also provide detailed results for individual tasks in Appendix E.
+Figure 3: Confusion matrix for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other, cf. Table 5).
 Model football gsmarena openweath.
 $F_{1}$ $\gamma$ $F_{1}$ $\gamma$ $F_{1}$ $\gamma$
-Llama 3.3 0.27 0.20 0.08 0.08 0.10 0.05
-GPT-4o 0.27 0.19 0.07 0.06 0.16 0.12
-Claude 3.7 0.36 0.25 0.13 0.14 0.27 0.19
-DeepS. R1 0.30 0.28 0.09 0.16 0.16 0.12
-o3-mini 0.38 0.33 0.11 0.19 0.29 0.26
-Gem. 2-FT 0.35 0.31 0.13 0.16 0.22 0.16
+Llama 3.3 0.29 0.20 0.11 0.08 0.12 0.05
+GPT-4o 0.30 0.19 0.09 0.06 0.18 0.12
+Claude 3.7 0.38 0.25 0.16 0.14 0.27 0.28
+DeepS. R1 0.34 0.28 0.12 0.16 0.18 0.12
+o3-mini 0.42 0.33 0.15 0.19 0.29 0.31
+Gem. 2-FT 0.38 0.31 0.16 0.16 0.22 0.24
 Table 3: Performance of models by domain (football, gsmarena, openweather) in D2T-Eval: $F_{1}$ score (hard), $\gamma$ score. The best results are bolded.
-Performance is task-dependent.
-LLMs provide annotations that are aligned with human annotations to a degree, with exact agreement depending on the task.
-For D2T-Eval, all models surpass the average $F_{1}$-scores (both soft and hard) between human annotators (0.13 and 0.21, respectively), although only o3-mini surpasses the $\gamma$ score (0.25).
-Performance is notably lower for Propaganda, where all models remain below the $\gamma$ agreement of human annotators.
-It should be noted that expert annotators were used in this task, which can make it more difficult for the models to reach their agreement level.
-This task also has the largest number of categories, manifested by a large increase in the soft $F_{1}$-score as opposed to its hard variant.
-On average, the models are also below human performance for MT-Eval, with Claude 3.7 showing the best results (0.22 $F_{1}$-score vs.
-0.25 of human annotators).
 Reasoning models outperform instruction-tuned models.
-DeepSeek-R1, which is a reasoning model, outperforms the same-sized instruction-tuned Llama 3.3.
-The superiority of DeepSeek-R1 is most pronounced on D2T-Eval ($F_{1}$-score of 0.19 vs.
-0.15, $\gamma$ score of 0.19 vs.
+DeepSeek-R1 generally outperforms the same-sized Llama 3.3.
+The superiority of DeepSeek-R1 is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs.
+0.18, $\gamma$ score of 0.19 vs.
 0.11).
-We hypothesize that this is because statements in texts generated from structured data often involve numerical reasoning.
 The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o despite o3-mini’s price per token being approximately 2x lower.
-Both findings demonstrate the potential of test-time scaling Snell et al.
-(2024); Welleck et al.
-(2024).
-A notable exception from this trend is Claude 3.7 Sonnet, which is mostly on par with o3-mini (although its price per token is approximately 3x higher).
-Models mostly confuse related categories.
-The results of category classification on D2T-Eval suggest that the models mostly tend to confuse annotation categories such as Contradictory, Not checkable and Misleading that are all related to semantic accuracy (cf.
-Figure 3 and Section G.8).
-This suggests that categorization errors may be less serious and may be related to category ambiguity or subjective understanding of category definitions.
-The results also show that the models use a less diverse distribution of categories, resorting to the Contradictory category more often than human annotators.
-LLMs are more cost-efficient than human annotators.
-Using LLMs, as opposed to human annotators, is notably different once the financial aspect is taken into account: For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance model o3-mini LLM costs \$3.60.
-We provide a more detailed cost- and time-wise analysis in Section A.1.
+A notable exception to this trend is the Claude 3.7 Sonnet, which is mostly on par with o3-mini and excels on MT-Eval.
+However, its price per token is approximately 3x higher (cf.
+Table 4 for model costs).
+LLMs reach human IAA on D2T-Eval, Propaganda is harder.
+For D2T-Eval, we can compare model results with an average IAA on examples annotated by two human annotators (cf.
+Section B.2).
+Here, o3-mini, Claude 3.7 and Gemini 2.0 mostly reach or surpass human agreement.
+For Propaganda, the upper bound of IAA is the result of Da San Martino et al.
+(2019), who report $\gamma=0.31$ for annotators before consolidation.
+This score is substantially higher (the best LLM score being $\gamma=0.16$ for the Claude 3.7 Sonnet).
+However, this task used expert annotators and has the largest number of categories.
+The latter manifests itself in the large difference between the soft and hard $F_{1}$ scores.
+[^12] The WMT24 dataset for MT-Eval contains examples annotated with a pair of annotators, but these examples form only a small fraction and exhibit high variance between language pairs.
+Therefore, we decided to omit the comparison with human IAA in MT-Eval.
+Models confuse related categories.
+Confusion matrices (cf.
+Figure 3 and Section E.8) suggest that the models tend mainly to confuse related categories such as Contradictory, Not checkable and Misleading for D2T-Eval, which may be related to ambiguity or subjective understanding of category definitions.
+The models also use a less diverse distribution of categories, e.g., resorting to the Contradictory category more often than human annotators.
+LLMs are more efficient than human annotators.
+For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance model o3-mini LLM costs \$3.60 (cf.
+Table 4).
+The crowdworkers take 129.1 seconds per output on average, which is better than DeepSeek-R1 on our local infrastructure, but order of magnitude slower than API-based models.
+[^13] Also note that unlike LLMs, the crowdworkers are not required to produce explanations for the annotations.
+This suggests that LLMs can be an effective alternative to human annotators both in terms of costs and time.
+Metric crowdw. Llama 3.3 DeepSeek-R1 Claude 3.7 Sonnet o3-mini
+Cost per 1k outputs (\$) 500 - - 10.5 3.6
+Time per output (s) 129.1 21.6 227.5 9.0 21.8
+Table 4: Estimate of costs and time requirements on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and proprietary models (Claude 3.7 Sonnet, o3-mini).
+Note that we omit the infrastructure and energy costs for the models deployed locally.
 5.3 Manual Analysis of LLM Annotations
-To answer RQ3, we manually analyzed the quality of LLM annotations on 216 samples.
-For each model, we sampled 3 annotations per error category in D2T-Eval and 1 annotation per error category in Propaganda.
+We manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda.
+For each model, we sampled three annotations per category in D2T-Eval and one annotation per category in Propaganda.
 We show the results in Figure 4.
-Overall, 49.5% of LLM-generated annotations and 50.5% of reasons were marked as correct for both tasks (with 9.2% of annotations and 12.5% of reasons additionally marked as partially correct).
+In total, we marked 49.5% of LLM-generated annotations and 50.5% of reasons as correct (with 9.2% of annotations and 12.5% of reasons additionally marked as partially correct).
 Reasoning models perform better, with 56.4% of their annotations and 58.3% marked as completely correct.
-The annotations made by Gemini 2.0 Flash Thinking and DeepSeek-R1 were the most accurate in D2T-Eval.
-o3-mini performed well across both tasks, even though Propaganda proved to be challenging for all models.
+The most accurate annotations on D2T-Eval were those made by Gemini 2.0 and DeepSeek-R1.
+o3-mini performed well on both tasks, although Propaganda proved to be challenging for all models.
 What are the sources of model errors?
@@
 All of these cases could be tackled by more descriptive guidelines or additional examples.
-However, in some cases, the models also misread or misinterpret data (e.g., claiming wind speed measurements do not exist when they do), which hints at deeper issues with understanding the data.
-Sources of incorrect reasons vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., counting omissions as sources of errors), to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors.
+However, in some cases, the models also misread or misinterpreted the data (e.g., claiming wind speed measurements do not exist when they do), which hints at deeper issues with understanding the data.
+Incorrect explanations vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., counting omissions as sources of errors) to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors.
 Occasionally, the model admits that it marked a correct span as an error, such as in “The description of the game’s duration aligns with the data, providing coherent information”.
-[^11] Note that this typically happens to GPT-4o.
-Since OpenAI API ensures JSON key ordering, the reason had to be generated before the annotation (same as for the other models).
+[^14] This typically happens to GPT-4o, even though OpenAI API ensures JSON key ordering so the explanation should have been generated before the annotation (cf.
+Section 3.2).
 How good are human annotations?
@@
 This fact led us to analyze the quality of human annotations in D2T-Eval.
-On a sample of 108 human annotations, we found that only 45.3% of annotations and 54.6% of reasons were marked as correct, which is comparable to LLM annotations.
+On a sample of 108 human annotations, we marked 45.3% of the annotations and 54.6% of the explanations correct, which is comparable to the LLM annotations.
 This suggests that the task is hard even for human annotators, and the quality of annotations from crowdworkers varies, even if they are preselected using a qualification task.
@@
 The color bars show annotations that we classified as correct, partially correct, assigned wrong category, incorrect, and undecidable cases.
-Detailed results are provided in Tables 22 and 23 in Appendix G.
+Detailed results are provided in Tables 22 and 23 in Appendix E.
 6 Discussion
 Can LLMs substitute human annotators?
-Our results paint a complex picture in this regard.
 The IAA between LLMs and human annotators is generally moderate.
 However, we also need to take into account a realistic upper bound: the average IAA between human annotators themselves.
-In that case, the strongest LLMs have already reached the level of human annotators in D2T-Eval and are not far behind in MT-Eval and Propaganda.
+In that case, the strongest LLMs have already reached the level of skilled crowdworkers in D2T-Eval.
 As we also showed, not all human annotations are “gold”, as even qualified crowdworkers make similar amounts of errors as LLMs.
-If we also factor in financial aspects, LLMs provide better flexibility, shorter response times, and much lower costs.
-In light of this, we argue that the benefits of automating span annotation with LLMs already outweigh its issues.
-Moreover, LLM-based span annotation will benefit from future increases in LLM capabilities.
+If we also factor in financial aspects, LLMs provide better flexibility, shorter response times, and lower costs.
+In light of this, we argue that the advantages of automating span annotation with LLMs outweigh its issues.
+LLM-based span annotation will also benefit from future increases in LLM capabilities.
 Recommendations
-To instantiate an LLM as a span annotator, it is beneficial to provide detailed guidelines describing how to handle all edge cases.
-In contrast, we note that providing specific examples (i.e., a few-shot setup) does not bring consistent improvements due to the length and complexity of typical examples for this task.
+To use an LLM as a span annotator, we strongly recommend providing detailed guidelines describing how to handle all edge cases.
+In contrast, we note that providing specific examples (i.e., a few-shot setup) does not bring consistent improvements, possibly due to the length and complexity of examples for this task.
 When using LLMs in a new setup, we recommend validating model’s annotations against examples hand-annotated by experts on a sample of the data.
-As for choosing the model, we recommend using a reasoning model (as opposed to an equivalent model without reasoning), since the reasoning models provide more reliable annotations, although at the cost of slightly higher response times and token count.
+Regarding model choice, we recommend choosing a reasoning model as opposed to an equivalent model without reasoning, since the reasoning models provide more reliable annotations at the cost of slightly higher response times and token count.
 Limitations
@@
 7 Conclusion
-We demonstrated that LLMs can serve as span annotators for three span annotation tasks: evaluation of data-to-text generation, evaluation of machine translation, and propaganda detection in human-written texts.
+We demonstrated that LLMs can serve as span annotators for three span annotation tasks: evaluating data-to-text generation, identifying errors in machine translation, and detecting propaganda in human-written texts.
 Our experiments show that LLMs achieve moderate agreement with skilled human annotators.
@@
 Ethics Statement
+The human evaluation study was approved by the internal ethics committee of our institution.
 Our human annotators were hired over Prolific and paid the platform-recommended wage of 9 GBP/hour (adjusted to slightly higher rates to account for real annotation times).
-The annotators were preselected based on their primary language (English).
-All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research (see Appendix D.2 for details).
-The domains were selected so that they do not contain any sensitive or potentially offensive content.
-We do not collect any demographic data about the participants.
+Annotators were pre-selected on the basis of their primary language (English).
+All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research (see Appendix B.2 for details).
+The domains were selected so that they do not contain sensitive or potentially offensive content.
+We do not collect demographic data about participants.
 Appendix A Implementation Details
-A.1 Costs and Time Requirements
-We estimate the cost and time requirements for a representative subset of annotators on D2T-Eval in Table 4.
-Metric crowdw. Llama 3.3 DeepSeek-R1 Claude 3.7 Sonnet o3-mini
-Cost per 1k outputs (\$) 500 - - 10.5 3.6
-Time per output (s) 129.1 21.6 227.5 9.0 21.8
-Table 4: Estimate of costs and time requirements for different kinds of annotators on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and two out of the proprietary models (Claude 3.7 Sonnet, and o3-mini).
-Note that we do not estimate the costs of the models we run locally.
-In practice, the costs of infrastructure acquisition and inference (GPUs, electricity) needs to be taken into account.
-The time estimates for the open models were computed on the infrastructure we describe in Section A.2.
-A.2 Open Models
+A.1 Open Models
 We run the local models using the ollama framework in 4-bit quantization. Specifically, we use the following local models for span annotations:
 We also use the following models for generating texts in D2T-Eval:
-For better reproducibility, we set the seed to 42 and temperature to 0 for the local models. We do not use these parameters for proprietary models, as these parameters are generally not supported.
+For better reproducibility, we set the seed to 42 and the temperature to 0 for the local models. We do not use these parameters for proprietary models as these parameters are generally not supported.
 We run the models on a set of GPUs, including NVIDIA H100 NVL (95G), AMD MI210 (64G), and NVIDIA RTX 3090 (24G).
-A.3 Proprietary Models
+A.2 Proprietary Models
 We use the following proprietary model versions:
@@
   • Gemini 2.0 Flash Thinking: gemini-2.0-flash-thinking-exp-01-21
-Appendix B Extracting Annotations from LLM Outputs
-B.1 Tagging the Input Text
-To the best of our knowledge, there is no established way of tagging the input text with decoder-only models.
-In the following, we discuss some of the existing approaches along with their pros and cons.
-  1. 1.
-    Labeling the text with special tokens Wang et al.
-(2023); Treviso et al.
-(2024): This approach instructs the LLM to generate the full copy of the input text.
-Spans are denoted using special XML-like tokens added to this copy.
-The approach is unambiguous in case the model copies the text correctly.
-However, it relies on the ability of the LLM to produce an exact copy of the (potentially very long) input text, and offers no fallback if the model fails to do that.
-  2. 2.
-    Listing start and end indices of spans Hasanain et al.
-(2024); Ramponi et al.
-(2025): This approach asks the LLM to generate numerical indices of the characters or tokens in the span.
-The approach would also make the output unambiguous, but it was discouraged by Hasanain et al., who notes that “[GPT-4] frequently generated indices not matching the corresponding spans in a paragraph”.
-Similarly, when models were asked to produce the output in CoNLL format Sang & Meulder (2003), “the models struggled to provide consistent outputs” Ramponi et al.
-(2025).
-A potential solution, providing the identifier for each token in the input text Ramponi et al.
-(2025), could be a confounding factor as it modifies the input text.
-  3. 3.
-    Listing textual content of spans Kasner & Dušek (2024); Kocmi & Federmann (2023); Zouhar et al.
-(2024): This approach asks the LLM to produce the textual content of the labeled spans along with their labels.
-The start and end indices are then found heuristically by case-insensitive string matching on the input text.
-This approach is more flexible, as the model does not need to copy the entire text, although it still relies on the ability of the models to literally copy text snippets.
-It can also misalign the model output by assigning the annotation to the first match in case there are multiple matches.
-Of the available options, we select listing textual content of spans as our method of choice.
-Although we are aware of its disadvantages, our preliminary experiments showed that the failures are infrequent: the models are generally able to copy text snippets literally and string matching collisions are rare.
-This approach is also the most robust, as failing to match a single annotation does not influence other matches.
-    "annotations": [{
-      "reason": [REASON],
-      "text": [SPAN_TEXT],
-      "type": [CATEGORY_IDX]
-Figure 5: Annotation output format for LLMs.
-B.2 Structured Output
-Having a high-level idea of the output format, we need to determine the exact format required from the model.
-Existing works collect model annotations using unstructured lists Kocmi & Federmann (2023); Zouhar et al.
-(2024) and structured output in the JSON format Kasner & Dušek (2024).
-Following our preliminary experiments, we decide to closely follow the setup of Kasner & Dušek (2024), using structured decoding with a fixed JSON scheme as we describe in Section 3.2 (see Figure 5).
-Our rationale is that JSON outputs are easier and more robust to parse, while being a common format that is natural for the LLMs.
-Structured output is now also getting available in LLM frameworks.
-[^12] See, e.g., https://ollama.com/blog/structured-outputs.
-Can structured output hurt model performance?
-Some authors voice concerns about the diminished performance of LLMs when using JSON output Tam et al.
-(2024); Beurer-Kellner et al.
-(2024).
-However, reactions to Tam et al.
-Kurt (2024); Castillo (2024) suggest that these issues do not manifest if several factors—such as describing the JSON format in the prompt—are controlled for.
-We leave experiments on this issue as a separate topic for future work.
-Parsing the output with reasoning chains
-At the time we performed our experiments, some of the reasoning models either did not support structured output (Gemini 2.0 Flash Thinking) or it would cancel out their ability to perform reasoning chains (ollama models [^13] https://github.com/ollama/ollama/issues/8529).
+A.3 Parsing the Output with Reasoning Traces
+At the time we performed our experiments, some of the reasoning models either did not support structured output (Gemini 2.0 Flash Thinking) or it would cancel out their ability to perform reasoning (ollama models [^15] https://github.com/ollama/ollama/issues/8529).
 Therefore, we use the following heuristic: we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response.
-This approach turns out to be able to robustly parse model responses containing reasoning chains with minimum failures.
-Appendix C Discussion on Data Contamination
-It is common practice not to disclose the data used for LLM training.
-That applies not only to the prorietary models, but also to the majority of the open models.
-As discussed in several recent works Balloccu et al.
-(2024); Dong et al.
-(2024); Jiang et al.
-(2024b), it is mostly certain that the LLMs were trained on some test sets of benchmarks that were available before their knowledge cut-off date.
-Memorization of these test sets may then artificially inflate the model performance on these benchmarks.
-In the context of D2T generation, Kasner & Dušek (2024) suggests to download a new set of structured data from publicly available sources – e.g., downloading up-to-date weather data for weather forecast generation.
-We use the method for D2T-Eval, where it creates only little overhead.
-For the other tasks, we are not able to follow the same process due to the costs of data collection. Nevertheless, we believe that it does not influence the validity of our results:
-  1. 1.
-    For MT-Eval, we use the WMT 2024 dataset that was released at the end of 2024, which is after the knowledge cut-off date of the LLMs we are using.
-  2. 2.
-    For Propaganda, the dataset is in a format that cannot be easily memorized.
-The model would need to match the word-level annotations for token ids with the text, which (as we also discuss in Section B.2) is not typically within model capabilities.
-  3. 3.
-    Lastly, D2T-Eval serves as a control task.
-If the model would underperform on D2T-Eval, but perform well on other tasks where the datasets were known, then we could suspect that the models memorized the data.
-In practice, we see the opposite: the models have the lowest performance on Propaganda which is the only task with a publicly available dataset.
-Appendix D D2T-Eval Details
-In this section, we provide more details on the process of collecting data (Section D.1) and annotations (Section D.2) for D2T-Eval.
-D.1 Collecting Data
-Our goal was to collect a novel dataset for D2T-Eval that is not subject to data contamination (cf.
-Appendix C).
-To this end, we follow the approach of Kasner & Dušek (2024) to collect structured data from publicly available APIs.
+Appendix B D2T-Eval Details
+In this section, we provide more details on the process of collecting data (Section B.1) and annotations (Section B.2) for D2T-Eval.
+B.1 Collecting Data
+To collect a novel dataset for D2T-Eval that is not subject to data contamination (cf.
+Section 4.4), we follow the approach of Kasner & Dušek (2024) to collect structured data from publicly available APIs.
 Domains
@@
 Additionally, we collect the new football domain from RapidAPI - API-Football.
-[^14] Our investigation showed that its structured data contain more detailed information about the game than ice_hockey and the sport itself is also more generally well-known.
+[^16] Its structured data contain more detailed information about the game than ice_hockey and the sport itself is also more generally well-known.
 Generating model outputs.
-We leverage LLMs to get output texts for the dataset.
-Specifically, we prompt two state-of-the-art models, Llama 3.3 Grattafiori et al.
+To obtain output texts, we prompt two state-of-the-art models, Llama 3.3 Grattafiori et al.
 (2024) and GPT-4o Hurst et al.
@@
 We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences.
-We provide the prompts in Appendix F.
+We provide the prompts in Appendix D.
 Note on annotating models’ own outputs.
-We point out that two of the four models that we use to collect annotations (Llama 3.3 and GPT-4o) are also later used to collect annotations.
-The fact that the models annotate their own outputs can reveal the self-preference bias Zheng et al.
+Two of the models (Llama 3.3 and GPT-4o) are also used later to collect annotations, which can reveal self-preference bias Zheng et al.
 (2023); Wataoka et al.
 (2024).
-Since our main point is not to rank the quality of generated outputs, we do not consider it a flaw, but rather a suitable ground for future investigation.
-We hypothesize that since span annotations are more grounded than direct assessment (and not directly ranking models), the self-preference bias will not be as prominent here.
-D.2 Collecting annotations
-Here we describe our process of writing annotation guidelines and selecting crowdworkers in order to ensure high quality of the collected annotations for D2T-Eval.
+Since we do not rank the quality of generated outputs, we do not consider it a flaw but rather a ground for future investigation.
+Moreover, we hypothesize that since span annotations are more grounded than direct assessment, the self-preference bias will not be as prominent.
+B.2 Collecting annotations
+Here we describe our process of writing annotation guidelines and selecting crowdworkers in order to ensure high quality of the annotations collected for D2T-Eval.
 Annotation guidelines
@@
 We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works Kasner & Dušek (2024); Thomson & Reiter (2020).
-Eventually, we settled on the list of annotation categories described in Appendix E (Table 5).
-Our list is an extended and updated version of the error taxonomy of Kasner & Dušek (2024), covering also coherence categories (Incoherent, Repetitive) on top of semantic accuracy.
+Eventually, we settled on the list of annotation categories described in Appendix C (Table 5).
+Our list is an extended and updated version of the error taxonomy of Kasner & Dušek (2024), also covering coherence categories (Incoherent, Repetitive) on top of semantic accuracy.
 Additionally, we agreed upon a list of general recommendations for collecting the annotations:
@@
   • The outputs in openweather can contain both precise numbers (e.g. 10.71°C) and the rounded ones (e.g. 11°C) as long as they agree with the data.
-We provide these additional recommendations in $\mathcal{P}_{\texttt{base}}$, while we omit them in $\mathcal{P}_{\texttt{noguide}}$ (cf. Figures 7 and 8).
+We provide these additional recommendations in $\mathcal{P}_{\texttt{base}}$, while we omit them in $\mathcal{P}_{\texttt{noguide}}$ (cf. Figures 6 and 7).
 Gold annotations
-Following the established annotation guidelines, we proceeded to collect our own internally-annotated gold data.
+With the annotation guidelines established, we proceeded to collect our own internally annotated gold data.
 The purpose of this data is two-fold: (1) to pre-select skilled crowdworkers, and (2) to create a smaller high-quality development set for the model prompting study.
-Since we also wanted to quantify our own internal IAA, we divided the data into two sets: $\mathcal{D}_{\texttt{dev}}$, containing 84 examples annotated individually by 7 annotators (12 examples per annotator), and $\mathcal{D}_{\texttt{dev}}$, containing 12 examples annotated commonly by each annotator.
-Our average IAA on $\mathcal{D}_{\texttt{dev}}$ was $F_{1}=0.433$ and $\gamma=0.399$ ([TODO: ] provides the average agreement of crowdworkers and models w.r.t.
-our annotations on this split).
+Since we also wanted to quantify our own internal IAA, we divided the data into two sets: $\mathcal{D}_{\texttt{dev}}$, containing 84 examples annotated individually by 7 annotators (12 examples per annotator), and $\mathcal{D}_{\texttt{iaa}}$, containing 12 examples annotated commonly by each annotator.
+Our average IAA on $\mathcal{D}_{\texttt{iaa}}$ was $F_{1}=0.444$ and $\gamma=0.399$.
 Qualification task
@@
 After the tutorial, we tested the worker performance on five manually pre-selected examples from $\mathcal{D}_{\texttt{dev}}$.
-For annotating the data in $\mathcal{D}_{\texttt{test}}$, we selected workers that had $F_{1}$-score higher than 0.5 w.r.t.
-our ground truth solution to the qualification task.
+For annotating the data in $\mathcal{D}_{\texttt{test}}$, we selected workers that had $F_{1}$ score higher than 0.5 w.r.t.
+our ground-truth solution to the qualification task.
 Of the 230 annotators who participated in the qualification task, 50 annotators (21.7%) passed this baseline.
@@
 Out of 50 selected participants, 45 participated in the main task (=90% turnover rate).
-We presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{dev}}$ (i.e, the examples from $\mathcal{D}_{\texttt{dev}}$ that we did not use for the qualification task).
-All the 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator.
-Additionally, 475 outputs (39.6%) were annotated by an extra annotator.
+We presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{dev}}$ (that is, the examples from $\mathcal{D}_{\texttt{dev}}$ that we did not use for the qualification task).
+All 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator.
+Furthermore, 475 outputs (39.6%) were annotated by an additional annotator.
 We payed all the annotators an average reward of 10.70 GBP / hour.
 We use examples with two annotators to compute the average IAA for D2T-Eval in Section 5.2.171717For other experiments, we use only the outputs from the first annotator as the reference data.
-D.3 Annotation Interface
+B.3 Annotation Interface
 We implement our annotation interface using factgenie. Figure 5 shows the main annotation interface, including an example from the openweather domain.
 Figure 5: Screenshot of the factgenie annotation interface used for collecting span annotations.
-Appendix E Annotation Categories
-Tables 5, 6 and 7 show an overview of annotation span categories that we used for our tasks along with their descriptions.
+Appendix C Annotation Categories
+Tables 5, 6 and 7 show an overview of the annotation span categories that we used for our tasks along with their descriptions.
 Category Name Description
@@
 Table 7: Annotation categories for the MT-Eval task. The categories are adopted from Da San Martino et al. (2019).
-Appendix F Prompts
-Figures 7, 8, 9 and 10 show the prompts for the D2T-Eval that we use for the experiments in Section 5.1.
+Appendix D Prompts
+Figures 6, 7, 8 and 9 show the prompts for the D2T-Eval that we use for the experiments in Section 5.1.
 Figure 10 shows the base prompt we used for MT-Eval, and Figure 11 shows the base prompt we used for Propaganda.
@@
 For more robust parsing, we initialize the model response with ’Sure, here is the summary: "’ .
-Appendix G Results
+Appendix E Results
 In this section, we provide detailed results of our experiments.
-The results for D2T-Eval are included in Sections G.1, G.2, G.3 and G.4:
-The results for MT-Eval are included in Sections G.5 and G.6:
-The results for Propaganda are included in Section G.7: Tables 20 and 21.
-The confusion matrices are included in Section G.8: Figures 14, 15 and 16.
-The manual evaluation results are included in Section G.9: Tables 22 and 23.
-G.1 D2T-Eval – Main Results
+The results for D2T-Eval are included in Sections E.1, E.2, E.3 and E.4:
+The results for MT-Eval are included in Sections E.5 and E.6:
+The results for Propaganda are included in Section E.7: Tables 20 and 21.
+The confusion matrices are included in Section E.8: Figures 13, 14 and 15.
+The manual evaluation results are included in Section E.9: Tables 22 and 23.
+E.1 D2T-Eval – Main Results
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
 Hard Soft Hard Soft Hard Soft $\Delta$
-Llama 3.3 0.307 0.132 0.276 0.185 0.388 0.154 0.323 0.169 0.109 0.418
-GPT-4o 0.346 0.178 0.300 0.180 0.303 0.179 0.301 0.122 0.130 0.429
-Claude 3.7 0.512 0.262 0.395 0.287 0.432 0.274 0.412 0.138 0.203 0.592
-DeepS. R1 0.453 0.293 0.493 0.154 0.259 0.202 0.340 0.138 0.185 0.645
-o3-mini 0.505 0.351 0.488 0.250 0.347 0.292 0.405 0.113 0.273 0.637
-Gem. 2-FT 0.458 0.259 0.434 0.236 0.395 0.247 0.414 0.167 0.209 0.612
+Llama 3.3 0.307 0.176 0.365 0.187 0.388 0.181 0.377 0.196 0.109 0.418
+GPT-4o 0.346 0.233 0.391 0.184 0.308 0.206 0.345 0.139 0.130 0.429
+Claude 3.7 0.512 0.294 0.442 0.304 0.457 0.299 0.449 0.150 0.203 0.592
+DeepS. R1 0.453 0.317 0.532 0.185 0.310 0.233 0.392 0.159 0.185 0.645
+o3-mini 0.505 0.392 0.542 0.285 0.395 0.330 0.457 0.127 0.273 0.637
+Gem. 2-FT 0.458 0.293 0.488 0.263 0.438 0.277 0.462 0.185 0.209 0.612
 Table 8: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. See Figure 2 for visualizaton of this table.
@@
 w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-G.2 D2T-Evalfootball domain
+E.2 D2T-Evalfootball domain
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
 Hard Soft Hard Soft Hard Soft $\Delta$
-Llama 3.3 0.512 0.241 0.408 0.295 0.500 0.266 0.449 0.183 0.197 0.455
-GPT-4o 0.422 0.271 0.394 0.263 0.382 0.267 0.388 0.121 0.190 0.427
-Claude 3.7 0.612 0.330 0.466 0.384 0.542 0.355 0.501 0.146 0.254 0.587
-DeepS. R1 0.561 0.414 0.643 0.237 0.368 0.301 0.468 0.167 0.276 0.659
-o3-mini 0.610 0.431 0.558 0.344 0.445 0.382 0.495 0.113 0.328 0.645
-Gem. 2-FT 0.588 0.387 0.571 0.314 0.464 0.347 0.512 0.165 0.307 0.647
+Llama 3.3 0.512 0.288 0.484 0.296 0.498 0.292 0.491 0.199 0.197 0.455
+GPT-4o 0.422 0.347 0.502 0.268 0.388 0.302 0.438 0.136 0.190 0.427
+Claude 3.7 0.612 0.360 0.511 0.391 0.555 0.375 0.532 0.157 0.254 0.587
+DeepS. R1 0.561 0.431 0.670 0.275 0.428 0.336 0.522 0.186 0.276 0.659
+o3-mini 0.610 0.470 0.608 0.384 0.496 0.422 0.547 0.125 0.328 0.645
+Gem. 2-FT 0.588 0.419 0.617 0.347 0.512 0.379 0.560 0.181 0.307 0.647
 Table 10: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the football domain of D2T-Eval.
@@
 w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-G.3 D2T-Evalgsmarena domain
+E.3 D2T-Evalgsmarena domain
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
 Hard Soft Hard Soft Hard Soft $\Delta$
-Llama 3.3 0.165 0.062 0.144 0.118 0.276 0.081 0.189 0.108 0.077 0.456
-GPT-4o 0.127 0.068 0.169 0.078 0.195 0.072 0.181 0.109 0.062 0.453
-Claude 3.7 0.214 0.146 0.290 0.121 0.240 0.133 0.263 0.130 0.140 0.656
-DeepS. R1 0.223 0.183 0.373 0.059 0.121 0.090 0.183 0.093 0.160 0.686
-o3-mini 0.172 0.201 0.429 0.078 0.167 0.113 0.240 0.127 0.192 0.689
-Gem. 2-FT 0.234 0.179 0.352 0.098 0.193 0.127 0.249 0.122 0.161 0.675
+Llama 3.3 0.165 0.107 0.247 0.122 0.282 0.114 0.263 0.149 0.077 0.456
+GPT-4o 0.127 0.107 0.269 0.082 0.205 0.093 0.233 0.140 0.062 0.453
+Claude 3.7 0.214 0.176 0.348 0.152 0.300 0.163 0.322 0.159 0.140 0.656
+DeepS. R1 0.223 0.199 0.405 0.089 0.180 0.123 0.249 0.126 0.160 0.686
+o3-mini 0.172 0.225 0.473 0.115 0.241 0.152 0.320 0.168 0.192 0.689
+Gem. 2-FT 0.234 0.218 0.426 0.128 0.249 0.161 0.314 0.153 0.161 0.675
 Table 12: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the gsmarena domain of D2T-Eval.
@@
 w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-G.4 D2T-Evalopenweather domain
+E.4 D2T-Evalopenweather domain
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
 Hard Soft Hard Soft Hard Soft $\Delta$
-Llama 3.3 0.209 0.092 0.270 0.118 0.346 0.103 0.303 0.200 0.053 0.272
-GPT-4o 0.404 0.164 0.298 0.158 0.288 0.161 0.293 0.132 0.122 0.377
-Claude 3.7 0.550 0.246 0.370 0.291 0.437 0.266 0.401 0.135 0.189 0.429
-DeepS. R1 0.454 0.215 0.391 0.129 0.234 0.162 0.293 0.131 0.124 0.533
-o3-mini 0.552 0.317 0.437 0.259 0.357 0.285 0.393 0.108 0.259 0.481
-Gem. 2-FT 0.458 0.197 0.366 0.241 0.447 0.217 0.403 0.186 0.155 0.368
+Llama 3.3 0.209 0.112 0.327 0.118 0.344 0.115 0.335 0.220 0.053 0.272
+GPT-4o 0.404 0.196 0.354 0.160 0.289 0.176 0.318 0.142 0.122 0.377
+Claude 3.7 0.550 0.274 0.411 0.292 0.437 0.283 0.424 0.141 0.189 0.429
+DeepS. R1 0.454 0.241 0.436 0.143 0.258 0.180 0.324 0.144 0.124 0.533
+o3-mini 0.552 0.362 0.496 0.267 0.366 0.307 0.422 0.115 0.259 0.481
+Gem. 2-FT 0.458 0.226 0.415 0.249 0.458 0.237 0.435 0.198 0.155 0.368
 Table 14: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the openweather domain of D2T-Eval.
@@
 w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-G.5 MT-Eval – Main Results
+E.5 MT-Eval – Main Results
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
 Hard Soft Hard Soft Hard Soft $\Delta$
-Llama 3.3 0.182 0.048 0.085 0.228 0.376 0.071 0.124 0.053 0.015 0.392
-GPT-4o 0.158 0.063 0.107 0.169 0.291 0.081 0.137 0.057 0.076 0.428
-Claude 3.7 0.301 0.117 0.174 0.304 0.435 0.151 0.223 0.071 0.131 0.628
-DeepS. R1 0.177 0.077 0.121 0.119 0.187 0.084 0.132 0.047 0.058 0.631
-o3-mini 0.197 0.085 0.143 0.137 0.234 0.096 0.162 0.066 0.100 0.646
-Gem. 2-FT 0.257 0.104 0.172 0.117 0.211 0.102 0.174 0.071 0.066 0.710
+Llama 3.3 0.182 0.121 0.200 0.229 0.378 0.155 0.257 0.102 0.014 0.392
+GPT-4o 0.158 0.141 0.240 0.195 0.327 0.156 0.266 0.110 0.076 0.428
+Claude 3.7 0.301 0.226 0.325 0.335 0.469 0.262 0.373 0.111 0.131 0.628
+DeepS. R1 0.177 0.169 0.268 0.183 0.280 0.168 0.262 0.094 0.058 0.631
+o3-mini 0.197 0.169 0.291 0.161 0.275 0.160 0.275 0.115 0.100 0.646
+Gem. 2-FT 0.257 0.184 0.312 0.180 0.339 0.173 0.304 0.130 0.066 0.710
 Table 16: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages.
 Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
-Human 2405 0.62 70.04 13.97
+Human 2090 0.7 66.0 14.5
 Llama 3.3 6361 2.3 6.2 17.4
@@
 w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-G.6 MT-Eval – By Language
-Language $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
+E.6 MT-Eval – By Language
+Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
 Hard Soft Hard Soft Hard Soft $\Delta$
-en-cs 0.303 0.085 0.130 0.205 0.323 0.117 0.180 0.063 0.084 0.582
-en-es 0.171 0.063 0.117 0.164 0.313 0.086 0.161 0.075 0.080 0.631
-en-hi 0.170 0.042 0.059 0.160 0.237 0.064 0.091 0.027 -0.0 0.552
-en-is 0.347 0.217 0.368 0.149 0.255 0.173 0.294 0.121 0.108 0.493
-en-ja 0.127 0.033 0.046 0.256 0.320 0.058 0.078 0.020 0.063 0.569
-en-ru 0.225 0.120 0.170 0.172 0.250 0.136 0.195 0.059 0.162 0.588
-en-uk 0.192 0.063 0.097 0.172 0.276 0.089 0.139 0.050 0.031 0.542
-en-zh 0.163 0.036 0.083 0.155 0.339 0.058 0.130 0.072 0.075 0.623
+en-cs 0.303 0.144 0.268 0.180 0.326 0.156 0.286 0.130 0.084 0.582
+en-es 0.171 0.161 0.243 0.236 0.362 0.190 0.288 0.098 0.080 0.631
+en-hi 0.170 0.173 0.265 0.208 0.327 0.173 0.269 0.096 -0.0 0.552
+en-is 0.347 0.136 0.246 0.187 0.361 0.145 0.269 0.124 0.108 0.493
+en-ja 0.127 0.193 0.302 0.249 0.363 0.209 0.318 0.109 0.063 0.569
+en-ru 0.225 0.178 0.256 0.273 0.386 0.208 0.298 0.090 0.162 0.588
+en-uk 0.192 0.166 0.254 0.214 0.339 0.184 0.286 0.102 0.031 0.542
+en-zh 0.163 0.196 0.346 0.163 0.294 0.169 0.302 0.133 0.075 0.623
 Table 18: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language (average across models).
@@
 w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-G.7 Propaganda – Main Results
+E.7 Propaganda – Main Results
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
 Hard Soft Hard Soft Hard Soft $\Delta$
-Llama 3.3 0.336 0.069 0.253 0.063 0.231 0.066 0.242 0.176 0.092 0.343
-GPT-4o 0.344 0.092 0.294 0.036 0.114 0.051 0.164 0.113 0.066 0.234
-Claude 3.7 0.460 0.109 0.277 0.088 0.225 0.098 0.249 0.151 0.155 0.113
-DeepS. R1 0.354 0.081 0.257 0.058 0.182 0.067 0.213 0.146 0.091 0.476
-o3-mini 0.418 0.152 0.431 0.065 0.183 0.091 0.257 0.166 0.154 0.517
-Gem. 2-FT 0.560 0.106 0.277 0.186 0.484 0.135 0.353 0.218 0.202 0.493
-Table 20: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on Propaganda.
+Llama 3.3 0.336 0.070 0.243 0.063 0.219 0.066 0.230 0.164 0.092 0.343
+GPT-4o 0.344 0.095 0.293 0.038 0.115 0.054 0.166 0.112 0.066 0.234
+Claude 3.7 0.460 0.110 0.274 0.096 0.239 0.103 0.255 0.152 0.155 0.113
+DeepS. R1 0.354 0.083 0.246 0.062 0.182 0.071 0.209 0.138 0.091 0.476
+o3-mini 0.418 0.152 0.411 0.066 0.179 0.092 0.249 0.157 0.154 0.517
+Gem. 2-FT 0.560 0.106 0.268 0.190 0.477 0.136 0.343 0.207 0.202 0.493
+Table 20: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the Propaganda.
 Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
@@
 Gem. 2-FT 1864 18.5 3.0 54.1
-Table 21: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on Propaganda.
+Table 21: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the Propaganda.
 Ann=# of annotations, Ann/Ex=annotations per example.
 w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-G.8 Confusion Matrices
+E.8 Confusion Matrices
 (a) Llama 3.3
@@
 Figure 15: Confusion matrix of category counts averaged across models for Propaganda (see Table 7 for the description of categories).
-G.9 Manual Analysis
+E.9 Manual Analysis
 Annotation Reason
@@
 Table 23: Manual evaluation results for Propaganda domain. Categories for annotation and reason: C=Correct, P=Partially correct, W=Wrong category, I=Incorrect, U=Undecidable.
-Appendix H Related Work (extended)
+Appendix F Related Work (extended)
 This section provides more discussion on related work following Section 2.
@@
 For these reasons, Kocmi et al.
-(2024b), based on prior work by Popovic (2020), introduce ESA (Error Span Annotation) for machine translation evaluation.
+(2024c), based on prior work by Popovic (2020), introduce ESA (Error Span Annotation) for machine translation evaluation.
 This protocol drops the error categorization, which is primarily useful for diagnosis and not evaluation, simplifies severities, and has the annotators also assign the final score after they have done the span annotations.
@@
 Although computationally more efficient than prompting larger LLMs, these two metrics either require reference texts, or are limited to predefined evaluation aspects.
-Kartáč et al.
+Kartác et al.
 (2025) apply an ensemble of open-weight LLMs as evaluators to a variety of tasks and aspects, generating both numeric scores and error spans with detailed explanations.
@@
 Goel et al. (2023) apply LLMs for medical text annotation, identifying and extracting spans containing specific named entities.
-Appendix I Examples
+Appendix G Examples
 In Tables 26, 25 and 24, we show examples of the annotated outputs for our tasks.

### v3

@@
-Large Language Models as Span Annotators
+LLMs as Span Annotators: A Comparative Study of LLMs and Humans
 Abstract
-Span annotation is the task of localizing and classifying text spans according to custom guidelines.
-Annotated spans can be used to analyze and evaluate high-quality texts for which single-score metrics fail to provide actionable feedback.
-Until recently, span annotation was limited to human annotators or fine-tuned models.
-In this study, we show that large language models (LLMs) can serve as flexible and cost-effective span annotation backbones.
-To demonstrate their utility, we compare LLMs to skilled human annotators on three diverse span annotation tasks: evaluating data-to-text generation, identifying translation errors, and detecting propaganda techniques.
-We demonstrate that LLMs achieve inter-annotator agreement (IAA) comparable to human annotators at a fraction of a cost per output annotation.
-We also manually analyze model outputs, finding that LLMs make errors at a similar rate to human annotators.
-We release the dataset of more than 40k model and human annotations for further research.
+Span annotation – annotating specific text features at the span level – can be used to evaluate texts where single-score metrics fail to provide actionable feedback.
+Until recently, span annotation was done by human annotators or fine-tuned models.
+In this paper, we study whether large language models (LLMs) can serve as an alternative to human annotators.
+We compare the abilities of LLMs to skilled human annotators on three span annotation tasks: evaluating data-to-text generation, identifying translation errors, and detecting propaganda techniques.
+We show that overall, LLMs have only moderate inter-annotator agreement (IAA) with human annotators.
+However, we demonstrate that LLMs make errors at a similar rate as skilled crowdworkers.
+LLMs also produce annotations at a fraction of the cost per output annotation.
+We release the dataset of over 40k model and human span annotations for further research.
 [^1] Project website: https://llm-span-annotators.github.io
+LLMs as Span Annotators: A Comparative Study of LLMs and Humans
+Zdeněk Kasner1   Vilém Zouhar2   Patrícia Schmidtová1 Ivan Kartáč1Kristýna Onderková1Ondřej Plátek1 Dimitra Gkatzia3Saad Mahamood4Ondřej Dušek1Simone Balloccu5 1Charles University 2ETH Zurich 3Edinburgh Napier University 4trivago N.V.
+5TU Darmstadt, Germany Contact: kasner@ufal.mff.cuni.cz
 1 Introduction
-Fine-grained aspects of texts, such as faithfulness or coherence, depend on local lexical choices.
-In order to reflect these aspects in quality judgments of texts, we need analysis techniques that provide the appropriate amount of detail.
-However, most automatic evaluation metrics for natural language generation (NLG) are based on direct assessment, producing holistic scores for each evaluated aspect Gkatzia & Mahamood (2015); Sai et al.
-(2023); Schmidtová et al.
-(2024).
-Although numerical values make it easy to rank systems, these metrics are too simplistic and susceptible to biases or miscalibration of the underlying models Gehrmann et al.
-(2023); Gao et al.
-(2024); Liu et al.
-(2024); Wang et al.
-(2024).
-The subject of our study, span annotation (Figure 1), offers an alternative approach.
-Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans and classify them according to task-specific guidelines.
-In contrast to numerical rating, annotations are aligned to the evaluated text, which makes them more explainable and actionable.
-The annotations can also be examined post hoc, which brings more accountability into the evaluation process.
+Fine-grained aspects of texts, such as semantic accuracy or coherence, depend on local lexical choices.
+To reflect these aspects in quality judgments of texts, techniques are needed that provide the appropriate amount of detail.
+However, most automatic evaluation metrics for Natural Language Generation (NLG) assign only singular scores for the whole text per each evaluated aspect gkatzia2015snapshot; sai2022survey; schmidtova2024automatic.
+Although numerical values make it easy to rank systems, these metrics are too crude and susceptible to biases or miscalibration of the underlying models gehrmann2022repairing; liu2023calibrating; wang2023large; gao2024llm.
+The subject of our study, span annotation (Figure˜1), offers an alternative approach.
+Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans of interest and classify them according to task-specific guidelines.
+Span annotations are aligned to specific parts of the evaluated text, which makes them more explainable and actionable than numerical ratings.
 Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation.
-The method traditionally required human annotators, making it costly and difficult to scale Da San Martino et al.
-(2019); Thomson & Reiter (2020); Popovic (2020); Kocmi et al.
-(2024c).
-The situation may change with the upcoming LLM-as-a-judge paradigm, in which LLMs follow instructions for the evaluation process similarly to human annotators.
-However, current LLM-as-a-judge applications still focus mainly on document-level classification or numerical ratings on the Likert scale Zheng et al.
-(2023); Gu et al.
-(2024).
-In this work, we investigate whether LLMs can be applied as a substitute for human annotators on span annotation tasks.
-First, we focus on how to produce span annotations with LLMs.
-We discuss input tagging methods and experiment with prompting techniques, looking for a setup that is robust and widely applicable.
-Next, we compare LLMs with skilled human annotators in terms of IAA.
-Where possible, we contrast the IAA between LLMs and human annotators with the upper bound of the IAA between the annotators themselves.
-Finally, we manually analyze a subset of LLM and human annotations to assess the validity of the annotations and their respective explanations.
-For our experiments, we select three diverse span annotation tasks (cf.
-Section 3.1): evaluating data-to-text generation Thomson & Reiter (2020), identifying errors in machine translation Kocmi et al.
-(2024a), and detecting propaganda techniques in human-written texts Da San Martino et al.
-(2019).
-Note that unlike established span annotation tasks such as named entity recognition Tjong Kim Sang & De Meulder (2003); Wang et al.
-(2025a), there are no general purpose fine-tuned models for these tasks since annotation guidelines are specified ad hoc, making their automation difficult.
-We base our span annotators on open and proprietary state-of-the-art LLMs, including recent reasoning LLMs that leverage test-time compute for better performance on challenging tasks Muennighoff et al.
-(2025); Marjanović et al.
-(2025).
+The method traditionally required human annotators, making it costly and difficult to scale dasanmartino2019finegrained; thomson2020gold; popovic2020informative; kocmi2024error.
+The LLM-as-a-judge paradigm recently emerged as a promising solution to this problem zheng2023judging; gu2025surveyllmasajudge, allowing task-specific applications kocmi2023gembamqm; hasanain2024large.
+However, to our knowledge, no study has systematically compared span annotation performance between LLMs and human annotators.
+Figure 1: Examples of span annotation tasks that we automate with LLMs. We unify the setup for evaluation tasks (D2T-Eval, MT-Eval) and text analysis tasks (Propaganda).
+The central focus of our investigation is comparing human annotators and state-of-the-art LLMs on span annotation tasks.
+We select three span annotation tasks (cf.
+Section˜3.1): evaluating data-to-text generation thomson2020gold, identifying errors in machine translation kocmi-etal-2024-findings, and detecting propaganda techniques in human-written texts dasanmartino2019finegrained.
 Our contributions are as follows:
   1. 1.
-    We demonstrate that zero-shot prompting in combination with structured outputs can be a robust way to set up LLMs as span annotators, given that the model is provided detailed annotation guidelines (Sections 3.2 and 5.1)
+    We establish that with structured outputs and detailed annotation guidelines, LLMs can serve as robust span annotators (Sections˜3.2 and 5.1).
   2. 2.
-    Using a combination of specialized IAA metrics and manual analysis, we show that performance of LLMs is close to skilled human annotators, in some scenarios leveling or surpassing the performance of qualified crowdworkers (Sections 5.2 and 5.3).
+    We show that LLMs have moderate inter-annotator agreement with human annotators overall, but can reach the agreement of qualified crowdworkers (Section˜5.2).
   3. 3.
-    We release a dataset of more than 40k human and model annotations for all our tasks and setups, including reasoning traces from reasoning LLMs and annotations collected from skilled crowdworkers.
-Figure 1: Examples of span annotation tasks that we automate with LLMs. We unify the setup for evaluation tasks (D2T-Eval, MT-Eval) and text analysis tasks (Propaganda).
+    We discover the sources of model errors by manually analyzing a subset of LLM outputs (Section˜5.3).
+  4. 4.
+    We release a dataset of more than 40k human and model annotations, including annotations collected from crowdworkers and reasoning traces from reasoning LLMs.
 2 Related Work
 LLMs for NLG Evaluation.
-Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts Sai et al.
-(2023); Schmidtová et al.
-(2024).
-As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgments Novikova et al.
-(2017); Reiter (2018).
-With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators across diverse tasks, either with simple numeric scoring (Bavaresco et al., 2024; Liu et al., 2023; Sottana et al., 2023; Leiter et al., 2023; Chiang & Lee, 2023) or with detailed feedback (Li et al., 2024; Kim et al., 2024a; b; Xu et al., 2023).
-However, the scores produced by LLMs still miss fine-grained text aspects and are influenced by LLM biases Stureborg et al.
-(2024); Koo et al.
-(2024); Wang et al.
-(2024).
+Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts sai2022survey; schmidtova2024automatic.
+As such, they are unable to quantify more fine-grained aspects (gehrmann2022repairing; 10.1162/tacl_a_00437) and do not correlate well with human judgments novikova-etal-2017-need; reiter-2018-structured.
+With the emerging LLM-as-a-judge paradigm (gu2025surveyllmasajudge), LLMs have been applied as evaluators on various tasks, using simple numeric scoring (bavaresco2024llms; liu2023geval; sottana2023evaluation; leiter2023eval4nlp; chiang2023can), or free-form feedback (li2024generative; kim2024prometheusa; kim2024prometheus; kartac2025openlgauge).
+However, as these outputs are not firmly grounded in text, they tend to miss fine-grained aspects and are influenced by LLM biases stureborg2024large; koo2023benchmarking; wang2023large.
 Span Annotation Protocol.
-In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024c), where annotators mark erroneous spans in translations.
-In data-to-text (D2T) generation, span annotation is applied by Thomson & Reiter (2020), who introduce a span-based evaluation protocol for annotation of generated basketball reports.
-Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection Da San Martino et al.
-(2019) and text summarization Subbiah et al.
-(2024).
-Note that unlike our work, these works perform span annotation only using human annotators.
-Automating Span Annotation.
-For all the aforementioned tasks, first attempts at automating span annotation were based on fine-tuned pre-trained encoder models.
-That includes MT (Guerreiro et al., 2024), D2T generation Kasner et al.
-(2021), text summarization Goyal et al.
-(2022), and propaganda detection Martino et al.
-(2020); Goffredo et al.
-(2023); Piskorski et al.
-(2023).
+In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (lommel2014multidimensional; mariana2014multidimensional; popovic2020informative; kocmi2024error), where human annotators mark erroneous spans in translations.
+In data-to-text (D2T) generation, span annotation was applied by thomson2020gold, who introduced a span-based evaluation protocol for annotation of generated basketball match reports.
+Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection dasanmartino2019finegrained and text summarization subbiah2024reading.
+Unlike our work, these works focus on span annotation with human annotators.
+Automatic Span Annotation.
+Early attempts at automating span annotation with ad-hoc guidelines were based on fine-tuned pre-trained encoder models.
+That includes evaluation of MT (10.1162/tacl_a_00683), D2T generation kasner2021textincontext or text summarization goyal2022snac, as well as propaganda detection martino2020semeval2020; goffredo2023argumentbased; piskorski2023multilingual.
 Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities.
-We build on recent work that applies LLMs as a task-specific evaluation tool (Kocmi & Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner & Dušek, 2024; Chang et al., 2024; Zouhar et al., 2025; Kartác et al., 2025; Ramponi et al., 2025).
+We build on recent work that applies LLMs as a task-specific evaluation tool (kocmi2023gembamqm; fernandes-etal-2023-devil; hasanain2024large; kasner2024traditional; chang2024booookscore; zouhar2024aiassisted; kartac2025openlgauge; ramponi2025finegrained).
 Our work is the first that unifies diverse span annotation tasks within a single framework and systematically compares the performance of LLMs to human annotators.
-See more discussion on related work in Appendix F.
 3 Automating Span Annotation with LLMs
-In this section, we first formally introduce the span annotation process in Section 3.1.
-Next, we discuss how to automate the process with LLMs in Section 3.2 and how to evaluate the quality of span annotations in Section 3.3.
+We first formally introduce the span annotation process in Section˜3.1.
+Next, we discuss how to automate the process with LLMs in Section˜3.2 and how to evaluate the quality of span annotations in Section˜3.3.
 3.1 Span Annotation: Task Definition
-The goal of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$ given:
+The aim of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$ given:
   • the set of categories $C=\{c_{1},\ldots,c_{k}\}$,
-  • the annotation guidelines $\mathcal{G}$ (including category definitions and other instructions),
+  • the annotation guidelines $\mathcal{G}$,
   • the source $X$ (such as the translation source; empty if we are annotating only intrinsic text aspects).
-The goal is to produce a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i},r_{i}\rangle$:
+The output is a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i},r_{i}\rangle$:
   • $s_{i},e_{i}\in\{1,\ldots,n\},s_{i}<e_{i}$ are the start and end indices of the annotated span,
   • $c_{i}\in C$ is the assigned annotation category,
-  • $r_{i}$ is an explanation for the annotation (optional). [^2] We collect the explanations from LLMs, but not from human annotators due to time considerations.
-3.2 Automating Span Annotation with LLMs
-To automate span annotation, we collect the annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$ from an LLM:
-$A=\textsc{LLM}(\text{prompt}(Y,C,\mathcal{G},X)).$ (1)
-How to Identify the Spans?
-An annotation needs to contain the start and end indices of the span.
-For encoder models, the task would be trivially solved by classifying tokens on top of the encoder’s hidden states.
-However, to the best of our knowledge, there is no established way of tagging the input text with decoder-only models.
-Of the methods from previous work, we discarded surrounding spans with special XML-like tokens Wang et al.
-(2025b); Treviso et al.
-(2024).
-This method requires the LLM to generate full copy of the input text and does not offer a fallback if the model fails to do so.
-We also rule out listing start and end indices of spans Hasanain et al.
-(2024); Ramponi et al.
-(2025), as Hasanain et al.
-note that the model frequently generates indices that do not match the corresponding spans in a paragraph.
-Instead, we adopted the method of listing textual content of the spans Kocmi & Federmann (2023); Zouhar et al.
-(2025); Kasner & Dušek (2024).
-Here, the start and end indices of the span are found heuristically by string matching of the span’s textual content.
-With this approach, we do not need to modify the input and failing to match a single annotation does not influence other matches.
-[^3] We do note that this approach can misalign the model output by assigning the annotation to the first match, but our preliminary experiments showed that the collisions are rare.
-Parsing Model Output
-To obtain the annotations, we request the list of annotations in JSON format, following the setup of Kasner & Dušek (2024).
-This setup allows us to use constrained decoding with a defined JSON scheme, ensuring that the output is syntactically valid.
+  • $r_{i}$ is a reason for the annotation (optional).
+3.2 Span Annotation with LLMs
+In our setup, annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$ are collected from an LLM:
+$A=\textsc{LLM}(\text{prompt}(Y,C,\mathcal{G},X)).$
+To obtain the annotations, we follow the setup of kasner2024traditional: we request the list of annotations in JSON format, using constrained decoding with a fixed JSON scheme to ensure that the output is syntactically valid.
 We require each annotation to contain the fields reason (the explanation $r_{i}$), text (the textual content of the span), and type (the integer index of the error category $c_{i}$).
-[^4] Given the considerations of Castillo (2024), we ensure that the reason field has precedence.
+[^2] Following castillo2024structured, we ensure that the reason field is generated first.
+For reasoning models not supporting structured output, we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response.
 3.3 Evaluating Span Annotations
 To compare annotations automatically, we need a notion of similarity between two sets of annotations $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\hat{\mathcal{A}}=\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $\mathcal{Y}=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$.
-We consider three similarity metrics:
+We consider the following similarity metrics:
 Pearson correlation $\rho$ over counts.
-This metric simply compares how many spans were annotated for each example:
-$\mathrm{CountCorrelation}(\mathcal{A},\hat{\mathcal{A}})=\mathrm{Pearson}(\langle|A_{Y}|,|\hat{A}_{Y}|\rangle_{Y\in\mathcal{Y}})$ (2)
-This metric serves as a sanity check: a low value would suggest that an annotator either skips examples or over-annotates, indicating unclear annotation guidelines.
+This metric compares how many spans were annotated for each example:
+$\operatorname{Pearson}(\mathcal{A},\hat{\mathcal{A}})=\operatorname{\rho}(\langle|A_{Y}|,|\hat{A}_{Y}|\rangle_{Y\in\mathcal{Y}})$ (1)
+The correlation serves as a sanity check: a low value would suggest that an annotator either skips examples or over-annotates, indicating unclear annotation guidelines.
 Precision, Recall, and $\mathbf{F_{1}}$.
-To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined in Da San Martino et al.
-(2019).
-These measures are on matching annotations, adjusted to give partial credit to imperfect matches: $\text{Precision}(A_{Y},\hat{A}_{Y})=\frac{1}{|A_{Y}|}\sum_{a\in A_{Y}}\frac{|a\cap\hat{a}|}{|a|},$ (3) $\text{Recall}(A_{Y},\hat{A}_{Y})=\frac{1}{|\hat{A}_{Y}|}\sum_{\hat{a}\in\hat{A}_{Y}}\frac{|a\cap\hat{a}|}{|\hat{a}|},$ (4) where $a\cap\hat{a}$ is the character overlap between two annotation spans and $|a|=e-s+1$ is the length of the annotation span in characters (see Section 3.1).
+To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined by dasanmartino2019finegrained.
+These measures are on matching annotations, adjusted to give partial credit to imperfect matches:
+$\hskip-31.29802pt\operatorname{Precision}(A_{Y},\hat{A}_{Y})=\frac{1}{|A_{Y}|}\sum_{a\in A_{Y}}\frac{|a\cap\hat{a}|}{|a|},\hskip-22.76219pt$ (2)
+$\operatorname{Recall}(A_{Y},\hat{A}_{Y})=\frac{1}{|\hat{A}_{Y}|}\sum_{\hat{a}\in\hat{A}_{Y}}\frac{|a\cap\hat{a}|}{|\hat{a}|}$, (3)
+where $a\cap\hat{a}$ is the character overlap between two annotation spans and |a|=es+1|a|=e-s+1 is the length of the annotation span in characters (see Section˜3.1).
 Subsequently, we compute the $F_{1}$-score as the harmonic mean of precision and recall.
@@
 The $F_{1}$ score is sensitive to varying span granularities and does not consider near-matches with no overlap or agreement by chance.
-To this end, we follow Da San Martino et al.
-(2019) and Hasanain et al.
-(2024) in using the $\gamma$ score (Mathet et al., 2015) as a complementary metric.
-In short, the metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations.
-The score ranges from negative infinity to 1, where 1 is achieved when the annotations are perfectly aligned.
-The $\gamma$ score extends Krippendorff’s $\alpha$ (Krippendorff, 1980), another popular metric, by computing the category-aware span alignments.
-We use the implementation of Titeux & Riad (2021).
+To this end, we follow dasanmartino2019finegrained and hasanain2024large in using the $\gamma$ score (mathet2015unified) as a complementary metric.
+The metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations.
+The score ranges from -$\infty$ to 1, where 1 is achieved when the annotations are perfectly aligned.
+The $\gamma$ score extends Krippendorff’s $\alpha$ (krippendorff1980content), another popular metric, by computing the category-aware span alignments.
+We use the implementation of Titeux2021.
 $S_{\emptyset}$ score.
-For an output $y$, one or both annotation sets $A_{Y},\hat{A}_{Y}$ can be empty.
-Having these cases is desirable: e.g.
-if the goal is to annotate errors in an output that is entirely correct.
-However, neither of the introduced scores accounts for these cases.
-[^5] The $\gamma$ score is undefined for less than two non-empty annotation sets: these examples thus need to be skipped during the $\gamma$ computation.
-The F1 score does not account for examples with no annotations.
-To compensate for that, we introduce a score $S_{\emptyset}$ that is computed for examples with less than two nonempty sets of annotations as $S_{\emptyset}=1/(1+|A|)$, where:
+For an output $y$, one or both annotation sets $A_{Y},\hat{A}_{Y}$ may be empty.
+This is in fact desirable, e.g., if the goal is to annotate errors in an output that is entirely correct.
+However, these cases are not properly reflected by the other scores we are using: the F1 score only focuses on counting error spans and is not affected by true negatives, and the $\gamma$ score is undefined if any of the two annotation sets is empty (these examples therefore need to be skipped during the $\gamma$ computation).
+To compensate for this, we introduce a score $S_{\emptyset}$that is computed for examples where any of $A_{Y}$, $\hat{A}_{Y}$ is empty:
+$S_{\emptyset}=1/(1+|A|),$ (4)
+where:
 $|A|=\begin{cases}|A_{Y}|&\text{if }|\hat{A}_{Y}|=0,\\ |\hat{A}_{Y}|&\text{otherwise}.\end{cases}$ (5)
@@
 # Cat.
-denotes the number of categories used in the task (see Appendix C for their listings), # Texts the number of texts annotated, and Avg.
-Len the average number of words/characters in the output.
-We cover three span annotation tasks of different quantitative and qualitative aspects: evaluating data-to-text generation (D2T-Eval; Section 4.1), identifying errors in machine translation (MT-Eval; Section 4.2), and detecting propaganda techniques (Propaganda; Section 4.3).
-In Table 1, we provide a quantitative overview of our tasks.
-[^6] See also Figure 1 and Appendix G for the visualization of examples of our tasks.
-4.1 Task: Evaluation of Data-to-text Generation
-In D2T-Eval, we focus on evaluating the semantic accuracy and stylistic aspects of texts generated from structured data Celikyilmaz et al.
-(2020); Sharma et al.
-(2022).
-The input is the text $Y$ generated from the structured data $X$.
-Subtasks and input data.
-We select three D2T generation subtasks: generating five-day weather forecasts (openweather), generating product descriptions (gsmarena), and generating reports of soccer games (football).
-We use the approach of Kasner & Dušek (2024) to download novel structured data from publicly available APIs to make sure that they are unseen for all LLMs (cf.
-Section 4.4 on data contamination).
-We provide more details on the data collection process in Section B.1.
-Collecting model outputs.
-We generate outputs for the structured inputs using LLMs.
-Specifically, we selected two state-of-the-art models, Llama 3.3 70B Grattafiori et al.
-(2024) and GPT-4o Hurst et al.
-(2024), and two smaller models, Gemma 2 2B Team et al.
-(2024) and Phi 3.5 3.8B Abdin et al.
-(2024).
-We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences.
-[^7] See Sections A.1 and A.2 for details on the models and Appendix D for the prompts.
-Data splits.
-For further experiments, we divide the data into multiple subsets:
-  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – the subset for LLM evaluation, annotated with crowdworkers,
-  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – the subset for the study of prompt variants, annotated internally,
-  • $\mathcal{D}_{\texttt{iaa}}$ (12 outputs) – the control subset for human crowdworkers, annotated internally.
-Crowdsourcing annotations
-We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com.
-We apply best practices for gathering human annotations, including an iterative process of refining annotation guidelines and preselecting best-performing annotators using a qualification task Tseng et al.
-(2020); Iskender et al.
-(2020); Huang et al.
-(2023); Zhang et al.
-(2023).
-Of the 230 annotators that participated in the qualification task, we selected 50 best annotators to annotate our data.
-We provide more details on the process of collecting human annotations in Section B.2.
-4.2 Task: Identifying Errors in Machine Translation
-For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024b).
-The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024c) by professional translators.
-We focus on character-level span annotations of major and minor translation errors provided in the dataset.
-The inputs $X$ in MT-Eval are the source texts used to generate the translation $Y$ [^8] Note unlike for the other tasks, the annotations in MT-Eval cannot overlap and need not to be aligned with word boundaries.
-We use data for the three textual domains present in the WMT 2024 shared task: news, literary, and social.
-We focus on data translated from English into other languages, including Chinese, Czech, German, Hindi, Icelandic, Japanese, Russian, Spanish, Ukrainian.
-The original dataset has nearly 50k model outputs, making it too costly to run the full evaluation campaign.
-Therefore, we used all available system outputs for 10 sampled inputs for each language, giving us 2,854 outputs that we use for the experiments.
-4.3 Task: Propaganda Technique Detection
-For the Propaganda task, we use the dataset of Da San Martino et al.
-(2019).
-The dataset contains news collected from mostly propagandistic on-line sources.
+denotes the number of categories used in the task (see Appendix˜C for their listings), # Texts the number of texts annotated, Avg.
+Len the average number of words/characters in the output, and Novel Data indicates newly collected data.
+4.1 Tasks
+We cover three span annotation tasks of different qualitative aspects.
+We focus on tasks that do not have extensive training data resources and cannot be readily solved by encoder models (such as, e.g., named entity tagging): evaluating data-to-text generation (D2T-Eval; Section˜4.1.1), identifying errors in machine translation (MT-Eval; Section˜4.1.2), and detecting propaganda techniques (Propaganda; Section˜4.1.3).
+See Table˜1 for an overview of our datasets.
+4.1.1 D2T-Eval: Evaluation of Data-to-text Generation
+In D2T-Eval, we use span annotation to evaluate semantic accuracy and stylistic aspects of data-to-text generation outputs sharma2022innovations; celikyilmaz2021evaluation.
+The inputs $X$ are the structured data used to generate the output text $Y$.
+We use D2T-Eval as a control task to mitigate the effects of data contamination: the fact that the performance of the model might be inflated by previous exposure to publicly available benchmarks balloccu2024leak; dong2024generalization; jiang2024investigating.
+Instead of using an existing dataset, we use the Quintd tool kasner2024traditional to download structured inputs from multiple public APIs.
+[^3] We selected two of the existing domains: openweather for generating weather forecasts and gsmarena for generating phone descriptions.
+We also add the football domain (using RapidAPI - API-Football) for generating soccer game reports.
+To obtain output texts for the structured data, we prompt LLMs in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences.
+See Appendix B.1 for more details.
+To gather annotations for our dataset, we use crowdworkers from Prolific.com.
+We apply best practices for gathering human annotations, including an iterative process to refine annotation guidelines and preselecting the best-performing annotators using a qualification task tseng2020best; iskender2020best; huang2023incorporating; zhang-etal-2023-needle.
+Our process of collecting annotations proceeded in two stages, following the setup of zhang-etal-2023-needle: (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations.
+See Appendix B.2 for more details on collecting annotations.
+For quality checks, we collect additional internal gold annotation (by the authors) for subsets of the data: $\mathcal{D}_{\texttt{dev}}$ for selecting the best prompt and $\mathcal{D}_{\texttt{iaa}}$ for validating the performance of human annotators (cf.
+Appendix B.2).
+Here is a complete overview of our data splits for D2T-Eval:
+  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – for LLM evaluation, annotated with crowdworkers,
+  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – for the study of prompt variants, annotated internally,
+  • $\mathcal{D}_{\texttt{iaa}}$ (12 outputs) – control for human crowdworkers, annotated internally.
+4.1.2 MT-Eval: Identifying Errors in Machine Translation
+For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (kocmi2024findings).
+The system outputs were annotated with the Error Span Annotation (ESA) protocol (kocmi2024error) by professional translators.
+The inputs $X$ for MT-Eval are the texts in the source language used to produce the translation $Y$.
+We follow the WMT 2024 annotation guidelines, focusing on character-level span annotations of Major and Minor translation errors (see Table˜6 for their definitions).
+Note that unlike for the other tasks, the annotations in MT-Eval cannot overlap and need not be aligned with word boundaries.
+We select the three textual domains present in the WMT 2024 shared task: news, literary, and social; using the data translated from English into other languages: Chinese, Czech, German, Hindi, Icelandic, Japanese, Russian, Spanish, and Ukrainian.
+The original dataset has nearly 50k model outputs, making it too extensive for our evaluation campaign.
+Therefore, we used a balanced sample of all available system outputs for 10 sampled inputs for each pair of languages, making up 2,854 examples in total.
+4.1.3 Propaganda: Propaganda Technique Detection
+For the Propaganda task, we use the dataset of dasanmartino2019finegrained containing news collected mostly from on-line propagandistic sources.
 The token-level annotations in the dataset created by expert annotators cover 18 categories of logical fallacies and persuasion techniques.
 We use the test split for our experiments.
-Note that there are no inputs $X$ for the task, as all the annotated categories are intrinsic to the evaluated text $Y$.
-4.4 Mitigating Data Contamination
-We took preventive measures to ensure that our results are not affected by data contamination Balloccu et al. (2024); Dong et al. (2024); Jiang et al. (2024b):
-  • For D2T-Eval, we follow the approach of Kasner & Dušek (2024) to download a new set of structured data from publicly available sources and generate new outputs.
-  • For MT-Eval, we use the WMT 2024 dataset that was released at the end of 2024, which is after the knowledge cut-off date of the LLMs we are using.
-  • For Propaganda, the dataset is publicly available since 2019, but due to its format cannot be easily memorized.
-[^9] The model would need to separately match the annotations for numerical token ids with the text.
+Inputs $X$ are empty for this task, as all annotated categories are intrinsic to the evaluated text $Y$.
 4.2 Collecting LLM annotations
 Models
-For our span annotator LLMs, we use a mix of open and proprietary state-of-the-art models: (1) instruction-tuned models Llama 3.3 70B Grattafiori et al.
-(2024), GPT-4o Hurst et al.
-(2024), Claude 3.7 Sonnet Anthropic (2025), and (2) reasoning models: DeepSeek-R1 70B DeepSeek-AI et al.
-(2025), o3-mini OpenAI (2025) and Gemini 2.0 Flash Thinking Deepmind (2025).
-We run the open models (Llama 3.3, DeepSeek-R1) locally in 4-bit quantization using ollama.
+For collecting span annotations with LLMs, we use a mixture of open and proprietary state-of-the-art models:
+  • instruction-tuned models: Llama 3.3 70B grattafiori2024llama, GPT-4o hurst2024gpt, and Claude 3.7 Sonnet anthropic2025claude,
+  • reasoning models: [^4] By reasoning models we understand the models that use extra inference time to generate a thinking trace before providing the answer marjanovic2025deepseek.
+DeepSeek-R1 distilled 70B version guo2025deepseek, o3-mini openai2025o3mini, and Gemini 2.0 Flash Thinking google2025geminithinking.
+See Appendix˜A for details on our experimental setup.
 Prompts
 We define several prompt variants for our experiments.
-$\mathcal{P}_{\texttt{base}}$ is the base prompt that includes the guidelines $\mathcal{G}$ as given to human annotators.
-By extending $\mathcal{P}_{\texttt{base}}$, we implement few-shot ($\mathcal{P}_{\texttt{5shot}}$) and chain-of-thought ($\mathcal{P}_{\texttt{cot}}$) prompts.
+$\mathcal{P}_{\texttt{base}}$ is the base prompt that includes the guidelines $\mathcal{G}$ as given to human annotators and asks the model to explain its annotation.
+By extending $\mathcal{P}_{\texttt{base}}$, we implement a few-shot prompt adding 5 examples ($\mathcal{P}_{\texttt{5shot}}$) and a chain-of-thought prompt simply asking the model to produce intermediate reasoning ($\mathcal{P}_{\texttt{cot}}$).
 We also ablate $\mathcal{P}_{\texttt{base}}$ by removing extended guidelines ($\mathcal{P}_{\texttt{noguide}}$) and not asking for explanations ($\mathcal{P}_{\texttt{noreason}}$).
-The full prompts can be found in Appendix D.
-Annotation Tool
-We use factgenie Kasner et al. (2024), a tool that supports collecting span annotations from humans via a web interface and from LLMs via API calls.
+The full prompts can be found in Appendix˜D.
 5 Results
-In this section, we describe the results of our experiments on span annotation with LLMs.
-We first investigate the effect of prompting techniques in Section 5.1.
-Next, we evaluate the LLM annotations using automatic metrics (Section 5.2) and manually analyze the errors in the model outputs (Section 5.3).
+We first investigate the effect of prompting techniques in Section˜5.1.
+Next, we evaluate the LLM annotations using automatic metrics (Section˜5.2) and manually analyze the errors in the model outputs (Section˜5.3).
 5.1 Prompting Techniques
-We perform a small-scale study on $\mathcal{D}_{\texttt{dev}}$ using open models (Llama 3.3 and DeepSeek-R1) to study the differences between prompting techniques. The results are shown in Table 2.
+We perform preliminary experiments on the D2T-Eval task $\mathcal{D}_{\texttt{dev}}$ set using open models (Llama 3.3 and DeepSeek-R1) to study the differences between prompting techniques.
+The results are shown in Table˜2.
 Llama 3.3 DeepSeek-R1
-Prompt $F_{1}$ $\gamma$a/o $F_{1}$ $\gamma$a/o
+Prompt $F_{1}$ $\gamma$ #a/o $F_{1}$ $\gamma$ #a/o
 $\mathcal{P}_{\texttt{base}}$ 0.20 0.13 2.4 0.25 0.20 1.0
+$\mathcal{P}_{\texttt{cot}}$ 0.09 0.10 0.8 0.24 0.19 1.1
+$\mathcal{P}_{\texttt{5shot}}$ 0.25 0.18 2.5 0.21 0.16 1.4
 $\mathcal{P}_{\texttt{noguide}}$ 0.11 0.08 3.4 0.20 0.16 1.6
 $\mathcal{P}_{\texttt{noreason}}$ 0.22 0.13 2.2 0.24 0.18 1.1
-$\mathcal{P}_{\texttt{cot}}$ 0.09 0.10 0.8 0.24 0.19 1.1
-$\mathcal{P}_{\texttt{5shot}}$ 0.25 0.18 2.5 0.21 0.16 1.4
 Table 2: Comparison of prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ (#a/o is the average number of annotations per output).
+Figure 2: Comparison of LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators.
+For $F_{1}$ score, the shadow bar denotes its soft variant.
+The dashed horizontal lines for D2T-Eval denote agreement between our human annotators.
+Detailed results are included in Tables˜11, 12, 13 and 14.
 Including detailed guidelines seems beneficial: omitting the guidelines ($\mathcal{P}_{\texttt{noguide}}$) lowers the performance of both models.
 In contrast, not letting the model explain the annotation ($\mathcal{P}_{\texttt{noreason}}$) does not have a substantial effect.
-For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce less annotations per example than the base variant (0.8 vs.
+For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce fewer annotations per example than the base variant (0.8 vs.
 2.4), leading to lower F1 and $\gamma$ scores.
-[^10] The model with $\mathcal{P}_{\texttt{cot}}$ tends to “overthink” the annotations, deciding not to annotate cases of errors against which it can find some arguments.
-We provide an example of this behavior in Table 27.
-Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.111111This observation is aligned with DeepSeek-AI et al.
-(2025), who note that few-shot prompting consistently degrades the performance of DeepSeek-R1.
-Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for our main experiments.
+Llama 3.3 with $\mathcal{P}_{\texttt{cot}}$ tends to “overthink” the annotations, deciding not to annotate cases of errors against which it can find some arguments.
+Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.
+This observation is aligned with guo2025deepseek, who note that few-shot prompting degrades the performance of DeepSeek-R1.
+Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for further experiments.
 5.2 LLM vs. Human Annotations
-Figure 2: Comparison of LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators.
-The stripe pattern denotes proprietary models, darker shades denote reasoning models.
-For $F_{1}$ score, the grey bar denotes its soft variant.
-The dashed horizontal lines for D2T-Eval denote agreement between our human annotators.
-Detailed results are included in Tables 8, 16 and 20.
-Next, we compute IAA between LLM and human annotations using the metrics described in Section 3.3.
-The overall results for all tasks are given in Figure 2.
-We also provide detailed results for individual tasks in Appendix E.
-Figure 3: Confusion matrix for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other, cf. Table 5).
-Model football gsmarena openweath.
-$F_{1}$ $\gamma$ $F_{1}$ $\gamma$ $F_{1}$ $\gamma$
-Llama 3.3 0.29 0.20 0.11 0.08 0.12 0.05
-GPT-4o 0.30 0.19 0.09 0.06 0.18 0.12
-Claude 3.7 0.38 0.25 0.16 0.14 0.27 0.28
-DeepS. R1 0.34 0.28 0.12 0.16 0.18 0.12
-o3-mini 0.42 0.33 0.15 0.19 0.29 0.31
-Gem. 2-FT 0.38 0.31 0.16 0.16 0.22 0.24
-Table 3: Performance of models by domain (football, gsmarena, openweather) in D2T-Eval: $F_{1}$ score (hard), $\gamma$ score. The best results are bolded.
+Next, we compare LLM and human annotations using the metrics described in Section˜3.3.
+The overall results for all tasks are given in Figure˜2.
+We provide detailed results for individual tasks in the Appendix F.
+Figure 3: Confusion matrix for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other), averaged across models (see Table˜5 for category descriptions).
 Reasoning models outperform instruction-tuned models.
 DeepSeek-R1 generally outperforms the same-sized Llama 3.3.
-The superiority of DeepSeek-R1 is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs.
+Its superiority is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs.
 0.18, $\gamma$ score of 0.19 vs.
 0.11).
-The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o despite o3-mini’s price per token being approximately 2x lower.
-A notable exception to this trend is the Claude 3.7 Sonnet, which is mostly on par with o3-mini and excels on MT-Eval.
-However, its price per token is approximately 3x higher (cf.
-Table 4 for model costs).
+The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o.
+A notable exception to this trend is the non-reasoning Claude 3.7 Sonnet, which scores mostly on par with o3-mini and excels at MT-Eval.
+Model Cost (\$/1k out) Time (s/out)
+crowdworkers 500 129.1
+Llama 3.3 - 21.6
+DeepSeek-R1 - 227.5
+Claude 3.7 Sonnet 10.5 9.0
+o3-mini 3.6 21.8
+Table 3: Estimate of costs and time requirements on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and proprietary models (Claude 3.7 Sonnet, o3-mini).
 LLMs reach human IAA on D2T-Eval, Propaganda is harder.
-For D2T-Eval, we can compare model results with an average IAA on examples annotated by two human annotators (cf.
-Section B.2).
+For D2T-Eval, we compare model results with an average IAA on a subset of examples annotated by two human annotators.
 Here, o3-mini, Claude 3.7 and Gemini 2.0 mostly reach or surpass human agreement.
-For Propaganda, the upper bound of IAA is the result of Da San Martino et al.
-(2019), who report $\gamma=0.31$ for annotators before consolidation.
-This score is substantially higher (the best LLM score being $\gamma=0.16$ for the Claude 3.7 Sonnet).
+For Propaganda, the upper bound of IAA is the result of dasanmartino2019finegrained, who report $\gamma=0.31$ for annotators before consolidation.
+This score is substantially higher than LLMs (the best LLM score being $\gamma=0.16$ for Claude 3.7 Sonnet).
 However, this task used expert annotators and has the largest number of categories.
-The latter manifests itself in the large difference between the soft and hard $F_{1}$ scores.
-[^12] The WMT24 dataset for MT-Eval contains examples annotated with a pair of annotators, but these examples form only a small fraction and exhibit high variance between language pairs.
-Therefore, we decided to omit the comparison with human IAA in MT-Eval.
+The latter property is reflected in the large difference between the soft and hard $F_{1}$ scores.
+[^5] We omit the comparison with human IAA in MT-Eval.
+While the WMT24 dataset for MT-Eval contains examples annotated with a pair of annotators, these examples take up only a small fraction and exhibit high variance between language pairs.
 Models confuse related categories.
-Confusion matrices (cf.
-Figure 3 and Section E.8) suggest that the models tend mainly to confuse related categories such as Contradictory, Not checkable and Misleading for D2T-Eval, which may be related to ambiguity or subjective understanding of category definitions.
-The models also use a less diverse distribution of categories, e.g., resorting to the Contradictory category more often than human annotators.
-LLMs are more efficient than human annotators.
-For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance model o3-mini LLM costs \$3.60 (cf.
-Table 4).
-The crowdworkers take 129.1 seconds per output on average, which is better than DeepSeek-R1 on our local infrastructure, but order of magnitude slower than API-based models.
-[^13] Also note that unlike LLMs, the crowdworkers are not required to produce explanations for the annotations.
-This suggests that LLMs can be an effective alternative to human annotators both in terms of costs and time.
-Metric crowdw. Llama 3.3 DeepSeek-R1 Claude 3.7 Sonnet o3-mini
-Cost per 1k outputs (\$) 500 - - 10.5 3.6
-Time per output (s) 129.1 21.6 227.5 9.0 21.8
-Table 4: Estimate of costs and time requirements on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and proprietary models (Claude 3.7 Sonnet, o3-mini).
-Note that we omit the infrastructure and energy costs for the models deployed locally.
+Confusion matrices (see Figure˜3 and Appendix˜F) suggest that the models tend mainly to confuse related categories, which may be related to ambiguity or subjective understanding of category definitions.
+The models also use a less diverse distribution of categories than human annotators.
+LLMs are more cost- and time-efficient than human annotators.
+An important factor when comparing LLMs and human annotators is efficiency with respect to cost and time per output.
+For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance o3-mini LLM costs \$3.60 (see Table˜3).
+In terms of time, the crowdworkers take 129.1 seconds per output on average, which is better than DeepSeek-R1 70B running on our local infrastructure, but an order of magnitude slower than the API-based models.
+[^6] Note that we do not ask the crowdworkers to give us a reason $r$ for the annotation, which would arguably make the responses of the crowdworkers slower.
+Therefore, LLMs are a more efficient alternative in terms of costs and time.
+Figure 4: Results of our manual analysis.
+We analyzed 18 annotations and their explanations for each model and task (216 annotations in total).
+The color bars show annotations that we classified as Correct, Partially correct, Wrong category, Incorrect, and Undecidable.
+Detailed results are provided in Tables˜19 and 20.
 5.3 Manual Analysis of LLM Annotations
-We manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda.
+To gain more insights into the qualitative aspects of LLM annotations, we manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda.
+[^7] We do not include MT-Eval in the manual analysis due to our insufficient expertise in most target languages.
 For each model, we sampled three annotations per category in D2T-Eval and one annotation per category in Propaganda.
-We show the results in Figure 4.
+Without access to the annotation source, we classified the annotations and their explanations as Correct, Partially correct, Wrong category, Incorrect, and Undecidable.
+We show the results in Figure˜4 and Tables˜19 and 20.
 In total, we marked 49.5% of LLM-generated annotations and 50.5% of reasons as correct (with 9.2% of annotations and 12.5% of reasons additionally marked as partially correct).
-Reasoning models perform better, with 56.4% of their annotations and 58.3% marked as completely correct.
+Reasoning models perform better, with 56.4% of their annotations and 58.3% reasons marked as completely correct.
 The most accurate annotations on D2T-Eval were those made by Gemini 2.0 and DeepSeek-R1.
-o3-mini performed well on both tasks, although Propaganda proved to be challenging for all models.
+o3-mini performed well on both tasks, although Propaganda proved challenging for all models.
 What are the sources of model errors?
-Qualitatively, we find that the models often select wrong error categories despite identifying real issues (e.g.
-labeling contradictory statements as "incoherent").
-Models also tend to be overly attentive, flagging noise in the data (quotes or non-article content) as errors, or marking slight numerical variations (such as rounded values) as misleading.
+We find that the models often select wrong error categories despite identifying real issues (e.g., labeling Contradictory statements as Incoherent).
+Models also tend to be overly attentive, flagging noise in the data (e.g., markup or off-topic content in Propaganda) as errors, or marking slight numerical variations (such as rounded values) as misleading.
 All of these cases could be tackled by more descriptive guidelines or additional examples.
 However, in some cases, the models also misread or misinterpreted the data (e.g., claiming wind speed measurements do not exist when they do), which hints at deeper issues with understanding the data.
-Incorrect explanations vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., counting omissions as sources of errors) to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors.
+Incorrect explanations vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., appealing to facts that are “missing” from the text) to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors.
 Occasionally, the model admits that it marked a correct span as an error, such as in “The description of the game’s duration aligns with the data, providing coherent information”.
-[^14] This typically happens to GPT-4o, even though OpenAI API ensures JSON key ordering so the explanation should have been generated before the annotation (cf.
+[^8] This typically happened to GPT-4o, even though OpenAI API ensures JSON key ordering so the explanation should have been generated before the annotation (cf.
 Section 3.2).
 How good are human annotations?
-Interestingly, the LLM annotations that were marked as correct have only 24% hard character-level overlap (51% soft) with human annotations.
-This fact led us to analyze the quality of human annotations in D2T-Eval.
-On a sample of 108 human annotations, we marked 45.3% of the annotations and 54.6% of the explanations correct, which is comparable to the LLM annotations.
-This suggests that the task is hard even for human annotators, and the quality of annotations from crowdworkers varies, even if they are preselected using a qualification task.
-Figure 4: Results of our manual analysis.
-We analyzed 18 annotations and their explanations for each model and task (216 annotations in total).
-The color bars show annotations that we classified as correct, partially correct, assigned wrong category, incorrect, and undecidable cases.
-Detailed results are provided in Tables 22 and 23 in Appendix E.
+Concerningly, the LLM annotations that were marked as correct have only 24% hard character-level overlap (51% soft) with human annotations.
+This fact led us to analyze the quality of human annotations in D2T-Eval (the task in which we had the necessary domain expertise).
+Using the same methodology as we used for the LLM annotations, we annotated a limited sample of 108 human annotations.
+We marked 45.3% of the annotations as Correct, which is comparable to the LLM annotations (see Table˜4 for the results).
+These findings suggest that the task is hard even for human annotators, and the quality of annotations from crowdworkers varies, even if they are preselected using a qualification task.
 6 Discussion
+Here, we summarize our findings and discuss the implications of our results.
 Can LLMs substitute human annotators?
-The IAA between LLMs and human annotators is generally moderate.
-However, we also need to take into account a realistic upper bound: the average IAA between human annotators themselves.
-In that case, the strongest LLMs have already reached the level of skilled crowdworkers in D2T-Eval.
-As we also showed, not all human annotations are “gold”, as even qualified crowdworkers make similar amounts of errors as LLMs.
-If we also factor in financial aspects, LLMs provide better flexibility, shorter response times, and lower costs.
-In light of this, we argue that the advantages of automating span annotation with LLMs outweigh its issues.
-LLM-based span annotation will also benefit from future increases in LLM capabilities.
-Recommendations
-To use an LLM as a span annotator, we strongly recommend providing detailed guidelines describing how to handle all edge cases.
-In contrast, we note that providing specific examples (i.e., a few-shot setup) does not bring consistent improvements, possibly due to the length and complexity of examples for this task.
-When using LLMs in a new setup, we recommend validating model’s annotations against examples hand-annotated by experts on a sample of the data.
-Regarding model choice, we recommend choosing a reasoning model as opposed to an equivalent model without reasoning, since the reasoning models provide more reliable annotations at the cost of slightly higher response times and token count.
-Limitations
-Span annotation as an evaluation method is not suitable for certain NLG evaluation tasks, such as annotating omissions or rating the overall text style.
-Although we tried to experiment with a representative sample of models, prompts, and tasks, our choice is necessarily constrained by our limited time frame and budget.
-Our estimates of the upper-bound IAA for each task are difficult to establish and depend on many factors, such as the chosen annotation categories, their inherent ambiguity, the annotation guidelines, or the qualification level of human annotators.
+The IAA between LLMs and human annotators is only moderate, suggesting LLMs cannot straightforwardly replace human annotators.
+However, using LLMs may be a reasonable option in scenarios such as D2T-Eval based on crowdworkers, where the strongest LLMs reach the average IAA between human annotators themselves.
+In other cases, when deciding whether to employ LLMs as span annotators, one needs to balance desired output quality with other practical aspects.
+Here, LLMs provide better flexibility, shorter response times, and lower costs.
+One should also consider the quality of available human annotators, as even qualified crowdworkers make similar amounts of errors as LLMs.
+It can be also assumed that LLM-based span annotation will benefit from future increases in LLM capabilities, while crowdworkers may increasingly rely on LLM to complete tasks veselovsky2023artificial.
+A promising solution seems to be a hybrid approach in which LLMs pre-annotate the text and humans post-edit the annotations zouhar2024aiassisted.
+How to deploy LLMs as span annotators?
+We recommend providing LLMs with detailed guidelines that describe conventions and how to handle ambiguous cases (cf.
+Figure˜5).
+In contrast, we do not recommend providing specific examples (cf.
+Figure˜8), as this approach did not bring consistent improvements.
+Arguably, this is due to the length and complexity of the examples, making them distracting to the model.
+When using LLMs with custom categories or guidelines, we recommend validating the model’s annotations against examples hand-annotated by experts on a sample of the data.
+In general, reasoning models tend to provide more reliable annotations at the cost of higher response times and token count.
+Annotations
+Source C P W I U
+Human annotators 49 4 17 31 7
+Table 4: Manual evaluation results for human annotators on D2T-Eval. Categories: C=Correct, P=Partially correct, W=Wrong category, I=Incorrect, U=Undecidable.
 7 Conclusion
-We demonstrated that LLMs can serve as span annotators for three span annotation tasks: evaluating data-to-text generation, identifying errors in machine translation, and detecting propaganda in human-written texts.
+We showed that LLMs can serve as span annotators for three span annotation tasks: evaluating data-to-text generation, identifying errors in machine translation, and detecting propaganda in human-written texts.
 Our experiments show that LLMs achieve moderate agreement with skilled human annotators.
@@
 Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions.
-Overall, automating span annotation with LLMs offers a promising alternative to human annotation, opening the way for more fine-grained NLG evaluation methods.
+Automating span annotation with LLMs seems to be a promising alternative to fine-grained human evaluation sourced from crowdsourcing platforms, opening the way towards scalable and actionable automatic NLG evaluation methods.
+Limitations
+Although we aimed to select a representative sample of models, prompts and tasks, our choice is constrained by our limited time frame and budget.
+Our estimates of the upper-bound IAA for each task are difficult to establish and depend on many factors, such as the chosen annotation categories, their ambiguity, the annotation guidelines, or the qualification level of human annotators.
+The estimates are also not readily available for existing datasets and require additional data collection.
+As an evaluation method, span annotation is not well-suited for certain NLG evaluation tasks such as annotating omissions or rating the overall text style.
+In these cases, it is best to combine span annotation with other evaluation methods.
 Author Contributions
@@
 Acknowledgments
-This work was funded by the European Union (ERC, NG-NLG, 101039303) and Charles University project SVV 260 698.
+This work was funded by the European Union (ERC, NG-NLG, 101039303) and the National Recovery Plan funded project MPO 60273/24/21300/21000 CEDMO 2.0 NPO.
 It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No.
@@
 Annotators were pre-selected on the basis of their primary language (English).
-All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research (see Appendix B.2 for details).
+All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research.
 The domains were selected so that they do not contain sensitive or potentially offensive content.
@@
 A.1 Open Models
-We run the local models using the ollama framework in 4-bit quantization. Specifically, we use the following local models for span annotations:
-We also use the following models for generating texts in D2T-Eval:
+We run the local models using the ollama framework in 4-bit quantization.
+Specifically, we use llama3.3:70b and deepseek-r1:70b for span annotations.
+We also use gemma2:2b and phi3.5:3.8b for generating texts in D2T-Eval.
 For better reproducibility, we set the seed to 42 and the temperature to 0 for the local models. We do not use these parameters for proprietary models as these parameters are generally not supported.
-We run the models on a set of GPUs, including NVIDIA H100 NVL (95G), AMD MI210 (64G), and NVIDIA RTX 3090 (24G).
+We run the models using several GPU variants, including NVIDIA H100 NVL (95G), AMD MI210 (64G), and NVIDIA RTX 3090 (24G).
 A.2 Proprietary Models
@@
   • Gemini 2.0 Flash Thinking: gemini-2.0-flash-thinking-exp-01-21
-A.3 Parsing the Output with Reasoning Traces
-At the time we performed our experiments, some of the reasoning models either did not support structured output (Gemini 2.0 Flash Thinking) or it would cancel out their ability to perform reasoning (ollama models [^15] https://github.com/ollama/ollama/issues/8529).
-Therefore, we use the following heuristic: we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response.
-Appendix B D2T-Eval Details
-In this section, we provide more details on the process of collecting data (Section B.1) and annotations (Section B.2) for D2T-Eval.
-B.1 Collecting Data
-To collect a novel dataset for D2T-Eval that is not subject to data contamination (cf.
-Section 4.4), we follow the approach of Kasner & Dušek (2024) to collect structured data from publicly available APIs.
-Domains
-The Quintd tool Kasner & Dušek (2024) enables downloading data for five domains: openweather, gsmarena, ice_hockey, our_world_in_data, and wikidata.
-We selected openweather and gsmarena as the most suitable for the data collection process.
-Additionally, we collect the new football domain from RapidAPI - API-Football.
-[^16] Its structured data contain more detailed information about the game than ice_hockey and the sport itself is also more generally well-known.
-Generating model outputs.
-To obtain output texts, we prompt two state-of-the-art models, Llama 3.3 Grattafiori et al.
-(2024) and GPT-4o Hurst et al.
-(2024), and two smaller models, Gemma 2 Team et al.
-(2024) and Phi-3.5 Abdin et al.
-(2024) (cf.
-Section A.1).
-We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences.
-We provide the prompts in Appendix D.
-Note on annotating models’ own outputs.
-Two of the models (Llama 3.3 and GPT-4o) are also used later to collect annotations, which can reveal self-preference bias Zheng et al.
-(2023); Wataoka et al.
-(2024).
-Since we do not rank the quality of generated outputs, we do not consider it a flaw but rather a ground for future investigation.
-Moreover, we hypothesize that since span annotations are more grounded than direct assessment, the self-preference bias will not be as prominent.
+A.3 Web Interface
+We implement our span annotation process using factgenie kasner2024factgenie: a tool that supports both collecting span annotations from humans via a web interface and from LLMs via API calls.
+Figure˜12 shows samples of our annotation interface implemented in factgenie for human annotators, including data visualizations from the football and openweather domains.
+Appendix B Annotating D2T-Eval
+B.1 Generating Outputs
+For generating the outputs for the structured inputs we collected, we use two larger models – Llama 3.3 70B grattafiori2024llama and GPT-4o hurst2024gpt – and two smaller models – Gemma 2 2B team2024gemma and Phi-3.5 3.8B abdin2024phi.
+See more details on the models in Appendix˜A and prompts in Appendix˜D.
 B.2 Collecting annotations
-Here we describe our process of writing annotation guidelines and selecting crowdworkers in order to ensure high quality of the annotations collected for D2T-Eval.
 Annotation guidelines
-For the annotation guidelines, we (the authors of this paper) went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$.
-We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works Kasner & Dušek (2024); Thomson & Reiter (2020).
-Eventually, we settled on the list of annotation categories described in Appendix C (Table 5).
-Our list is an extended and updated version of the error taxonomy of Kasner & Dušek (2024), also covering coherence categories (Incoherent, Repetitive) on top of semantic accuracy.
-Additionally, we agreed upon a list of general recommendations for collecting the annotations:
-  • The annotator should always annotate the longest continuous span (i.e., the whole fact instead of a single word).
-  • The annotator should annotate according to their own knowledge and only the spans that they are sure about.
-  • The annotator should ignore subjective statements such as "a lightweight smartphone".
-  • The outputs in openweather can contain both precise numbers (e.g. 10.71°C) and the rounded ones (e.g. 11°C) as long as they agree with the data.
-We provide these additional recommendations in $\mathcal{P}_{\texttt{base}}$, while we omit them in $\mathcal{P}_{\texttt{noguide}}$ (cf. Figures 6 and 7).
+For the annotation guidelines, we went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$.
+We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works kasner2024traditional; thomson2020gold.
+We settled on the following annotation categories (see Table˜5 for details): semantic accuracy errors due to information Contradictory to the input, Not Checkable, or Misleading; any Incoherent and Repetitive content, and any Other errors.
 Gold annotations
-With the annotation guidelines established, we proceeded to collect our own internally annotated gold data.
-The purpose of this data is two-fold: (1) to pre-select skilled crowdworkers, and (2) to create a smaller high-quality development set for the model prompting study.
-Since we also wanted to quantify our own internal IAA, we divided the data into two sets: $\mathcal{D}_{\texttt{dev}}$, containing 84 examples annotated individually by 7 annotators (12 examples per annotator), and $\mathcal{D}_{\texttt{iaa}}$, containing 12 examples annotated commonly by each annotator.
+With the annotation guidelines established, we proceeded to collect our own internally annotated gold data: $\mathcal{D}_{\texttt{dev}}$, which contains 84 examples annotated individually by one of 7 annotators (12 examples per annotator) and $\mathcal{D}_{\texttt{iaa}}$, which contains 12 examples annotated commonly by all annotators.
+[^9] We selected an example for each of the 4 domains and 3 models.
+The purpose of $\mathcal{D}_{\texttt{dev}}$ is to create a high-quality development set for the model prompting study, while the purpose of $\mathcal{D}_{\texttt{iaa}}$ is to pre-select skilled crowdworkers and quantify the performance of crowdworkers during data collection.
 Our average IAA on $\mathcal{D}_{\texttt{iaa}}$ was $F_{1}=0.444$ and $\gamma=0.399$.
-Qualification task
-To gather a pool of skilled annotators, we set up a qualification task on Prolific.com.
-For the study, we pre-selected workers whose first language is English, with >99% approval rate and more than 100 submissions.
-The workers were first presented with a detailed tutorial with annotation guidelines and examples of individual errors.
-After the tutorial, we tested the worker performance on five manually pre-selected examples from $\mathcal{D}_{\texttt{dev}}$.
-For annotating the data in $\mathcal{D}_{\texttt{test}}$, we selected workers that had $F_{1}$ score higher than 0.5 w.r.t.
-our ground-truth solution to the qualification task.
-Of the 230 annotators who participated in the qualification task, 50 annotators (21.7%) passed this baseline.
-We pay all the annotators an average reward of 9.58 GBP / hour, regardless of the qualification outcome.
-Annotating test data
-We invited annotators who passed the qualification task to annotate $\mathcal{D}_{\texttt{test}}$.
-Out of 50 selected participants, 45 participated in the main task (=90% turnover rate).
-We presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{dev}}$ (that is, the examples from $\mathcal{D}_{\texttt{dev}}$ that we did not use for the qualification task).
-All 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator.
+Crowdsourcing annotations
+We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com.
+Our process of collecting annotations proceeded in two stages, following the setup of zhang-etal-2023-needle: (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations.
+  • Qualification task: For the qualification task, we pre-selected workers whose first language is English, with >99% approval rate and more than 100 submissions.
+The workers were presented with a detailed tutorial with annotation guidelines and examples of individual errors.
+After the tutorial, we tested the worker performance on five manually pre-selected examples from $\mathcal{D}_{\texttt{iaa}}$.
+We invited annotators with the $F_{1}$ score higher than 0.5 w.r.t.
+our internal annotations for the main task.
+  • Main task: Of the 230 annotators who participated in the qualification task, 50 annotators (21.7%) qualified.
+Of these, 45 annotators accepted (=90% turnover rate).
+For annotating the data in $\mathcal{D}_{\texttt{test}}$, we presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{iaa}}$ (that is, the examples that we did not use for the qualification task).
+All the 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator.
 Furthermore, 475 outputs (39.6%) were annotated by an additional annotator.
-We payed all the annotators an average reward of 10.70 GBP / hour.
-We use examples with two annotators to compute the average IAA for D2T-Eval in Section 5.2.171717For other experiments, we use only the outputs from the first annotator as the reference data.
-B.3 Annotation Interface
-We implement our annotation interface using factgenie. Figure 5 shows the main annotation interface, including an example from the openweather domain.
-Figure 5: Screenshot of the factgenie annotation interface used for collecting span annotations.
+[^10] We use examples with two annotators to compute the average IAA for D2T-Eval in Section 5.2.
+For other experiments, we use only the outputs from the first annotator as reference data.
+For the qualification task, we paid all the annotators an average reward of 9.58 GBP / hour regardless of the qualification outcome.
+For the main task, we pay all the annotators an average reward of 10.70 GBP / hour.
 Appendix C Annotation Categories
-Tables 5, 6 and 7 show an overview of the annotation span categories that we used for our tasks along with their descriptions.
+Tables˜5, 6 and 7 show an overview of the annotation span categories that we used for our tasks along with their descriptions.
 Category Name Description
@@
 Whataboutism A technique that attempts to discredit an opponent“s position by charging them with hypocrisy without directly disproving their argument
-Table 7: Annotation categories for the MT-Eval task. The categories are adopted from Da San Martino et al. (2019).
+Table 7: Annotation categories for the Propaganda task. The categories are adopted from dasanmartino2019finegrained.
 Appendix D Prompts
-Figures 6, 7, 8 and 9 show the prompts for the D2T-Eval that we use for the experiments in Section 5.1.
-Figure 10 shows the base prompt we used for MT-Eval, and Figure 11 shows the base prompt we used for Propaganda.
-Finally, Figure 12 shows the prompt we used for generating the outputs for D2T-Eval.
+Here, we provide the model prompts:
+  • Figures˜5, 6, 7 and 8 show the prompts for the D2T-Eval that we use for the experiments in Section˜5.1.
+  • Figure˜9 shows the base prompt we used for MT-Eval.
+  • Figure˜10 shows the base prompt we used for Propaganda.
+  • Figure˜11 shows the prompt we used for generating the outputs for D2T-Eval.
 Your task is to identify errors in the text and classify them.
@@
 For more robust parsing, we initialize the model response with ’Sure, here is the summary: "’ .
-Appendix E Results
-In this section, we provide detailed results of our experiments.
-The results for D2T-Eval are included in Sections E.1, E.2, E.3 and E.4:
-The results for MT-Eval are included in Sections E.5 and E.6:
-The results for Propaganda are included in Section E.7: Tables 20 and 21.
-The confusion matrices are included in Section E.8: Figures 13, 14 and 15.
-The manual evaluation results are included in Section E.9: Tables 22 and 23.
-E.1 D2T-Eval – Main Results
+(a) Data visualization – openweather
+(b) Interface for highlighting spans
+Figure 12: Samples from the factgenie annotation interface used for collecting span annotations.
+Appendix E Examples
+In Tables˜8, 9 and 10, we show examples of the annotated outputs for our tasks. Figure˜12 shows our annotation interface.
+Source (part of the visualization) [Uncaptioned image]
+Annotated by Llama 3 Bear Creek will experience mostly overcast skies for the next five days, with a chance of occasional showers on Wednesday and Thursdaynot check..
+Temperatures will remain relatively cool, hovering around freezing point or slightly below.
+Winds will be light to moderate throughout the period, with some gusts expected on Fridaynot check..
+The overall weather pattern suggests a generally cloudy and potentially wet week.
+Annotated by o3-mini Bear Creek will experience mostly overcast skies for the next five days, with a chance of occasional showers on Wednesday and Thursdaycont..
+Temperatures will remain relatively cool, hovering around freezing point or slightly below.
+Winds will be light to moderate throughout the period, with some gusts expected on Fridaycont..
+The overall weather pattern suggests a generally cloudy and potentially wet weekcont..
+Annotated by human Bear Creek will experience mostly overcast skies for the next five days, with a chance of occasional showers on Wednesday and Thursdaycont..
+Temperatures will remain relatively cool, hovering around freezing point or slightly below.
+Winds will be light to moderate throughout the period, with some gusts expected on Fridaymislead..
+The overall weather pattern suggests a generally cloudy and potentiallycont.
+wetcont.
+weekcont..
+Table 8: Example for span annotation of D2T-Eval in the weather domain with contradictorycont, misleadingmislead., and not-checkeablenot check.
+as error span categories.
+In the annotation interface, the visualization is interactive.
+The top part (not visible here) shows the place name and date timeline.
+Source “It’s your birthday soon, isn’t it?” Ivory asked, remembering that the princess’ birthday was in a few days. Kari nodded, a sad glint in her light blue eyes.
+Annotated by Llama 3 “¿Es su cumpleaños pronto, no lo es?”major Ivory le preguntó,minor recordando que el cumpleaños de la princesa era en unos días.
+Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
+Annotated by o3-mini “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días.
+Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
+Annotated by human “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días.
+Kari nodómajor, un deslumbramiento triste en sus ojos azules claros.
+[…]
+Annotated by Llama 3 When the left made Linda Sarsour into its role model, it climbed into bed withfear the worst of the worstlabelling.
+The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundloaded to commit school shootings, prosecutors alleged in court documents Wednesday.
+Annotated by o3-mini When the left made Linda Sarsour into its role model, it climbed into bed with the worst of the worst.loaded The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundloaded to commit school shootings, prosecutors alleged in court documents Wednesday.
+Annotated by human When the left made Linda Sarsour into its role modellabelling, it climbed into bedloaded with the worst of the worst.exag.
+The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundlabelling to commit school shootings, prosecutors alleged in court documents Wednesday.
+Table 10: Two examples for span annotation of Propaganda outputs with appeal-to-fearfear, name-calling-labellinglabelling, loaded-languageloaded, and exaggerationexag. as span categories.
+Appendix F Results
+Here, we provide detailed results of our experiments:
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
@@
 Gem. 2-FT 0.458 0.293 0.488 0.263 0.438 0.277 0.462 0.185 0.209 0.612
-Table 8: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. See Figure 2 for visualizaton of this table.
-Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
-Human 2981 2.5 28.8 50.3
-Llama 3.3 3214 2.7 7.4 65.5
-GPT-4o 2284 1.9 4.8 66.3
-Claude 3.7 2865 2.4 22.5 57.2
-DeepS. R1 1387 1.2 44.2 56.8
-o3-mini 1836 1.5 35.6 58.0
-Gem. 2-FT 2517 2.1 28.9 54.3
-Table 9: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval.
-Ann=# of annotations, Ann/Ex=annotations per example.
-w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-E.2 D2T-Evalfootball domain
-Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
-Hard Soft Hard Soft Hard Soft $\Delta$
-Llama 3.3 0.512 0.288 0.484 0.296 0.498 0.292 0.491 0.199 0.197 0.455
-GPT-4o 0.422 0.347 0.502 0.268 0.388 0.302 0.438 0.136 0.190 0.427
-Claude 3.7 0.612 0.360 0.511 0.391 0.555 0.375 0.532 0.157 0.254 0.587
-DeepS. R1 0.561 0.431 0.670 0.275 0.428 0.336 0.522 0.186 0.276 0.659
-o3-mini 0.610 0.470 0.608 0.384 0.496 0.422 0.547 0.125 0.328 0.645
-Gem. 2-FT 0.588 0.419 0.617 0.347 0.512 0.379 0.560 0.181 0.307 0.647
-Table 10: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the football domain of D2T-Eval.
-Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
-Human 1053 2.6 26.8 54.1
-Llama 3.3 977 2.4 6.2 71.5
-GPT-4o 761 1.9 2.8 72.8
-Claude 3.7 1051 2.6 13.8 63.1
-DeepS. R1 486 1.2 41.2 67.1
-o3-mini 642 1.6 30.5 70.8
-Gem. 2-FT 755 1.9 29 61.3
-Table 11: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the football domain of D2T-Eval.
-Ann=# of annotations, Ann/Ex=annotations per example.
-w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-E.3 D2T-Evalgsmarena domain
-Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
-Hard Soft Hard Soft Hard Soft $\Delta$
-Llama 3.3 0.165 0.107 0.247 0.122 0.282 0.114 0.263 0.149 0.077 0.456
-GPT-4o 0.127 0.107 0.269 0.082 0.205 0.093 0.233 0.140 0.062 0.453
-Claude 3.7 0.214 0.176 0.348 0.152 0.300 0.163 0.322 0.159 0.140 0.656
-DeepS. R1 0.223 0.199 0.405 0.089 0.180 0.123 0.249 0.126 0.160 0.686
-o3-mini 0.172 0.225 0.473 0.115 0.241 0.152 0.320 0.168 0.192 0.689
-Gem. 2-FT 0.234 0.218 0.426 0.128 0.249 0.161 0.314 0.153 0.161 0.675
-Table 12: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the gsmarena domain of D2T-Eval.
-Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
-Human 809 2.0 40.2 42.5
-Llama 3.3 890 2.2 15 73.9
-GPT-4o 670 1.7 10.5 59.2
-Claude 3.7 584 1.5 47.2 48.9
-DeepS. R1 208 0.5 69.2 53.8
-o3-mini 272 0.7 66.2 49.1
-Gem. 2-FT 419 1.1 54.2 45.1
-Table 13: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the gsmarena domain of D2T-Eval.
-Ann=# of annotations, Ann/Ex=annotations per example.
-w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-E.4 D2T-Evalopenweather domain
-Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
-Hard Soft Hard Soft Hard Soft $\Delta$
-Llama 3.3 0.209 0.112 0.327 0.118 0.344 0.115 0.335 0.220 0.053 0.272
-GPT-4o 0.404 0.196 0.354 0.160 0.289 0.176 0.318 0.142 0.122 0.377
-Claude 3.7 0.550 0.274 0.411 0.292 0.437 0.283 0.424 0.141 0.189 0.429
-DeepS. R1 0.454 0.241 0.436 0.143 0.258 0.180 0.324 0.144 0.124 0.533
-o3-mini 0.552 0.362 0.496 0.267 0.366 0.307 0.422 0.115 0.259 0.481
-Gem. 2-FT 0.458 0.226 0.415 0.249 0.458 0.237 0.435 0.198 0.155 0.368
-Table 14: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the openweather domain of D2T-Eval.
-Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
-Human 1119 2.8 19.2 52.2
-Llama 3.3 1347 3.4 1 55.7
-GPT-4o 853 2.1 1.2 66.2
-Claude 3.7 1230 3.1 6.5 56.1
-DeepS. R1 693 1.7 22.2 50.5
-o3-mini 922 2.3 10 51.8
-Gem. 2-FT 1343 3.4 3.5 53.2
-Table 15: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the openweather domain of D2T-Eval.
-Ann=# of annotations, Ann/Ex=annotations per example.
-w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-E.5 MT-Eval – Main Results
+Table 11: Evaluation of human and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. See Figure˜2 for visualizaton of this table.
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
@@
 Gem. 2-FT 0.257 0.184 0.312 0.180 0.339 0.173 0.304 0.130 0.066 0.710
-Table 16: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages.
-Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
-Human 2090 0.7 66.0 14.5
-Llama 3.3 6361 2.3 6.2 17.4
-GPT-4o 4866 1.7 7.0 15.9
-Claude 3.7 3782 1.4 30.6 15.9
-DeepS. R1 2586 0.9 36.3 15.1
-o3-mini 3039 1.1 35.8 13.8
-Gem. 2-FT 2181 0.8 50.0 15.2
-Table 17: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages.
-Ann=# of annotations, Ann/Ex=annotations per example.
-w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-E.6 MT-Eval – By Language
+Table 12: Evaluation of human and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages.
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
@@
 en-zh 0.163 0.196 0.346 0.163 0.294 0.169 0.302 0.133 0.075 0.623
-Table 18: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language (average across models).
-Language Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
-en-cs Model avg. 600 1.4 27.0 16.6
-Human 399 0.7 66.1 13.0
-en-es Model avg. 417 1.1 38.9 18.8
-Human 248 0.6 70.3 10.3
-en-hi Model avg. 396 1.3 26.2 19.0
-Human 222 0.5 71.2 10.7
-en-is Model avg. 563 1.9 14.3 15.7
-Human 752 2.5 18.3 16.6
-en-ja Model avg. 471 1.3 28.7 11.1
-Human 118 0.2 87.5 14.8
-en-ru Model avg. 500 1.3 25.9 18.2
-Human 287 0.7 58.7 19.4
-en-uk Model avg. 436 1.5 25.4 17.8
-Human 208 0.7 64.3 12.3
-en-zh Model avg. 420 1.2 34.6 7.2
-Human 171 0.2 85.1 6.6
-Table 19: Statistics of models (averaged) and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language.
-Ann=# of annotations, Ann/Ex=annotations per example.
-Ex.
-w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-E.7 Propaganda – Main Results
+Table 13: Evaluation of human and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language (average across models).
 Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
@@
 Gem. 2-FT 0.560 0.106 0.268 0.190 0.477 0.136 0.343 0.207 0.202 0.493
-Table 20: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the Propaganda.
-Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
+Table 14: Evaluation of human and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the Propaganda.
+Annotator Ann Ann/Ex w/o% Char/Ann
+Human 2981 2.5 28.8 50.3
+Llama 3.3 3214 2.7 7.4 65.5
+GPT-4o 2284 1.9 4.8 66.3
+Claude 3.7 2865 2.4 22.5 57.2
+DeepS. R1 1387 1.2 44.2 56.8
+o3-mini 1836 1.5 35.6 58.0
+Gem. 2-FT 2517 2.1 28.9 54.3
+Table 15: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval.
+Ann=# of annotations, Ann/Ex=ann.
+per example.
+w/o=% ex.
+without annotations, Char/Ann=# chars per ann.
+Annotator Ann. Ann/Ex w/o% Char/Ann
+Human 2090 0.7 66.0 14.5
+Llama 3.3 6361 2.3 6.2 17.4
+GPT-4o 4866 1.7 7.0 15.9
+Claude 3.7 3782 1.4 30.6 15.9
+DeepS. R1 2586 0.9 36.3 15.1
+o3-mini 3039 1.1 35.8 13.8
+Gem. 2-FT 2181 0.8 50.0 15.2
+Table 16: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on MT-Eval. See Table˜15 for the legend.
+Lang. Annot. Ann. Ann/Ex w/o% Char/Ann
+en-cs Model 600 1.4 27.0 16.6
+Human 399 0.7 66.1 13.0
+en-es Model 417 1.1 38.9 18.8
+Human 248 0.6 70.3 10.3
+en-hi Model 396 1.3 26.2 19.0
+Human 222 0.5 71.2 10.7
+en-is Model 563 1.9 14.3 15.7
+Human 752 2.5 18.3 16.6
+en-ja Model 471 1.3 28.7 11.1
+Human 118 0.2 87.5 14.8
+en-ru Model 500 1.3 25.9 18.2
+Human 287 0.7 58.7 19.4
+en-uk Model 436 1.5 25.4 17.8
+Human 208 0.7 64.3 12.3
+en-zh Model 420 1.2 34.6 7.2
+Human 171 0.2 85.1 6.6
+Table 17: Statistics of models (averaged) and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language. See Table˜15 for the legend.
+Annotator Ann. Ann/Ex w/o% Char/Ann
 Human 1439 14.2 4.0 40.2
@@
 Gem. 2-FT 1864 18.5 3.0 54.1
-Table 21: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the Propaganda.
-Ann=# of annotations, Ann/Ex=annotations per example.
-w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
-E.8 Confusion Matrices
-(a) Llama 3.3
-(b) DeepSeek-R1
-(c) GPT-4o
-(d) o3-mini
-(e) Gemini 2.0 Flash Thinking
-(f) Claude 3.7 Sonnet
-Figure 13: Confusion matrices comparing human annotations (rows) with model predictions (columns) for D2T-Eval.
-Categories: Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other (cf.
-Table 5).
-For the model average, see Figure 3.
-(a) Llama 3.3
-(b) DeepSeek-R1
-(c) GPT-4o
-(d) o3-mini
-(e) Gemini 2.0 Flash Thinking
-(f) Claude 3.7 Sonnet
-Figure 14: Confusion matrices comparing human annotations (rows) with model predictions (columns) for MT-Eval. Categories: Major and Minor errors (cf. Table 6).
-Figure 15: Confusion matrix of category counts averaged across models for Propaganda (see Table 7 for the description of categories).
-E.9 Manual Analysis
-Annotation Reason
+Table 18: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the Propaganda. See Table˜15 for the legend.
+Annotations Explanations
 Model C P W I U C P W I U
@@
 Table 19: Manual evaluation results for D2T-Eval domain. Categories for annotation and reason: C=Correct, P=Partially correct, W=Wrong category, I=Incorrect, U=Undecidable.
-Annotation Reason
+Annotations Explanations
 Model C P W I U C P W I U
@@
 Total 55 6 14 31 2 56 9 9 32 2
-Table 23: Manual evaluation results for Propaganda domain. Categories for annotation and reason: C=Correct, P=Partially correct, W=Wrong category, I=Incorrect, U=Undecidable.
-Appendix F Related Work (extended)
-This section provides more discussion on related work following Section 2.
-Due to the complexity and often open-ended nature of generated text, NLG evaluation is inherently difficult (Celikyilmaz et al., 2020).
-Machine translation (MT) has historically played an important role in the development of both automated evaluation metrics and human annotation protocols.
-Although evaluation practices are less standardized in other tasks, human evaluation typically involves ranking outputs or assigning numeric scores for specific aspects such as fluency or faithfulness (Novikova et al., 2018; van der Lee et al., 2021; Kocmi et al., 2022).
-Error Span Annotation.
-As simple numeric scores can be subjective and unreliable, the field is shifting towards annotating erroneous spans.
-The MT community has adopted a detailed error taxonomy, known as MQM (Lommel et al., 2014; Mariana, 2014, Multidimensional Quality Metrics).
-In MQM, each erroneous span in the translation has an assigned category, such as accuracy/terminology and severity, such as minor/major/critical.
-The final translation score is defined as a weighted sum of the error spans based on the error categories and severities.
-However, this leads to many ties and is problematic when judging longer translations.
-In addition, this protocol requires a very expensive expert human labor.
-For these reasons, Kocmi et al.
-(2024c), based on prior work by Popovic (2020), introduce ESA (Error Span Annotation) for machine translation evaluation.
-This protocol drops the error categorization, which is primarily useful for diagnosis and not evaluation, simplifies severities, and has the annotators also assign the final score after they have done the span annotations.
-To make large-scale annotation more efficient, Zouhar et al.
-(2025) use a quality estimation system that pre-highlights potential error spans in the translation.
-The annotators then “post-edit” the error annotations and assign the final score.
-In data-to-text, Thomson & Reiter (2020) introduce an evaluation protocol for high-quality annotation of semantic accuracy errors in generated basketball reports.
-Similarly to MQM, such an approach involves significant costs and time requirements, as the authors report.
-Beyond NLG evaluation, span annotation has also been applied in other areas, such as fine-grained propaganda detection.
-Da San Martino et al.
-(2019) build an annotated corpus with span annotations for 18 different propaganda techniques, followed by Hasanain et al.
-(2024) and Ramponi et al.
-(2025) who apply similar approach for propaganda detection in Arabic and Italian, respectively.
-Automated Metrics.
-Although human evaluation is considered the gold standard for assessing quality in NLG tasks, its substantial demands in terms of time and cost make it less scalable and practical.
-This has led to an overreliance on traditional automatic metrics (Schmidtová et al., 2024), which typically assess text quality by measuring lexical overlap with human-written reference texts (e.g., Papineni et al., 2002; Lin, 2004).
-Although efficient and scalable, these metrics focus on surface-level properties and are unable to measure fine-grained aspects of text quality (Gehrmann et al., 2023).
-In addition, traditional automatic metrics have been shown to generally correlate weakly with human judgements (Novikova et al., 2017; Reiter, 2018).
-Model-based metrics (Rei et al., 2020; Yuan et al., 2021; Zhong et al., 2022) have addressed these issues to some extent by fine-tuning pre-trained models to evaluate text beyond lexical overlap.
-While these metrics generally correlate better with human judgments compared to overlap-based metrics, they still suffer from several limitations.
-Specifically, they provide limited explainability by providing only numeric scores, require large amounts of labeled training data, and are often limited to evaluating specific tasks or aspects (e.g., Kryscinski et al., 2020; Mehri & Eskénazi, 2020).
-To automate error span annotation in data-to-text, Kasner et al.
-(2021) introduce a combination of a rule-based NLG system with a neural retriever model and fine-tuned RoBERTa (Liu et al., 2019).
-Goyal et al.
-(2022) propose a protocol for fine-grained annotation of coherence errors in narrative summarization, applying it to both human and automatic span annotations.
-In MT evaluation, xCOMET (Guerreiro et al., 2024) is a recent model-based metric specialized in error span annotation.
-LLM-as-a-judge.
-With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators across diverse tasks.
-Although their generative capabilities initially have been applied mainly to simple numeric scoring Liu et al.
-(2023); Sottana et al.
-(2023), other approaches include generating textual feedback before their final decision (Li et al., 2024; Kim et al., 2024a; b).
-Relatively few works have addressed evaluation as error span annotation with LLMs.
-To obtain error spans for translation evaluation, GEMBA-MQM (Kocmi & Federmann, 2023) promps an LLM with annotation guidelines based on MQM evaluation protocol.
-InstructScore (Xu et al., 2023) and TigerScore (Jiang et al., 2024a) fine-tune smaller LLMs on synthetic data generated by GPT-4 to generate error spans.
-Although computationally more efficient than prompting larger LLMs, these two metrics either require reference texts, or are limited to predefined evaluation aspects.
-Kartác et al.
-(2025) apply an ensemble of open-weight LLMs as evaluators to a variety of tasks and aspects, generating both numeric scores and error spans with detailed explanations.
-To address potential data contamination, Kasner & Dušek (2024) collect novel structured data and model outputs in five domains, and collect span annotations for semantic accuracy errors with both crowdworkers and GPT-4.
-Goel et al. (2023) apply LLMs for medical text annotation, identifying and extracting spans containing specific named entities.
-Appendix G Examples
-In Tables 26, 25 and 24, we show examples of the annotated outputs for our tasks.
-In Table 27, we show an example of a reasoning trace of Llama 3.3 prompted with $\mathcal{P}_{\texttt{cot}}$.
-In Tables 28 and 29, we show the examples of reasoning traces of DeepSeek-R1.
-Annotated by Llama 3 When the left made Linda Sarsour into its role model, it climbed into bed withfear the worst of the worstlabelling.
-The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundloaded to commit school shootings, prosecutors alleged in court documents Wednesday.
-Annotated by DeepSeek r1 When the left made Linda Sarsour into its role model, it climbed into bed with the worst of the worst.loaded The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compound to commit school shootings, prosecutors alleged in court documents Wednesday.exag.
-Annotated by o3-mini When the left made Linda Sarsour into its role model, it climbed into bed with the worst of the worst.loaded The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundloaded to commit school shootings, prosecutors alleged in court documents Wednesday.
-Annotated by human When the left made Linda Sarsour into its role modellabelling, it climbed into bedloaded with the worst of the worst.exag.
-The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundlabelling to commit school shootings, prosecutors alleged in court documents Wednesday.
-Table 24: Two examples for span annotation of Propaganda outputs with appeal-to-fearfear, name-calling-labellinglabelling, loaded-languageloaded, and exaggerationexag. as span categories.
-Source Stealthily I made my way towards the back of the store, murmurs and talking heard beyond a metal grate which took a few yanks to tear off.
-Climbing into the vent was the safest and quietest option for me right now; climbing my way through, I’d pick up on conversations and talks between strangely dressed men.
-I noticed Al inside, chained.
-Figures he’d get caught eventually.
-Annotated output by Llama3.3 Plížil jsem semajor ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemiminor, které se podařilo po několika pokusech strhnout.
-Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorům mezi podivně oblečenými mužiminor.
-Všiml jsem si Ala uvnitř, v řetězech.
-Je jasné, že se nakonec chytil.major
-Annotated output by DeepSeek r1 Plížil jsem se ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemi, které se podařilo po několika pokusech strhnout.
-Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorůmminor mezi podivně oblečenými muži.
-Všiml jsem si Ala uvnitř, v řetězech.
-Je jasné, že se nakonec chytil.minor
-Annotated output by o3-mini Plížil jsem se ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemiminor, které se podařilo po několika pokusech strhnout.
-Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorůmminor mezi podivně oblečenými muži.
-Všiml jsem si Ala uvnitř, v řetězech.
-Je jasné, že se nakonec chytil.major
-Annotated output by human Plížil jsem se ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemi, které se podařilo po několika pokusech strhnout.
-Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorům mezi podivně oblečenými muži.
-Všiml jsem si Ala uvnitř, v řetězech.
-Je jasné, že se nakonec chytil.minor
-Source “It’s your birthday soon, isn’t it?” Ivory asked, remembering that the princess’ birthday was in a few days. Kari nodded, a sad glint in her light blue eyes.
-Annotated output by Llama 3 “¿Es su cumpleaños pronto, no lo es?”major Ivory le preguntó,minor recordando que el cumpleaños de la princesa era en unos días.
-Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
-Annotated output by DeepSeek r1 “¿Es su cumpleaños pronto, no lo es?”major Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días.
-Kari nodó, un deslumbramientominor triste en sus ojos azules claros.
-Annotated output by o3-mini “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días.
-Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
-Annotated output by human “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días.
-Kari nodómajor, un deslumbramiento triste en sus ojos azules claros.
-[…]
-Annotated by Llama 3 Borsec will experience a period of overcast skies with light snow expected for the next few days.not check.
-Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.not check.
-There’s a chance of scattered clouds and occasional clear periods throughout the week.
-Winds will be moderate, ranging from 1-3 m/s.mislead.
-Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
-Annotated by DeepSeek r1 Borsec will experience a period of overcast skiescont.
-with light snow expected for the next few days.
-Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.mislead.
-There’s a chance of scattered clouds and occasional clear periods throughout the week.
-Winds will be moderate, ranging from 1-3 m/s.
-Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
-Annotated by o3-mini Borsec will experience a period of overcast skies with light snow expected for the next few days.
-Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.mislead.
-There’s a chance of scattered clouds and occasional clear periods throughout the week.
-Winds will be moderate, ranging from 1-3 m/s.
-Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
-Annotated by human Borsec will experience a period of overcast skies with light snow expected for the next few days.
-Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.cont.
-There’s a chance of scattered clouds and occasional clear periods throughout the week.
-Winds will be moderate,cont.
-ranging from 1-3 m/s.
-Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
-Table 26: Example for span annotation of D2T-Eval in the weather domain with contradictorycont, misleadingmislead., and not-checkeablenot check.
-as error span categories.
-In the annotation interface, the weather panel is interactive and hovering over specific time point shows the temperature and wind conditions.
-Annotated by human annotator The Ituano team won 4-3 against Vila Nova in a match played on November 17th, 2024.
-Ituano scored four goals throughout the game, with Leozinho and Salatiel contributing to their victory.mislead.
-Vila Nova fought back with three goals from Luciano Naninho, Alesson, and Jemmes, but ultimately fell short.
-Annotated by Llama 3 with $\mathcal{P}_{\texttt{base}}$ The Ituano team won 4-3 against Vila Nova in a match played on November 17th, 2024.
-Ituano scored four goals throughout the game, with Leozinho and Salatiel contributing to their victory.mislead.
-Vila Nova fought back with three goals from Luciano Naninho, Alesson, and Jemmesmislead., but ultimately fell short.
-Annotated by Llama 3 with $\mathcal{P}_{\texttt{cot}}$ The Ituano team won 4-3 against Vila Nova in a match played on November 17th, 2024.
-Ituano scored four goals throughout the game, with Leozinho and Salatiel contributing to their victory.
-Vila Nova fought back with three goals from Luciano Naninho, Alesson, and Jemmes, but ultimately fell short.
-Reasoning trace of Llama 3 with $\mathcal{P}_{\texttt{cot}}$ (emphasis added) To identify errors in the text, we need to compare it with the provided data.
-The data contains information about a football match between Ituano and Vila Nova, including the score, goals, and events during the game.
-First, let’s check if the score mentioned in the text matches the data.
-According to the data, the score is indeed 4-3 in favor of Ituano, so this part of the text is correct.
-Next, we look at the goals scored by each team.
-The text mentions that Ituano scored four goals and attributes contributions to their victory to Leozinho and Salatiel.
-From the data, we see that Ituano’s goals were scored by José Aldo, Leozinho, Salatiel, and Vinícius Paiva.
-While Leozinho and Salatiel are correctly mentioned as contributors, the text does not mention José Aldo and Vinícius Paiva, who also scored for Ituano.
-However, this omission does not necessarily make the statement incorrect or misleading since it doesn’t claim to list all scorers.
-The text then mentions that Vila Nova scored three goals from Luciano Naninho, Alesson, and Jemmes.
-According to the data, these players indeed scored for Vila Nova, with their goals being part of the events listed.
-Therefore, this information is accurate based on the provided data.
-Given the information in the text and comparing it with the data, there doesn’t seem to be any direct contradictions (Contradictory), unverifiable claims (Not checkable), misleading statements (Misleading), incoherent phrasing (Incoherent), repetitive information (Repetitive), or other issues that would require annotation.
-Since the text accurately reflects the data provided without introducing any errors or problematic statements, there are no annotations to make.
-Table 27: Example of a reasoning trace of Llama 3.3 prompted with $\mathcal{P}_{\texttt{cot}}$ in the D2T-Eval football domain.
-The model prompted with $\mathcal{P}_{\texttt{cot}}$ dismisses the annotation of the sentence “with Leozinho” that the model prompted with $\mathcal{P}_{\texttt{base}}$ (and the human annotator) annotate as misleadingmislead..
-The screenshot of the data input is truncated.
-Annotated by DeepSeek-R1 The Posh Equal Plus X700 is a 7-inch smartphone with an IPS LCD display and a resolution of 600 x 1024 pixels.
-It features a 5MP main camera with LED flash and supports video recording.
-The device runs on Android 6.0 Marshmallow and has a Qualcomm MT8321 chipsetcont., along with 1GB RAM and 16GB internal storage.
-The Posh Equal Plus X700 is powered by a Li-Po battery with a capacity of 2650 mAh.
-Reasoning trace of DeepSeek-R1 (truncated) Alright, I need to identify errors in the given text based on the provided data.
-Let me go through each sentence one by one.
-First sentence: "The Posh Equal Plus X700 is a 7-inch smartphone with an IPS LCD display and a resolution of 600 x 1024 pixels." Checking the data, under ’Display’, it says size is 7.0 inches and resolution 600x1024.
-So that’s correct.
-Next: "It features a 5MP main camera with LED flash and supports video recording." Data shows Main Camera has 5 MP and LED flash, Video is ’Yes’.
-Correct again.
-Then: "The device runs on Android 6.0 Marshmallow…" Wait, in the data under Platform, OS is Android 6.0 (Marshmallow), correct.
-But chipset is Mediatek MT8321, not Qualcomm.
-The text says Qualcomm MT8321, which is wrong.
-So that’s an error.
-Next part: "…and has a Qualcomm MT8321 chipset, along with 1GB RAM and 16GB internal storage." As mentioned, chipset is incorrect.
-Also, Memory section shows Internal as 16GB 1GB RAM, so that part is correct.
-Last sentence: "The Posh Equal Plus X700 is powered by a Li-Po battery with a capacity of 2650 mAh." Battery details match exactly.
-So no issue here.
-So the only error is the chipset being incorrectly stated as Qualcomm instead of Mediatek.
-This contradicts the data, so it’s a Contradictory error.
-Table 28: The reasoning trace of DeepSeek-R1 prompted with $\mathcal{P}_{\texttt{base}}$ in D2T-Eval gsmarena domain with a contradictorycont annotation.
-The visualization of the input data is truncated, the relevant field is highlighted.
-Input text annotated by DeepSeek-R1 (truncated) Dallas Chief Says She Can’t Fire Amber Guyger for Killing Botham Jean - Dept Rules Imply Otherwiseb-w-fall.
-Dallas, TX — As citizens call for justice in the case of Dallas police officer Amber Guyger — the cop who killed the innocentBotham Jean in his own home — the controversy continues to grow.
-In a town hall meeting this week, citizens showed that they still have a lot of questions for the Dallas Police Department in regards to the cold-bloodedkilling of Botham Jean.
-On top of questions as to why it took several days to arrest the cop — despite the fact that Guyger admitted to it — were questions like why this person who has proven to be a threat to societyhas yet to be fired from the police force.
-During the meeting, as Chief Renee Hall dodged several questions over the details of the arrest, she did go out of her way to answer the question of why Guyger is still collecting a paycheck from their department.
-The short of it, according to Hall, is she can’t fire her.
-[…2,849 characters skipped …]
-Reasoning trace of DeepSeek-R1 (truncated) Okay, so I need to analyze the given text and identify spans that use propaganda techniques.
-Let me go through each paragraph step by step.
-Starting with the title: "Dallas Chief Says She Can’t Fire Amber Guyger for Killing Botham Jean - Dept Rules Imply Otherwise." Hmm, this seems to present two opposing viewpoints without a middle ground.
-It’s implying either she can’t fire her or the dept rules say otherwise, which might be a Black-and-White Fallacy.
-In the first paragraph, it mentions citizens calling for justice and the controversy growing.
-Not sure about propaganda here yet.
-The second paragraph talks about a town hall meeting with questions about the arrest delay and why Guyger hasn’t been fired.
-Still more informative than propagandistic.
-Moving on, Chief Hall dodges questions but explains she can’t fire Guyger due to laws.
-Then experts say this isn’t true.
-This sets up a conflict where the authority (Chief) is contradicted by other experts, which might relate to Appeal_to_Authority if the article is using the experts’ opinions to discredit the Chief without evidence.
-[…1,385 characters skipped …]
-So, I think several propaganda techniques are present: Appeal_to_Authority (using experts), Black-and-White Fallacy in the title, Exaggeration/Minimisation with other officers’ cases, Doubt when questioning Hall’s actions, and Loaded_Language throughout.
-Table 29: Excerpt from the reasoning trace of DeepSeek-R1 prompted with $\mathcal{P}_{\texttt{base}}$ in Propaganda.
-Both the input text (containing the annotated Black-and-White Fallacyb-w-fall.) and the reasoning trace are truncated.
+Table 20: Manual evaluation results for Propaganda domain. See Table˜19 for the legend.
+Figure 13: Confusion matrix for MT-Eval, averaged across models (see Table˜6 for category descriptions).
+Figure 14: Confusion matrix comparing human annotations (rows) with model predictions (columns) for Propaganda, averaged across models. (see Table˜7 for the description of categories).

### v4

@@
 To reflect these aspects in quality judgments of texts, techniques are needed that provide the appropriate amount of detail.
-However, most automatic evaluation metrics for Natural Language Generation (NLG) assign only singular scores for the whole text per each evaluated aspect gkatzia2015snapshot; sai2022survey; schmidtova2024automatic.
-Although numerical values make it easy to rank systems, these metrics are too crude and susceptible to biases or miscalibration of the underlying models gehrmann2022repairing; liu2023calibrating; wang2023large; gao2024llm.
+However, most automatic evaluation metrics for Natural Language Generation (NLG) assign only singular scores for the whole text per each evaluated aspect Gkatzia and Mahamood (2015); Sai et al.
+(2023); Schmidtová et al.
+(2024).
+Although numerical values make it easy to rank systems, these metrics are too crude and susceptible to biases or miscalibration of the underlying models Gehrmann et al.
+(2023); Liu et al.
+(2024); Wang et al.
+(2024); Gao et al.
+(2024).
 The subject of our study, span annotation (Figure˜1), offers an alternative approach.
@@
 Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation.
-The method traditionally required human annotators, making it costly and difficult to scale dasanmartino2019finegrained; thomson2020gold; popovic2020informative; kocmi2024error.
-The LLM-as-a-judge paradigm recently emerged as a promising solution to this problem zheng2023judging; gu2025surveyllmasajudge, allowing task-specific applications kocmi2023gembamqm; hasanain2024large.
+The method traditionally required human annotators, making it costly and difficult to scale Da San Martino et al.
+(2019); Thomson and Reiter (2020); Popovic (2020); Kocmi et al.
+(2024c).
+The LLM-as-a-judge paradigm recently emerged as a promising solution to this problem Zheng et al.
+(2023); Gu et al.
+(2024), allowing task-specific applications Kocmi and Federmann (2023); Hasanain et al.
+(2024).
 However, to our knowledge, no study has systematically compared span annotation performance between LLMs and human annotators.
@@
 We select three span annotation tasks (cf.
-Section˜3.1): evaluating data-to-text generation thomson2020gold, identifying errors in machine translation kocmi-etal-2024-findings, and detecting propaganda techniques in human-written texts dasanmartino2019finegrained.
+Section˜3.1): evaluating data-to-text generation Thomson and Reiter (2020), identifying errors in machine translation Kocmi et al.
+(2024a), and detecting propaganda techniques in human-written texts Da San Martino et al.
+(2019).
 Our contributions are as follows:
   1. 1.
-    We establish that with structured outputs and detailed annotation guidelines, LLMs can serve as robust span annotators (Sections˜3.2 and 5.1).
+    We establish that with structured outputs and detailed annotation guidelines, LLMs can serve as robust span annotators, yielding relevant spans for all three annotation tasks we work with (Sections˜3.2 and 5.1).
   2. 2.
-    We show that LLMs have moderate inter-annotator agreement with human annotators overall, but can reach the agreement of qualified crowdworkers (Section˜5.2).
+    We show that LLMs have moderate inter-annotator agreement with human annotators overall, but can reach the agreement level among verified crowdworkers who passed a qualification task (Section˜5.2).
   3. 3.
-    We discover the sources of model errors by manually analyzing a subset of LLM outputs (Section˜5.3).
+    We discover the sources of model errors by manually analyzing a subset of LLM annotation outputs (Section˜5.3).
   4. 4.
@@
 LLMs for NLG Evaluation.
-Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts sai2022survey; schmidtova2024automatic.
-As such, they are unable to quantify more fine-grained aspects (gehrmann2022repairing; 10.1162/tacl_a_00437) and do not correlate well with human judgments novikova-etal-2017-need; reiter-2018-structured.
-With the emerging LLM-as-a-judge paradigm (gu2025surveyllmasajudge), LLMs have been applied as evaluators on various tasks, using simple numeric scoring (bavaresco2024llms; liu2023geval; sottana2023evaluation; leiter2023eval4nlp; chiang2023can), or free-form feedback (li2024generative; kim2024prometheusa; kim2024prometheus; kartac2025openlgauge).
-However, as these outputs are not firmly grounded in text, they tend to miss fine-grained aspects and are influenced by LLM biases stureborg2024large; koo2023benchmarking; wang2023large.
+Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts Sai et al.
+(2023); Schmidtová et al.
+(2024).
+As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgments Novikova et al.
+(2017); Reiter (2018).
+With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators on various tasks, using simple numeric scoring (Bavaresco et al., 2025; Liu et al., 2023; Sottana et al., 2023; Leiter et al., 2023; Chiang and Lee, 2023), or free-form feedback (Li et al., 2024; Kim et al., 2024a, b; Kartáč et al., 2025).
+However, as these outputs are not firmly grounded in text, they tend to miss fine-grained aspects and are influenced by LLM biases Stureborg et al.
+(2024); Koo et al.
+(2024); Wang et al.
+(2024).
 Span Annotation Protocol.
-In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (lommel2014multidimensional; mariana2014multidimensional; popovic2020informative; kocmi2024error), where human annotators mark erroneous spans in translations.
-In data-to-text (D2T) generation, span annotation was applied by thomson2020gold, who introduced a span-based evaluation protocol for annotation of generated basketball match reports.
-Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection dasanmartino2019finegrained and text summarization subbiah2024reading.
+In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024c), where human annotators mark erroneous spans in translations.
+In data-to-text (D2T) generation, span annotation was applied by Thomson and Reiter (2020), who introduced a span-based evaluation protocol for annotation of generated basketball match reports.
+Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection Da San Martino et al.
+(2019) and text summarization Subbiah et al.
+(2024).
 Unlike our work, these works focus on span annotation with human annotators.
@@
 Early attempts at automating span annotation with ad-hoc guidelines were based on fine-tuned pre-trained encoder models.
-That includes evaluation of MT (10.1162/tacl_a_00683), D2T generation kasner2021textincontext or text summarization goyal2022snac, as well as propaganda detection martino2020semeval2020; goffredo2023argumentbased; piskorski2023multilingual.
+That includes evaluation of MT (Guerreiro et al., 2024), D2T generation Kasner et al.
+(2021) or text summarization Goyal et al.
+(2022), as well as propaganda detection Martino et al.
+(2020); Goffredo et al.
+(2023); Piskorski et al.
+(2023).
 Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities.
-We build on recent work that applies LLMs as a task-specific evaluation tool (kocmi2023gembamqm; fernandes-etal-2023-devil; hasanain2024large; kasner2024traditional; chang2024booookscore; zouhar2024aiassisted; kartac2025openlgauge; ramponi2025finegrained).
-Our work is the first that unifies diverse span annotation tasks within a single framework and systematically compares the performance of LLMs to human annotators.
+We build on work that applies LLMs as a task-specific evaluation tool (Kocmi and Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner and Dušek, 2024; Chang et al., 2024; Zouhar et al., 2025; Kartáč et al., 2025; Ramponi et al., 2025).
+Furthermore, Semin et al.
+(2026) recently investigated various strategies for automatic span annotation with LLMs.
+Our work is the first that systematically compares the performance of LLMs to human annotators.
 3 Automating Span Annotation with LLMs
@@
 $A=\textsc{LLM}(\text{prompt}(Y,C,\mathcal{G},X)).$
-To obtain the annotations, we follow the setup of kasner2024traditional: we request the list of annotations in JSON format, using constrained decoding with a fixed JSON scheme to ensure that the output is syntactically valid.
+To obtain the annotations, we follow the setup of Kasner and Dušek (2024): we request the list of annotations in JSON format, using constrained decoding with a fixed JSON scheme to ensure that the output is syntactically valid.
 We require each annotation to contain the fields reason (the explanation $r_{i}$), text (the textual content of the span), and type (the integer index of the error category $c_{i}$).
-[^2] Following castillo2024structured, we ensure that the reason field is generated first.
+[^2] Following Castillo (2024), we ensure that the reason field is generated first.
 For reasoning models not supporting structured output, we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response.
@@
 To compare annotations automatically, we need a notion of similarity between two sets of annotations $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\hat{\mathcal{A}}=\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $\mathcal{Y}=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$.
-We consider the following similarity metrics:
+Note that basic inter-annotator agreement metrics such as Cohen’s $\kappa$ (Cohen, 1960) are not applicable in our case, as they require a fixed set of annotation units, while the number and position of spans in the span annotation task may differ Mathet et al.
+(2015).
+Therefore, we consider the following similarity metrics:
 Pearson correlation $\rho$ over counts.
@@
 Precision, Recall, and $\mathbf{F_{1}}$.
-To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined by dasanmartino2019finegrained.
+To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined by Da San Martino et al.
+(2019).
 These measures are on matching annotations, adjusted to give partial credit to imperfect matches:
@@
 The $F_{1}$ score is sensitive to varying span granularities and does not consider near-matches with no overlap or agreement by chance.
-To this end, we follow dasanmartino2019finegrained and hasanain2024large in using the $\gamma$ score (mathet2015unified) as a complementary metric.
+To this end, we follow Da San Martino et al.
+(2019) and Hasanain et al.
+(2024) in using the $\gamma$ score (Mathet et al., 2015) as a complementary metric.
 The metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations.
 The score ranges from -$\infty$ to 1, where 1 is achieved when the annotations are perfectly aligned.
-The $\gamma$ score extends Krippendorff’s $\alpha$ (krippendorff1980content), another popular metric, by computing the category-aware span alignments.
-We use the implementation of Titeux2021.
+The $\gamma$ score extends Krippendorff’s $\alpha$ (Krippendorff, 1980), another popular metric, by computing the category-aware span alignments.
+We use the implementation of Titeux and Riad (2021).
 $S_{\emptyset}$ score.
@@
 4.1.1 D2T-Eval: Evaluation of Data-to-text Generation
-In D2T-Eval, we use span annotation to evaluate semantic accuracy and stylistic aspects of data-to-text generation outputs sharma2022innovations; celikyilmaz2021evaluation.
+In D2T-Eval, we use span annotation to evaluate semantic accuracy and stylistic aspects of data-to-text generation outputs Sharma et al.
+(2022); Celikyilmaz et al.
+(2020).
 The inputs $X$ are the structured data used to generate the output text $Y$.
-We use D2T-Eval as a control task to mitigate the effects of data contamination: the fact that the performance of the model might be inflated by previous exposure to publicly available benchmarks balloccu2024leak; dong2024generalization; jiang2024investigating.
-Instead of using an existing dataset, we use the Quintd tool kasner2024traditional to download structured inputs from multiple public APIs.
+We use D2T-Eval as a control task to mitigate the effects of data contamination: the fact that the performance of the model might be inflated by previous exposure to publicly available benchmarks Balloccu et al.
+(2024); Dong et al.
+(2024); Jiang et al.
+(2024).
+Instead of using an existing dataset, we use the Quintd tool Kasner and Dušek (2024) to download structured inputs from multiple public APIs.
 [^3] We selected two of the existing domains: openweather for generating weather forecasts and gsmarena for generating phone descriptions.
@@
 To obtain output texts for the structured data, we prompt LLMs in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences.
+Note that we do not need to deal with the factuality of outputs here, as the sole purpose of the texts is being the input to the annotation process (in fact, having some number of errors is desirable).
 See Appendix B.1 for more details.
 To gather annotations for our dataset, we use crowdworkers from Prolific.com.
-We apply best practices for gathering human annotations, including an iterative process to refine annotation guidelines and preselecting the best-performing annotators using a qualification task tseng2020best; iskender2020best; huang2023incorporating; zhang-etal-2023-needle.
-Our process of collecting annotations proceeded in two stages, following the setup of zhang-etal-2023-needle: (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations.
+We apply best practices for gathering human annotations, including an iterative process to refine annotation guidelines and preselecting the best-performing annotators using a qualification task Tseng et al.
+(2020); Iskender et al.
+(2020); Huang et al.
+(2023); Zhang et al.
+(2023).
+Our process of collecting annotations proceeded in two stages, following the setup of Zhang et al.
+(2023): (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations.
 See Appendix B.2 for more details on collecting annotations.
@@
 4.1.2 MT-Eval: Identifying Errors in Machine Translation
-For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (kocmi2024findings).
-The system outputs were annotated with the Error Span Annotation (ESA) protocol (kocmi2024error) by professional translators.
+For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024b).
+The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024c) by professional translators.
 The inputs $X$ for MT-Eval are the texts in the source language used to produce the translation $Y$.
@@
 The original dataset has nearly 50k model outputs, making it too extensive for our evaluation campaign.
-Therefore, we used a balanced sample of all available system outputs for 10 sampled inputs for each pair of languages, making up 2,854 examples in total.
+Therefore, we used a balanced subsample: For each of the nine language pairs, we randomly sample ten input translation segments.
+We then take all available system outputs for these 90 input segments, making up 2,854 examples in total.
 4.1.3 Propaganda: Propaganda Technique Detection
-For the Propaganda task, we use the dataset of dasanmartino2019finegrained containing news collected mostly from on-line propagandistic sources.
+For the Propaganda task, we use the dataset of Da San Martino et al.
+(2019) containing news collected mostly from on-line propagandistic sources.
 The token-level annotations in the dataset created by expert annotators cover 18 categories of logical fallacies and persuasion techniques.
@@
 For collecting span annotations with LLMs, we use a mixture of open and proprietary state-of-the-art models:
-  • instruction-tuned models: Llama 3.3 70B grattafiori2024llama, GPT-4o hurst2024gpt, and Claude 3.7 Sonnet anthropic2025claude,
-  • reasoning models: [^4] By reasoning models we understand the models that use extra inference time to generate a thinking trace before providing the answer marjanovic2025deepseek.
-DeepSeek-R1 distilled 70B version guo2025deepseek, o3-mini openai2025o3mini, and Gemini 2.0 Flash Thinking google2025geminithinking.
+  • instruction-tuned models: Llama 3.3 70B Grattafiori et al. (2024), GPT-4o Hurst et al. (2024), and Claude 3.7 Sonnet Anthropic (2025),
+  • reasoning models: [^4] By reasoning models we understand the models that use extra inference time to generate a thinking trace before providing the answer Marjanović et al.
+(2025).
+DeepSeek-R1 70B DeepSeek-AI (2025), o3-mini OpenAI (2025), and Gemini 2.0 Flash Thinking Deepmind (2025).
 See Appendix˜A for details on our experimental setup.
@@
 Table 2: Comparison of prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ (#a/o is the average number of annotations per output).
-Figure 2: Comparison of LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators.
-For $F_{1}$ score, the shadow bar denotes its soft variant.
-The dashed horizontal lines for D2T-Eval denote agreement between our human annotators.
-Detailed results are included in Tables˜11, 12, 13 and 14.
+Figure 2: Comparison between LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators.
+Rows represent different tasks (see Section 4.1), columns show different annotation similarity metrics (see Section 3.3).
+For the $F_{1}$ score, the shadow bar denotes its soft variant.
+The dashed horizontal lines denote agreement between our human annotators for D2T-Eval (the agreement is not available for the external datasets).
+More detailed results are included in Tables˜11, 12, 13 and 14.
 Including detailed guidelines seems beneficial: omitting the guidelines ($\mathcal{P}_{\texttt{noguide}}$) lowers the performance of both models.
@@
 Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.
-This observation is aligned with guo2025deepseek, who note that few-shot prompting degrades the performance of DeepSeek-R1.
+This observation is aligned with DeepSeek-AI (2025), who note that few-shot prompting degrades the performance of DeepSeek-R1.
 Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for further experiments.
@@
 Reasoning models outperform instruction-tuned models.
-DeepSeek-R1 generally outperforms the same-sized Llama 3.3.
+DeepSeek-R1 generally outperforms the same-sized Llama 3.3.555The models are comparable as the 70B distilled variant of DeepSeek-R1 is based on Llama 3.3 70B.
+See Section A.1 for details.
 Its superiority is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs.
@@
 Here, o3-mini, Claude 3.7 and Gemini 2.0 mostly reach or surpass human agreement.
-For Propaganda, the upper bound of IAA is the result of dasanmartino2019finegrained, who report $\gamma=0.31$ for annotators before consolidation.
+For Propaganda, the upper bound of IAA is the result of Da San Martino et al.
+(2019), who report $\gamma=0.31$ for annotators before consolidation.
 This score is substantially higher than LLMs (the best LLM score being $\gamma=0.16$ for Claude 3.7 Sonnet).
@@
 To gain more insights into the qualitative aspects of LLM annotations, we manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda.
-[^7] We do not include MT-Eval in the manual analysis due to our insufficient expertise in most target languages.
+[^8] The analysis was split among 7 authors of this paper.
+While we did not do double annotation due to lack of time, we discussed any unclear cases throughout the process.
+We do not include MT-Eval in the manual analysis due to our insufficient expertise in most target languages.
 For each model, we sampled three annotations per category in D2T-Eval and one annotation per category in Propaganda.
@@
 The IAA between LLMs and human annotators is only moderate, suggesting LLMs cannot straightforwardly replace human annotators.
-However, using LLMs may be a reasonable option in scenarios such as D2T-Eval based on crowdworkers, where the strongest LLMs reach the average IAA between human annotators themselves.
+However, using LLMs may be a reasonable option in scenarios based on crowdworkers, where the strongest LLMs reach the average IAA between human annotators themselves.
 In other cases, when deciding whether to employ LLMs as span annotators, one needs to balance desired output quality with other practical aspects.
 Here, LLMs provide better flexibility, shorter response times, and lower costs.
-One should also consider the quality of available human annotators, as even qualified crowdworkers make similar amounts of errors as LLMs.
-It can be also assumed that LLM-based span annotation will benefit from future increases in LLM capabilities, while crowdworkers may increasingly rely on LLM to complete tasks veselovsky2023artificial.
-A promising solution seems to be a hybrid approach in which LLMs pre-annotate the text and humans post-edit the annotations zouhar2024aiassisted.
+One should also consider the quality of available human annotators, as even qualified crowdworkers (i.e., those who passed a qualification task) make similar amounts of errors as LLMs.
+It can be also assumed that LLM-based span annotation will benefit from future increases in LLM capabilities, while crowdworkers may increasingly rely on LLM to complete tasks Veselovsky et al.
+(2023).
+A promising solution seems to be a hybrid approach in which LLMs pre-annotate the text and humans post-edit the annotations Zouhar et al.
+(2025).
 How to deploy LLMs as span annotators?
@@
 In general, reasoning models tend to provide more reliable annotations at the cost of higher response times and token count.
+Is the task meaningful despite the low scores?
+As pointed out by an anonymous reviewer of this paper, the low annotation accuracy – as found by our manual analysis – may indicate a fundamental limitation of the proposed evaluation setup.
+Span annotation is indeed complex and leaves more room for subjectivity than more straightforward annotation such as simple labels or scores.
+However, we argue that the detailed actionable feedback gained through span annotation outweighs the increased noise, both in terms of explainability of the outputs and their potential for further processing.
 Annotations
@@
 Our experiments show that LLMs achieve moderate agreement with skilled human annotators.
-The models perform best in D2T-Eval, where they are comparable to qualified crowdworkers.
+The models perform best in D2T-Eval, where they are comparable to verified crowdworkers who passed a qualification task.
 Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions.
@@
 The estimates are also not readily available for existing datasets and require additional data collection.
+Due to our insufficient expertise in the target languages, we also do not provide language-specific manual error analysis of results.
 As an evaluation method, span annotation is not well-suited for certain NLG evaluation tasks such as annotating omissions or rating the overall text style.
@@
 Acknowledgments
-This work was funded by the European Union (ERC, NG-NLG, 101039303) and the National Recovery Plan funded project MPO 60273/24/21300/21000 CEDMO 2.0 NPO.
+This work was funded by the European Union (ERC, NG-NLG, 101039303).
+It was additionally supported by the National Recovery Plan funded project MPO 60273/24/21300/21000 CEDMO 2.0 NPO and the Charles University Research Centre program No.
+24/SSH/009.
 It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No.
-LM2018101).
-We thank David M. Howcroft for his early input and contributions to the research methodologies adopted in this study.
+LM2023062).
+We thank David M.
+Howcroft for his early input and contributions to the research methodologies adopted in this study.
 Ethics Statement
@@
 We run the local models using the ollama framework in 4-bit quantization.
-Specifically, we use llama3.3:70b and deepseek-r1:70b for span annotations.
+Specifically, we use llama3.3:70b and deepseek-r1:70b (which is based on Llama 3.3 70B) for span annotations.
 We also use gemma2:2b and phi3.5:3.8b for generating texts in D2T-Eval.
@@
 A.3 Web Interface
-We implement our span annotation process using factgenie kasner2024factgenie: a tool that supports both collecting span annotations from humans via a web interface and from LLMs via API calls.
+We implement our span annotation process using factgenie Kasner et al. (2024): a tool that supports both collecting span annotations from humans via a web interface and from LLMs via API calls.
 Figure˜12 shows samples of our annotation interface implemented in factgenie for human annotators, including data visualizations from the football and openweather domains.
@@
 B.1 Generating Outputs
-For generating the outputs for the structured inputs we collected, we use two larger models – Llama 3.3 70B grattafiori2024llama and GPT-4o hurst2024gpt – and two smaller models – Gemma 2 2B team2024gemma and Phi-3.5 3.8B abdin2024phi.
+For generating the outputs for the structured inputs we collected, we use two larger models – Llama 3.3 70B Grattafiori et al.
+(2024) and GPT-4o Hurst et al.
+(2024) – and two smaller models – Gemma 2 2B Team et al.
+(2024) and Phi-3.5 3.8B Abdin et al.
+(2024).
 See more details on the models in Appendix˜A and prompts in Appendix˜D.
@@
 For the annotation guidelines, we went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$.
-We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works kasner2024traditional; thomson2020gold.
+We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works Kasner and Dušek (2024); Thomson and Reiter (2020).
 We settled on the following annotation categories (see Table˜5 for details): semantic accuracy errors due to information Contradictory to the input, Not Checkable, or Misleading; any Incoherent and Repetitive content, and any Other errors.
@@
 We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com.
-Our process of collecting annotations proceeded in two stages, following the setup of zhang-etal-2023-needle: (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations.
+Our process of collecting annotations proceeded in two stages, following the setup of Zhang et al.
+(2023): (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations.
   • Qualification task: For the qualification task, we pre-selected workers whose first language is English, with >99% approval rate and more than 100 submissions.
@@
 Whataboutism A technique that attempts to discredit an opponent“s position by charging them with hypocrisy without directly disproving their argument
-Table 7: Annotation categories for the Propaganda task. The categories are adopted from dasanmartino2019finegrained.
+Table 7: Annotation categories for the Propaganda task. The categories are adopted from Da San Martino et al. (2019).
 Appendix D Prompts